### v1

# VAMP-Net: An Interpretable Multi-Path Framework of Genomic Permutation-Invariant Set Attention and Quality-Aware 1D-CNN for MTB Drug Resistance

Abstract
Genomic prediction of drug resistance in Mycobacterium tuberculosis remains challenging due to complex epistatic interactions and highly variable sequencing data quality. We present a novel Interpretable Variant-Aware Multi-Path Network (VAMP-Net) that addresses both challenges through complementary machine learning pathways. Path-1 employs a Set Attention Transformer processing permutation-invariant variant sets to capture epistatic interactions between genomic loci. Path-2 utilizes a 1D Convolutional Neural Network that analyzes Variant Call Format quality metrics to learn adaptive confidence scores. A fusion module combines both pathways for final resistance classification. We conduct comparative evaluations of unmasked versus padding-masked Set Attention Blocks, and demonstrate that our multi-path architecture achieves superior performance over baseline CNN and MLP models, with accuracy exceeding 95% and AUC of $\sim$97% for rifampicin (RIF) and rifabutin (RFB) resistance prediction. The framework provides dual-layer interpretability: Attention Weight Analysis reveals Epistatic networks, and Integrated Gradients (IG) was applied for critical resistance loci (notably rpoB), while gradient-based feature importance from the CNN pathway uncovers drug-specific dependencies on data quality metrics. This architecture advances clinical genomics by delivering state-of-the-art predictive performance alongside auditable interpretability at two distinct levels—genetic causality of mutation sets and technical confidence of sequencing evidence—establishing a new paradigm for robust, clinically-actionable resistance prediction.
Keywords Set Attention Transformer Block (SAB) $\cdot$ Convolutional Neural Networks (1D-CNN) $\cdot$ Multi-Path Deep Learning $\cdot$ Interpretable Machine Learning $\cdot$ Genomic Variant Analysis $\cdot$ MTB Drug Resistance
1 Introduction
The rapid growth of high-throughput sequencing technologies has transformed genomics into a data-rich discipline, affording unprecedented opportunities to understand complex biological phenomena, such as genotype-phenotype associations and drug resistance mechanisms (sharma2022tuberculosis). Due to the contemporary requirement for personalized therapy within critical applications in Mycobacterium tuberculosis (MTB), drug resistance prediction from whole-genome sequencing (WGS) must be both accurate and interpretable. The prediction of drug resistance involves a number of significant modeling challenges. First, the genetic basis of drug resistance is determined by complex, high-order epistatic interactions among variants, meaning that the models need to learn set-based relationships rather than simple sequential dependencies. Second, raw WGS data from clinical samples are inherently of varying quality, with different coverages and allele frequencies, making it crucial to robustly distinguish the true pathogenic variants from low-confidence sequencing artifacts. In other words, the model should consider the reliability (confidence) of the input data in addition to the causality (biology) thereof. Overcoming the challenges of high-dimensional, variable-length genetic variant sequences and low-quality sequencing input is crucial for developing robust and biologically valid predictive systems (james2025whole; kim2022machine; cosgun2020exploring).
Recent advances in deep learning offer promising avenues for capturing the complex, non-local relationships inherent in genomic data (perelygin2025deep). Although CNNs have been widely applied to capture local patterns in genomic sequences, most often they require fixed-length representations and cannot model higher-order, long-range epistatic interactions among genetic variants that dictate complex phenomena such as drug resistance. This is partly overcome by Transformer-based architectures, which model non-linear, long-range dependencies using self-attention mechanisms with great success in sequence-based domains (wu2023transformer); however, applications to genomic variant sequences remain sparse. Moreover, existing approaches have fallen short, either treating the genome as a dense, image-like sequence (convolutional methods) or by being heavily reliant on expert-defined features, thereby limiting the discovery of novel epistatic relationships. Crucially, most current models treat critical VCF quality metrics as binary hard filters, lacking their treatment as continuous, learnable features that could modulate the model’s confidence in a variant’s biological contribution (eren2023improving).
In this work, we propose a novel Interpretable Multi-Path Fusion Architecture that strategically integrates a Set Attention Block (SAB) pathway with a dedicated, quality-aware 1D-CNN pathway to model drug resistance in MTB. We present a principled integration of a Set Transformer for learning permutation-invariant genetic causality and a 1D CNN for adaptive data quality assessment to overcome the limitations inherent in conventional sequential and feature-engineered models. The set of genomic variants (the reference-to-alternate change) is explicitly encoded with its precise position as a singular, biologically meaningful token (ChromPosi_Ref>AltChromPosi\_Ref>Alt) processed by the SAB pathway to capture complex, high-order epistatic variant-variant interactions under the principle of permutation invariance. The 1D-CNN pathway processes multi-channel representations of critical features in VCF format, (such as FRS and Genotype Confidence), enabling the model to learn an optimal weighting of data quality rather than relying on brittle, fixed thresholds for this task. A fusion module adaptively combines the outputs of these two pathways, leveraging both global, order-agnostic biological insights and local, quality-aware representations for superior predictive performance. This integrated approach extends the Set Transformer methods to genomic variant sequences while embedding a high degree of interpretability via the attention mechanism —a crucial requirement for clinical deep learning research.
Our work introduces the novel application of the Set Transformer architecture to analyze genetic variant sequences, culminating in a robust multi-path framework that significantly outperforms traditional MLP and CNN-only baselines by treating variants as an unordered set. We adapt the SAB in a two-fold manner: first, we establish a theoretically sound baseline by applying a purely permutation-invariant model, SAB without masking (lee2019set), and second, investigate how the same architecture behaves with the incorporation of padding masks while sustaining permutation invariance. Through extensive experiments and an in-depth interpretability analysis, we show how this architectural synergy enables the Set Transformer pathway to identify and map resistance loci accurately, while the parallel 1D-CNN provides a layer of confidence. Taken altogether, our contributions uniquely marry set-based attention modeling, multimodal fusion, and interpretability, bridging a critical gap between high-performance deep learning and actionable genomic modeling for precision medicine.
The rest of the paper is organized as follows. Section 2 summarizes the related work, including some deep learning frameworks and recent progress on deep genomic learning for drug resistance. Section 3 describes the dataset of genetic variants and VCF data used in our research. Section 4 describes the proposed method and strategy in detail, including the novel interpretable two-path transformer and 1D CNN framework. The experimental results and further discussion are presented in Section 5. Finally, Section 6 summarizes our contributions and perspectives on the possible future development of this research area.
2 Related Work
Genome-Wide Association Studies (GWAS) are powerful tools for identifying genetic variants, particularly single nucleotide polymorphisms (SNPs), linked to complex traits or diseases. However, traditional GWAS methods often face limitations in handling the complexity and scale of genomic data, especially in detecting non-linear interactions and epistatic relationships between SNPs. Recent advancements in machine learning (ML) and deep learning (DL) approaches have shown great promise in overcoming these limitations, offering improved accuracy and insights in post-GWAS analysis  (sun2021revisiting; sigala2023machine).
One of the primary advantages of ML over traditional statistical methods is its ability to identify complex interactions and hidden relationships between SNPs. However, traditional approaches such as linear regression (LR) are often inadequate in handling the high dimensionality and non-linearity of genomic data. Several works (an2020genome; maciukiewicz2018gwas; zhang2012identify) have used penalized regression techniques, like LASSO (L1) and Ridge regression (L2), to improve model interpretation and performance by shrinking coefficients, which helps in identifying significant SNPs with greater accuracy (sigala2023machine). But as mentioned previously and due to the Curse of Dimensionality problem, the computational complexity of applying these methods to genome-wide data was inevitable, consequently leading to the development of hybrid approaches that first pre-select SNPs using statistical methods before applying ML models. Random forests (RF) and gradient boosting machines (GBM) have been widely applied in genotype–phenotype association studies due to their ability to capture non‑linear relationships. For example,(arshadi2009predictive) showed that a GBM model was relatively robust to the inclusion of irrelevant SNPs, making it more tolerant to noise. Bayesian network methods have also been explored for modeling SNP–SNP and SNP–phenotype associations (sigala2023machine; kang2015inferring; ma2022sts). Support Vector Machines (SVMs) and Artificial Neural Networks (ANNs) are also used; while they show promise, benchmarking suggests they do not consistently outperform other methods, and their performance depends strongly on the size and quality of input data (nicholls2020reaching; sigala2023machine). For example, studies on GWAS‑case–control data have shown that encoding schemes (e.g., additive vs. homozygous/heterozygous) may affect classifier performance (mittag2015influence). Artificial Neural Networks have also been applied to SNP‑based classification, with some studies discussing how SNP encoding (e.g., homozygous/heterozygous representation) may affect performance (sigala2023machine). Deep learning models, such as the ExPecto framework (zhou2018deep) have demonstrated significant improvements in predicting the functional effects of SNPs across various tissues. ML methods are particularly effective in post-GWAS prioritization, where the goal is to narrow down the list of significant SNPs and genes identified during the GWAS stage. A variety of models, including logistic regression (LR), SVM, RF, GB, and ANN, have been applied to prioritize disease loci. (nicholls2020reaching) highlighted that random forests were top performers for diseases like amyotrophic lateral sclerosis and chronic kidney disease, emphasizing the importance of ensemble methods in improving prediction accuracy.
Other works adopt the approach of merging several techniques to enhance the performance of genotype–phenotype analysis and SNP prioritization. Combining machine learning models like random forests with traditional statistical methods has proven effective in improving prediction accuracy and scalability. (silva2022machine) demonstrated that random forest-based dimensionality reduction techniques can efficiently shrink the dataset size while preserving key genetic information. In their approach, random forest was used to rank SNPs according to their relevance, followed by clustering of the top-ranked SNPs for more focused analysis. Other studies combined RF (for SNP selection) and used KNN and SVM algorithms to classify the identified SNPs for their association with asthma (gaudillo2019machine). Studies in (boutorh2016complex; boutorh2015classication; boutorh2014grammatical) focused on dimensionality reduction in the genomic context, where their hybrid framework based on association rule mining (ARM), grammatical evolution, and neural networks addressed the “large p, small n” challenge (i.e., large features, relatively few samples) by selecting informative SNP‐sets and reducing feature space prior to classification.
The application of deep learning (DL) models presents significant advancements, particularly in overcoming the limitations associated with ML methods and traditional statistical methods, such as linear models, that struggle with high-dimensional genomic data. One of the primary advantages of DL models (especially CNNs) is their capacity to model non-linear interactions between SNPs, as well as their ability to incorporate epistatic relationships, which can enhance the accuracy of disease association predictions. For instance, (liu2018chromatin) introduced DEOPEN, an artificial neural network that combines a deep convolutional network with a three-layer feed-forward neural network to predict chromatin accessibility and evaluate genetic variants in breast cancer. Authors in (liu2019phenotype) further developed an independent deep CNN model for predicting phenotypic traits from SNP data, employing a saliency map visualization method to identify significant biomarkers and demonstrating improved accuracy over traditional statistical methods. Similarly, (romagnoni2019comparative) utilized dense neural networks with various architectures, achieving competitive mean AUC scores around 0.80 in classification tasks. (wang2019deep) introduced a deep mixed model that integrates CNN and Long Short-Term Memory (LSTM) architectures. Another key example of this is the ExPecto framework developed by (zhou2018deep), which utilizes deep convolutional neural networks, L2-regularized linear models, and spatial feature transformation techniques for in silico prediction of gene expression and disease risk. Similarly, DeepSEA (zhou2015predicting) offers a deep learning framework that uses convolutional layers to predict the effects of noncoding variants by learning regulatory sequences from chromatin-profiling data. It integrates various chromatin information to capture complex interactions, enabling precise single-nucleotide predictions and aiding in prioritizing functional variants like eQTLs and disease-associated variants. DeepWAS (arloth2020deepwas) integrates deep learning predictions of regulatory effects into a multivariate GWAS framework, addressing the challenge of linking genetic variants to their functional roles. By annotating SNPs for regulatory potential and using regularized regression models, DeepWAS reduces the multiple testing burden and incorporates current regulatory predictions. This method has been successfully applied to GWAS data for conditions like multiple sclerosis and major depressive disorder, effectively identifying functionally relevant SNPs. Another model is DeepCombi (mieth2021deepcombi) which addresses the issue of missing heritability in traditional GWAS methods by combining convolutional neural networks (CNNs) with GWAS analyses to capture complex genetic interactions and multi-locus effects. Others like (kwon2022genome) also utilized CNNs to develop CNN-GWAS that focus on enhancing the early diagnosis of atrial fibrillation (AF). Another example is the GenNet framework (van2021gennet) designed to tackle the complexities of interpreting genetic data in GWAS while addressing computational challenges. It managed to outperform traditional methods, achieving significant accuracy and identifying both well-replicated and novel genes linked to traits such as hair color and schizophrenia.
Transformer-based architectures are also widely used with genomic data, particularly for their ability to handle large sequence data, which is often the case with long DNA sequences. One of the leading models is DNABERT, introduced in (ji2021dnabert), a pre-trained Bidirectional Encoder consisting of 12 transformer blocks, designed to capture k-mer representations of DNA for downstream tasks such as promoter and enhancer recognition. Another work introduced LOGO (yang2022integrating), a lightweight language model that combines self-attention and convolutional layers to efficiently interpret non-coding regions of the human genome at high resolution, demonstrating improved performance with fewer parameters. More recent developments have extended these approaches by integrating hybrid convolution–transformer designs, hierarchical attention mechanisms, and domain-specific pretraining strategies to better capture long-range dependencies while maintaining computational efficiency.
Research in (green2022convolutional) trained convolutional neural networks on variants from 18 genomic loci across $10\,\text{k}$ M. tuberculosis isolates to predict resistance to 13 drugs; their multi-drug CNN (MD-CNN) and single-drug CNNs achieved high AUCs and, using saliency mapping, recovered known resistance sites while nominating novel candidate positions for further validation. (kuang2022accurate) compared traditional ML classifiers and CNNs on whole-genome sequence–derived features to build a fast, practical pipeline for calling TB drug resistance; they showed that carefully tuned classical models can match deep models on many tasks and presented workflows optimized for clinical turnaround. GWAS-derived SNP sets and deep learning (CNN) were used in (kwon2022genome) to predict phenotype (atrial fibrillation) from genotype, demonstrating how GWAS hits can be repurposed as structured inputs for predictive models and highlighting pitfalls and opportunities of transferring association signals into individual-level predictors. A recent research work (saliba2025enhanced) introduced group-association modeling to jointly model correlated resistance phenotypes across pathogens, leveraging shared patterns to boost sensitivity for multi-drug resistance detection while keeping model complexity low. (deelder2019machine) present flexible machine-learning frameworks that take large numbers of WGS-derived predictors to classify resistant vs susceptible M. tuberculosis isolates; their results demonstrate robust predictive performance across drugs and emphasize scalable feature handling and interpretability for clinical usage . (groschel2021gentb) deliver GenTB, a user-facing ML pipeline and web tool that integrates WGS inputs with trained models to predict TB resistance profiles; they benchmark the system against other tools, prioritize usability for labs, and provide transparent outputs to aid interpretation in clinical workflows . TB-DROP (wang2024tb), an end-to-end deep learning model that consumes whole-genome mutation matrices to predict drug resistance and provide a lightweight GUI; the work focuses on clinical applicability with fast per-sample runtime and validate performance on diverse cohorts. An MDPI Diagnostics study systematically evaluates multiple ML strategies on mutation-based inputs for resistance prediction, comparing feature engineering pipelines and classifiers, and discussing limits in generalization that underscore the need for diverse training cohorts and robust external validation (paredes2025predicting).
In summary, recent advancements in genomic data analysis reflect a clear evolution from traditional GWAS and classical machine learning methods toward more sophisticated hybrid and deep learning architectures. These modern approaches increasingly integrate statistical modeling, feature selection, and end-to-end representation learning to capture the complex, non-linear relationships inherent in genomic data. Despite these advances, several key challenges remain unresolved—particularly those concerning model interpretability, cross-population generalizability, and the scalability of learning algorithms to high-dimensional, heterogeneous variant data. Addressing these challenges requires the development of architectures capable of unifying biological interpretability with data-driven representation power. The following sections describe the dataset and our proposed interpretable multi-path deep learning framework, designed specifically to meet these needs.
3 Data Description
Data for this study were obtained from Comprehensive Resistance Prediction for Tuberculosis: an International Consortium (CRyPTIC) (cryptic2022data). This globally important resource consists of a robust set of 12,289 clinical isolates of Mycobacterium tuberculosis, collected from 23 countries. Importantly, the dataset provides matched whole-genome sequencing (WGS) information with quantitative Minimum Inhibitory Concentration (MIC) measurements for 13 essential antitubercular drugs, covering the entire treatment landscape: first-line agents (rifampicin, isoniazid, ethambutol), key second-line drugs (amikacin, kanamycin, rifabutin, levofloxacin, moxifloxacin, ethionamide), and newer/repurposed therapeutics (bedaquiline, clofazimine, delamanid, linezolid).
The genetic data is provided in variant call format (VCF) files with variants mapped against the M. tuberculosis H37Rv reference genome (NC000962.3). Each VCF file contains several key features:
  • GT (genotype) indicates the called allele at each position;
  • DP (depth) represents total read depth at the variant site;
  • COV_REF and COV_ALT provide reference and alternate allele coverage respectively;
  • DPF (depth fraction) indicates the proportion of reads supporting the variant;
  • GT_CONF (genotype confidence) gives the likelihood ratio for the called genotype;
  • GT_CONF_PERCENTILE normalizes this confidence score; and
  • FRS (a custom score) reflects variant quality.
These features allow for detailed quality assessment of variant calls and enable filtering based on sequencing depth, allele balance, and call confidence.
The phenotypic data includes both continuous MIC values and binary resistance classifications (resistant/susceptible) based on epidemiological cutoff values, with 88% of isolates having complete profiles across all 13 drugs. The dataset is enriched for drug resistance, with 55.4% of isolates resistant to at least one drug. The complete dataset is publicly available at ftp.ebi.ac.uk/pub/databases/cryptic/release_june2022/.
4 Methodology
Our approach introduces a novel Variant-Aware Multi-Path Network (VAMP-Net) that synergistically couples a Set Transformer architecture with a quality-aware 1D-Convolutional Neural Network for genetic variant analysis. This framework is designed as a two-branch fusion model (Figure 1).
The first branch, a Set Transformer, is specifically engineered to process the genetic variant sequence as an unordered set of tokens. This branch employs the Set Attention Block (SAB) to capture the complex non-sequential relationships between the variants, reflecting the inherent permutation-invariant nature of a set.
The second branch of our framework is a quality-aware 1D-CNN. This path focuses on local sequential features. For each variant, it takes as input a corresponding quality score sequence. The 1D-CNN is efficient in capturing the pattern in these local quality signals, thus offering the model critical information on the confidence and reliability of a variant call.
Figure 1: Overview of the Variant-Aware Multi-Path Network (VAMP-Net) Architecture
Finally, the outputs of the Set Transformer and the 1D-CNN branches are fused through a late-fusion mechanism, as pesented in the Figure 1. The representations from both branches are concatenated and passed to a final classification layer. This fusion allows the model to make predictions based on a rich, multi-faceted understanding of the data: the global, set-based relationships between variants from the SAB, and the local, sequential quality signals from the 1D-CNN. This multi-scale learning (local + global) provides both robustness (resistance to noisy data) and biological interpretability (attention highlighting key variants).
To formulate the problem addressed, a genomic sample is represented as a set of $N$ genetic variants: $V=\{v_{1},v_{2},...,v_{N}\}$ where each variant $v_{i}$ is described by $(\textbf{g}_{i},\textbf{f}_{i})$, with $\textbf{g}_{i}$ representing a sequence of genetic alteration and $\textbf{f}_{i}\in\mathbb{R}^{d}$ denoting the associated feature vector extracted from Variant Call Format (VCF) files. The prediction task is binary classification: $y=f_{\theta}(V)$, where $y\in\{0,1\}$ indicates resistance or susceptibility to a given drug, and $f_{\theta}$ is a parameterized deep learning model.
4.1 Data Representation: Genetic Variants as a Set of Tokens
The dataset used in this study comprises genetic variant sequences of Mycobacterium tuberculosis (MTB), as described in Section 3. Each sample is represented by a set of all identified genetic variants, resulting in sequences of variable length.
Transformer-based architectures have fundamentally transformed the processing of sequential data, exhibiting remarkable ability to capture long-range dependencies and contextual relationships. In the genomics domain, recent transformer models have achieved promising results; however, most existing approaches primarily rely on reference DNA sequences or K-mer tokenization as inputs (consens2025interpreting). While such representations effectively capture local sequence patterns, they often treat variants as abstract tokens embedded within a continuous sequence. This abstraction can obscure the explicit biological relationships and combinatorial interactions among individual genetic alterations—relationships that are essential for understanding complex traits such as drug resistance.
To address this limitation, we propose a variant-centric representation that explicitly encodes both the molecular identity and genomic position of each variant as a single, biologically meaningful token. In this framework, each token directly represents a specific reference-to-alternate allele substitution, enabling the model to learn variant-level interactions rather than merely sequence-level dependencies. This transition from sequence-level attention to variant-level interaction attention provides a more interpretable and biologically grounded foundation for modeling genotype–phenotype associations. Specifically, we employ an attention mechanism that operates on a ChromPos_Ref>AltChromPos\_Ref>Alt representation, designed to capture biologically meaningful variant–variant interactions. In this representation, each genetic variant $\textbf{a}_{k}$ is uniquely identified by its chromosome, precise genomic position, reference allele, and alternate allele (e.g., chr1:12345:A>>G). This compact and canonical identifier serves as a semantically rich token, allowing each sample’s genomic profile $\textbf{g}_{i}$ to be expressed as a discrete set of variant tokens T: $\textbf{g}_{i}=\{a_{1},a_{2},...,a_{T}\}$. Such a representation enables the transformer to reason directly over meaningful biological entities—facilitating the discovery of high-order, non-linear relationships among variants that underlie phenotypic outcomes such as drug resistance.
This representation is particularly well-suited for a permutation-invariant modeling framework. By adopting this symbolic encoding, the entire collection of genetic variants for each sample can be treated as a set rather than an ordered sequence. This design allows the model to learn complex, non-linear dependencies and interactions among variants independent of their order in the dataset or genome file. The proposed tokenization strategy therefore serves as a key enabler of our set-based learning paradigm, marking a distinct shift from conventional sequence-oriented approaches and focus on the intrinsic relationships between variants themselves.
4.2 Model Architecture: Variant-Aware Multi-Path Network (VAMP-Net)
Variant-Aware Multi-Path Network (VAMP-Net), a dual-path deep learning framework specifically designed to address the limitations of conventional single-stream models in genomic data analysis. VAMP-Net is architected to process two distinct sources of information essential for robust variant-based prediction ( Figure 1): (1) the symbolic identity and biological context of each genetic variant and (2) the quality metrics of quantitative sequencing that reflect data reliability.
The rationale for this dual-path design is twofold: first, to leverage the inherent set-based nature of a sample’s variant, where order is non-critical, by using a modern attention mechanism; and second, to capture local patterns and dependencies within the associated quality signals using convolutional processing.
The framework is structured as follows:
  • Path-1 – Symbolic and Topological Stream: Processes the set of genomic variants using a symbolic tokenization scheme at variant-level, allowing the model to learn relationships regardless of input order.
  • Path-2 – Feature and Sequential Stream: Processes a structured sequence of quantitative and contextual features (e.g., read depth, genotype quality) using one-dimensional convolutional layers to extract local quality patterns.
The outputs from both pathways are subsequently fused to form an integrated latent representation, which is passed to a final prediction head. This fusion allows VAMP-Net to jointly reason over global, biologically meaningful variant interactions and local, quality-aware signals, thereby offering a comprehensive and interpretable foundation for genomic variant modeling.
4.2.1 Path-1: Set Attention Transformer Block (SAB) on Genetic Variants
We address the modeling of genetic variant sequences for downstream tasks of drug resistance prediction. Each sample $\mathcal{X}_{i}$ consists of a variable-sized set $\mathcal{X}_{i}$ = $\{x_{1},x_{2},...,x_{T}\}$ where each element $x_{k}$ represent a specific genetic variant. Since $\mathcal{X}_{i}$ is an unordered set, the architecture must satisfy permutation invariance, formulated in Equation 1:
$f(\pi\mathcal{X}_{i})=f(\mathcal{X}_{i}),\quad\forall\pi\in S_{n},$ (1)
where $\pi$ denotes a permutation operator on the sequence elements. The first branch of our model is a Set Transformer, a deep learning architecture specifically designed to operate on sets presented in Figure 2. Its core component, the Set Attention Block (SAB), replaces the standard transformer’s multi-head attention mechanism. Unlike standard transformers which rely on positional encodings and often causal masks for sequential data, the SAB is inherently permutation-invariant. This means the model’s output is independent of the order of the input tokens. For our genetic variant sequences, where the order is arbitrary, the attention mechanism’s ability to model interactions among all elements in a set without relying on their positions is a critical advantage. This enables the model to learn a unified representation of the set, a property that is essential for our application.
Figure 2: Path-1 Architecture: Set Attention Transformer Block (SAB) for Symbolic Variant Processing
The SAB pathway models variant-variant interactions while maintaining permutation invariance. For an input sequence represented by a matrix $\textbf{X}\in\mathbb{R}^{n\times d_{model}}$, where $n$ is the sequence length and $d_{model}$ is the embedding dimension, the SAB computes attention scores between all pairs of tokens in the input sequence, followed by a weighted sum of the values. Mathematically (Equation 2), for an input set of tokens:
$\text{Attention}(\textbf{Q},\textbf{K},\textbf{V})=\text{softmax}\left(\frac{\textbf{Q}\textbf{K}^{\top}}{\sqrt{d_{k}}}\right)\textbf{V}$ (2)
where $d_{k}$ is the dimension of the keys. Unlike standard transformers, no causal mask is applied to preserve permutation invariance. The output of SAB is aggregated using a pooling operator $\rho$ as formulated in Equation 3:
$\textbf{z}_{\text{SAB}}=\rho(\text{SAB}(\textbf{X}))$ (3)
We apply the Set Attention Block (SAB) in two distinct configurations, which were investigated to assess the impact of explicit padding utilization:
I - Unmasked Set Attention:
The first contribution adapts the Set Transformer framework (lee2019set) of the unmasked SAB for modeling genetic variant sequences represented as unordered tokens. Unlike standard transformers for sequential data, no attention masks —neither causal masks is applied, ensuring each variant can attend to every other variant without positional or causal constraints, nor padding masks is applied. This model serves as a theoretically pure baseline for evaluating the efficacy of a purely permutation-invariant approach. The input to the model is a batch of padded sequences, where each variant is represented by its token embedding. Within the SAB, the self-attention mechanism computes attention scores for every token pair, including those involving the meaningless padding tokens. The model learns to encode the relationships between all tokens, with no explicit instruction to ignore the padded ones. This approach tests the model’s capacity to learn a meaningful representation of the data despite the presence of noisy, non-informative tokens. The training objective is to perform a specific task of Drug Resistance, based on their set of variants. The hyperparameters for this model, including the number of SAB layers, hidden dimensions, and learning rate, are provided in the Experimental Section 5.4 .
II- Padding-Masked Set Attention :
The second contribution is the introduction of a padding mask to the attention mechanism, a crucial modification that enables efficient batching and improves computational performance. The padding mask is a binary matrix that masks out attention to the padded tokens. It is applied by adding a large negative value to the attention scores of padded tokens prior to the softmax operation.
To handle sequences of varying lengths, we introduce in Equation 4 a padding mask $\textbf{M}\in\{0,-\infty\}^{N\times N}$, where:
[…]
The attention computation becomes as defined in Equation 5:
$\text{softmax}\left(\frac{\textbf{QK}^{\top}+\textbf{M}}{\sqrt{d_{k}}}\right)\textbf{V}$ (5)
which masks out contributions from padded positions but preserves invariance over valid variants. Since $M$ only removes invalid tokens (not reordering them), the equivariance property still holds over the subset of valid elements (Equation 6):
$\text{SAB}_{\text{masked}}(\textbf{P}\textbf{X})=\textbf{P}\text{SAB}_{\text{masked}}(\textbf{X}),$ (6)
$\mathbf{P}$ a permutation matrix that reorders the tokens in $\mathbf{X}$. Provided $M$ is permuted identically as formulated in Equation 7. $\mathbf{M}^{\prime}$ represents the permuted version of the padding mask $\mathbf{M}$ that must be used when the input tokens $\mathbf{X}$ are permuted by the matrix $\mathbf{P}$:
$M^{\prime}=PMP^{\top}.$ (7)
The application of the padding mask $\mathbf{M}$ provides crucial theoretical and practical advantages for processing variable-length variant sets:
  • Elimination of Irrelevant Signal: It ensures irrelevant (padded) tokens do not contribute to attention score computation, thereby preventing the model from dedicating capacity to learning representations for non-existent variants.
  • Mitigation of Attention Dilution: By zeroing out the influence of padding, the mask prevents artificial self-attention dilution within smaller variant sets. This maintains the true relative importance of valid tokens, particularly when a large portion of the input is padding.
  • Preservation of Set Invariance: Since the mask is content-based (identifying the padding symbol) and not positionally fixed, it does not interfere with the core permutation invariance of the Set Attention Block, provided the mask is permuted identically with the input tokens.
4.2.2 Path-2 : Quality-Aware 1D-CNN
A second branch of our model consists of a 1D-CNN that processes the quality scores associated with each variant (Figure 3). These scores, which are typically sequential and local, provide crucial information on the confidence of each variant call. A 1D-CNN is an ideal choice for this task, as it excels at capturing local patterns, such as the relationship between a variant’s read depth (DP) and its genotype confidence (GT CONF). The 1D-CNN branch processes the variant feature matrix $\textbf{F}\in\mathbb{R}^{N\times d}$.
Figure 3: Path-2 Architecture: Quality-Aware 1D Convolutional Neural Network (1D-CNN) for VCF Feature Processing
While the core genetic information (CHROM, POS, REF, ALT) is fundamental, the true richness of VCF files for deep learning lies within their FORMAT fields. These fields provide critical quantitative and qualitative information about the confidence and context of each variant call, moving beyond a simple binary presence or absence of a mutation. This depth of information, often underutilized in conventional genomic deep learning approaches, represents a powerful, biologically relevant multi-channel input for sophisticated models. Using these detailed metrics, a deep learning model can gain a more nuanced understanding of the reliability and characteristics of each detected variant, which is paramount for accurate genotype-phenotype association studies, especially when dealing with potentially noisy or low-coverage sequencing data from real-world clinical samples.
The proposed framework specifically harnesses seven VCF FORMAT fields as distinct input channels for the Convolutional Neural Network (1D-CNN). When extended to multi-channel inputs, a 1D-CNN can simultaneously process multiple features per genomic position, with each VCF field representing a distinct channel. This allows CNN to learn intricate local patterns and dependencies across these different quality and genotype metrics for each variant, rather than relying solely on the raw genotype information.
4.2.3 Multi-Path Fusion Mechanism
The final stage of the VAMP-Net architecture involves the integration of the learned representations from the two specialized paths. The outputs of the two branches—the SAB’s permutation-invariant set representation ($\mathbf{z}_{\text{SAB}}$) from Path-1, and the 1D-CNN’s quality-aware feature vector ($\mathbf{z}_{\text{CNN}}$) from Path-2 —are concatenated to form a comprehensive, fused representation $\mathbf{z}_{\text{fused}}$ (Equation 8):
$\mathbf{z}_{\text{fused}}=[\mathbf{z}_{\text{SAB}};\mathbf{z}_{\text{CNN}}]$ (8)
This late-fusion approach is crucial, as it allows each branch to specialize independently in its unique data modality (symbolic variant topology vs. quantitative feature sequence) before their insights are combined. The fused vector is then passed through a fully connected network to a final output layer for the prediction of drug resistance.
To determine the optimal strategy for combining these disparate signals, we investigate three distinct modulation mechanisms between the output of Path-1 (transformer) and Path-2 (1D-CNN):
  1. 1.
    Suppression (De-emphasis of Quality):
    The first fusion mechanism is inspired by LSTM’s forget gate control. The purpose of this gate in LSTM is to control how much of the old state is remembered in the new state. By taking the sigmoid of the current input $C_{t}$ and the hidden state $H_{t-1}$, values are transformed into the probability domain, and then multiplied element-wise with $C_{t-1}$, the past internal state.
    We modify the purpose and formula of this gate to suit our case: by taking the sigmoid of Path-2, we use the additional features to control the confidence of the current variants, and then multiply the result by the output of Path-1. The Suppression fusion is defined in Equation 9, where $\odot$ denotes element-wise multiplication.
    $\mathbf{g}=\sigma(\mathbf{z}_{\text{CNN}})$
    $\mathbf{z}_{\text{fused}}=\mathbf{g}\odot\mathbf{z}_{\text{SAB}}$ (9)
    This formulation applies a multiplicative suppression to the output of Path-1 ($\mathbf{z}_{\text{SAB}}$). Since the modulation score $\mathbf{g}$ is constrained to $[0,1]$, the resulting fused vector $\mathbf{z}_{\text{fused}}$ will have element values ranging between 0 and the corresponding values in $\mathbf{z}_{\text{SAB}}$. In effect, Path-2’s quality features provide a scalar factor that selectively scales the influence of each variant, and if the quality metrics indicate low confidence, the variant’s representation can be driven toward 0, effectively performing a dynamic, data-driven form of variant feature selection during the fusion process.
  2. 2.
    Amplification (Emphasis on Quality):
    The second investigated fusion strategy, Amplification, is designed to explicitly test the hypothesis that quality and low-level features (Path-2) provide a critical signal gain to the core symbolic variant representation (Path-1). This mechanism ensures that high-quality, high-confidence variants can have proportionally increase in impact, thereby emphasizing a robust subset of the data.
    Unlike the Suppression mechanism, this formulation guarantees that the symbolic representation $\mathbf{z}_{\text{SAB}}$ is always passed through to the final output, regardless of the quality metrics. The Quality-Aware vector $\mathbf{z}_{\text{CNN}}$ acts as an additive gain control. We adapt the gating mechanism as defined in Equation 10, where $\odot$ denotes element-wise multiplication.
    $\mathbf{g}=\sigma(\mathbf{z}_{\text{CNN}})$
    $\mathbf{z}_{\text{fused}}=(1+\mathbf{g})\odot\mathbf{z}_{\text{SAB}}$ (10)
    This formulation achieves a soft integration that preserves the full contribution of the core variant signal. By introducing the constant $1$ into the scalar factor $(1+\mathbf{g})$, the mechanism guarantees a baseline preservation, ensuring that the fused output $\mathbf{z}_{\text{fused}}$ is never less than the original symbolic representation $\mathbf{z}_{\text{SAB}}$. As the modulation score $\mathbf{g}$ is constrained to the range $[0,1]$, the total scalar factor $(1+\mathbf{g})$ remains bounded within the interval $[1,2]$. This results in a controlled form of additive gain, where the most robustly supported variants—those with high-quality metrics—receive a maximum scaling of $200\%$ of their original magnitude. This approach avoids the risk of zeroing out potentially relevant variants, allowing all identified mutations to participate in the final prediction while prioritizing those with the highest associated confidence.
  3. 3.
    Adaptive Amplification & Suppression
    The final fusion strategy is Adaptive Amplification & Suppression, designed to provide the model with a dynamic, context-aware control over the Path-1 signal. This mechanism combines the feature selection capability of Suppression with the signal enhancement capability of Amplification, allowing the influence of the symbolic variant representation ($\mathbf{z}_{\text{SAB}}$) to be either scaled down or scaled up based on the quality metrics ($\mathbf{z}_{\text{CNN}}$).
    We synthesize the previous two modulation strategies by modifying the gating factor. Using the same modulation score $\mathbf{g}=\sigma(\mathbf{z}_{\text{CNN}})$, the fused vector is calculated as Equation 11.
    $\mathbf{z}_{\text{fused}}=(2\odot\mathbf{g}-1)\odot\mathbf{z}_{\text{SAB}}$ (11)
    where $\odot$ denotes element-wise multiplication.
    This adaptive mechanism creates a bipolar scaling factor which introduces a powerful, new operational dynamic. As the modulation score $\mathbf{g}$, derived from the quality features, is constrained to the range $[0,1]$, the resulting scaling factor $(2\odot\mathbf{g}-1)$ is dynamically mapped to the closed interval [1,1][-1,1]. This allows the model to continuously transition between Suppression Zone (noise filtering) and Amplification Zone (signal gain) based on the input quality. Specifically, when quality features are poor ($\mathbf{g}\approx 0$), the scaling factor approaches 1-1, resulting in a negative scaling of the variant representation, effectively acting as an aggressive noise filter by reversing the polarity of low-confidence signal. Conversely, when quality features are excellent ($\mathbf{g}\approx 1$), the factor approaches $1$, leading to maximum positive scaling and thus prioritizing the most robustly supported variants. This ability to span both positive and negative scalar values allows the model to not only ignore noise but to actively penalize uncertain variants, providing a comprehensive, context-aware control signal to the final prediction layers .
4.3 Interpretability
Following the development of the Variant-Aware Multi-Path Network (VAMP-Net), we conducted a critical interpretability analysis to identify genomic variants and sequencing features most significantly associated with RIF and RFB resistance in Mycobacterium tuberculosis. Our approach is uniquely tailored to the dual-pathway model architecture, employing methods that respect the inherent structure of each stream .
4.3.1 Dual-Path Interpretability Framework
We analyzed both pathways of the VAMP-Net architecture separately to isolate the contribution of symbolic variant identity from that of quantitative quality metrics:
Path-1 (Set Transformer Analysis): For the SAB pathway, which processes the symbolically tokenized and permutation-invariant set of variants ($\mathbf{z}_{\text{SAB}}$), we applied Integrated Gradients (IG). This method calculated attribution scores by integrating gradients from baseline input to actual variant embeddings, measuring how each genetic variant contributed to the final resistance prediction irrespective of its ordering in the input set.
Path-2 (CNN Analysis): For the convolutional pathway processing the ordered VCF feature vector ($\mathbf{z}_{\text{CNN}}$), we employed gradient-based saliency analysis in conjunction with systematic feature ablation. The gradient analysis quantified the relative importance of feature channels (e.g., DP, GT, FRS) across the sequential input. Subsequently, systematic feature ablation confirmed the findings by quantifying the drop in predictive performance when specific VCF quality metrics were masked or perturbed. This framework reliably quantifies both the positional importance of variants along the input sequence and the absolute contribution of specific VCF quality metrics.
4.3.2 Variant Importance Quantification
For Path-1, which utilizes the Set Attention Transformer, we quantified the contribution of each individual variant by applying Integrated Gradients (IG) with respect to the final classification layer output. We first extracted the variant-level embeddings from the trained symbolic tokenizer. The attribution scores were then calculated by integrating the gradients along a linear path over $N=20$ integration steps, ensuring robust and stable gradient estimates. Finally, the calculated attribution scores were aggregated across all instances in the test cohort to obtain the mean importance measure for each unique ChromPos_Ref>>Alt variant. This yields an interpretable ranking of variants based on their model-predicted association with drug resistance.
4.3.3 Feature Ablation Studies
To robustly quantify the influence of sequencing quality and context features (processed by Path-2), we conducted systematic feature ablation on individual VCF channels. This approach, designed to isolate the causal contribution of each metric, involved methodically setting the values of a single VCF feature channel to a neutral baseline (zero) across all test samples . We then rigorously measured the resulting change in the model’s prediction confidence and overall performance (e.g., AUC or $F_{1}$ score).
This systematic process allowed us to identify which VCF quality metrics were most critical for accurate resistance classification and, conversely, which features introduced minimal signal. The ablation analysis specifically covered the following eight VCF feature channels, reflecting key data quality and genotype metrics: Genotype Calls (GT), Read Depth (DP), Depth Fraction (DPF), Reference Coverage (COV_REF), Alternative Coverage, (COV_ALT), Fraction of Supporting Reads (FRS), Genotype Confidence (GT_CONF), Confidence Percentiles (GT_CONF_PERCENTILE).
The resulting performance decrement for each ablated feature serves as its direct measure of importance within the VAMP-Net architecture.
4.3.4 Attention-Based Variant Interaction Analysis
To move beyond individual variant importance and uncover complex epistatic relationships between resistance-associated variants, we leveraged the intrinsic relational mechanism of the Set Attention Block (SAB) layer from Path-1. The SAB’s self-attention matrix inherently encodes the model’s learned pairwise dependency structure between every variant in the input set.
For each test sample, we systematically extracted the attention weights from the first SAB layer and averaged them across all attention heads to compute the raw variant relationships. The interaction strength between any two variants was thus quantified as the bidirectional attention weight between their respective tokens, which was then aggregated across the entire test cohort for statistical reliability.
This aggregated matrix was subsequently used to construct a variant interaction network. Network construction was restricted to variants exhibiting sufficient frequency and strong, recurring interaction occurrences to ensure statistical reliability. Finally, we applied modularity optimization algorithms to the network to perform community detection, thereby identifying functional modules (groups of variants) that likely operate synergistically to confer drug resistance. The highly connected nodes identified in this process were designated as hub variants, representing key drivers of the overall epistatic effect.
5 Experimentation and Discussion
This section presents the empirical validation of the Variant-Aware Multi-Path Network (VAMP-Net) and its core components, followed by a detailed discussion of the biological insights derived from our interpretability framework. We first detail the preparation of the dataset and the specific encoding utilized for the ChromPos_Ref>Alt variant tokens. Subsequent subsections systematically evaluate three critical aspects of the design of VAMP-Net: the selection of the optimal fusion mechanism, a comprehensive performance comparison, and a deep-dive interpretation of the predictions of the final model, including variant importance and epistatic interaction networks. The goal of this analysis is not only to demonstrate superior classification performance but to validate the translational utility of our dual-path approach for genetic resistance prediction.
5.1 Data Preprocessing
Prior to use, the data is subjected to multiple preparation procedures. After downloading the CRyPTIC_reuse_table_20240917.csv file from the official compendium website, we chose to focus on two drugs: Rifampicin (RIF), which is primarily used to treat tuberculosis by inhibiting RNA polymerase, and Rifabutin (RFB), a related rifamycin derivative with similar therapeutic applications.
First, we filter the dataset to include only high and medium quality samples, based on the {drug}_PHENOTYPE_QUALITY field. This results in 10,310 samples for RIF and 11,392 samples for RFB. For each selected sample, we download its corresponding VCF file, which contains variant information. Then we remove columns with missing data and retain only entries marked as PASS in the FILTER field, indicating reliable variant calls.
To reduce dimensionality and reliability, we focus on positions where the ALT allele differs from the REF allele, discarding entries such as 0/0 and 0|0, which indicate homozygous reference (no variant), as well as ./. and .|., which denote missing or uncalled genotypes, it should be noted that the slash (/) unphased genotype, while the pipe ($|$) indicates a phased genotype. Variants are represented as pos_ref>>alt for all alleles separated by commas; we refer to this representation as the Path-1 Input, after this filtering, we end up with a total of 417,635 unique variants.
We also retain additional descriptive features for each variant: [’GT’, ’DP’, ’DPF’, ’COV’, ’FRS’, ’GT_CONF’, ’GT_CONF_PERCENTILE’]. Numeric features are scaled using min–max normalization, while categorical features are encoded. The FRS field, representing the fraction of reads supporting the genotype call (#reads supporting GT / total reads), ranges between 0 and 1; Higher FRS values indicate stronger confidence in the genotype, whereas missing or low values typically correspond to low coverage or filtered calls, and hence we chose to replace its missing values with zeros. For each sample, these features are organised into an array of shape (number of variants × 8 features × 1), referred to as the Path-2 Input, which will be used later in a 1D-CNN model. Arrays are padded as needed to ensure uniform length across samples.
It should be noted that the resulting samples are not aligned with each other. For example, sample one may have 20_A>G at index 0, while sample two has 13_C>T at the same index. This indicates that the variants are not ordered strictly by their genomic positions in the input vectors. Originally, the variants were sorted by position in the VCF file, but after filtering and preprocessing, some variants shifted in the vectors. This positional inconsistency across samples motivates the use of a Permutation-Invariant Block (SAB) in the model architecture.
Figure 4: Distribution of classes for the selected drugs.
Figure 4 shows the class distribution of the two drugs, with a difference of $\sim$17. 6% between the susceptible and resistant class for RIF, and $\sim$30. 4% for RFB, requiring us to employ a weighted cross-entropy loss to handle this class imbalance.
We will first use the RIF data in all subsequent experiments, navigating several design choices to determine the optimal approach, and subsequently apply the selected method to RFB for validation.
5.2 Variants Encoding
As mentioned in Section 5.1, variants are represented as Pos_REF>ALT, with the ALT for each allele separated by a comma. We compare a Static Encoding method , where each variant is given a unique ID to represent it, against training a BERT Tokenizer on all unique variants , and then using it to encode each sample’s variants into vectors before training.
The results showed that Static Encoding led to overfitting in the long run, with early stopping triggered at epoch 3. The training curve clearly showed signs of overfitting, with a $\sim$6.73% gap between training and validation accuracy.In contrast, using BERT proved effective in delivering more stable learning and improved validation performance. Although it also showed signs of overfitting, with a 4.83% difference between training and validation accuracy at the end of training, it maintained a smoother learning curve and reached early stopping at epoch 16.
The BERT-based encoding was more robust and critical for stable learning as shown in Figure 5. Consequently, we selected BERT for the encoding phase in all subsequent experiments.
Figure 5: Comparison of Static Encoding vs BERT Tokeniser metrics over epochs
5.3 The Choice of Fusion Mechanism
We tested three types of fusion strategies between Path-1 and Path-2, namely: Suppression, Amplification, and Suppression-Amplification (see Section 4.2.3). The training and validation curves are shown in Figure 6 and the test results are summarizsed in Table 1.
All three types achieved comparable performance in terms of validation and test accuracy. However, in general, the Amplification fusion yielded the best results and, therefore, it was selected as the preferred mechanism for the final configuration, and we will refer to this model as - Model A -.
Table 1: Comparison of fusion mechanisms. Best results in bold
Fusion Type Accuracy Precision Recall F1 AUC
Suppression & Amplification 0.9267 0.9322 0.9439 0.9381 0.9572
Amplification 0.9523 0.9510 0.9598 0.9554 0.9690
Suppression 0.9438 0.9460 0.9591 0.9525 0.9710
Figure 6: Training (top) and validation (bottom) performance curves for fusion methods
5.4 Models Experimentation
Starting from - Model A - , we evaluated a set of configurations to measure the robustness of the padding/masking strategy, data augmentation, and to assess the benefits of including Path-2 in the architecture. Moreover, we compared with a traditional approach, adopted in many studies, which involved representing the data as a matrix of shape (Number of VVariants) $\times$ (Number of Samples), where cell $i,j$ is 1 if Variant-$i$ is present in the sample, otherwise 0. This representation is typically trained with classifiers using either classical machine learning or deep learning techniques such as CNNs or MLP-based models. This hybrid of simple binary encoding and deep models has been widely surveyed in antimicrobial resistance (AMR) and clinical genomics literature and has been effective in practice (e.g., CNN and ML studies for tuberculosis and other pathogens) (sakagianni2024data). However, one limitation of this approach is the large resulting sparse matrix due to the high number of variants, i.e., the curse of dimensionality, requiring extensive computational resources. Some studies addressed this by pre-eliminating variants using statistical methods, or dimensionality reduction techniques like PCA (reshetnikov2025feature; pikalyova2024predicting; xu2025machine).
To select informative variants, we applied a Chi-squared test to assess the association between each variant and the target phenotype. Variants with $p$-value <0.001<0.001 were retained, resulting in 11,567 ($\sim$36.1% reduction), which were then used to train and evaluate both a multi-layer perceptron (MLP) and a convolutional neural network (CNN).
To systematically evaluate the contributions of the proposed components, we developed and tested several distinct model architectures:
  1. 1.
    Model-A: SAB with Attention Masks for Padding, without Augmentation.
  2. 2.
    Model-B: SAB without Attention Masks for Padding, with Augmentation.
  3. 3.
    Model-C: SAB without Attention Masks for Padding, without Augmentation.
  4. 4.
    Model-D: SAB without Attention Masks for Padding, with Augmentation.
  5. 5.
    Model-E: Model-A settings, without the inclusion of Path-2.
  6. 6.
    Model-F: Model-B settings, without the inclusion of Path-2.
  7. 7.
    Model-G: MLP for Presence/absence Variants data.
  8. 8.
    Model-H: CNN for Presence/Absence Variants data.
Table 2: Best hyperparameters of -Path-1 - for each proposed Multi-Path model applied on RIF Drug + Model A on RFB Drug
Model Emb Dim Hidden Dim Num Layers
Model A 64 32 3
Model B 64 32 2
Model C 128 32 1
Model D 64 32 3
Model E 128 32 4
Model F 64 32 1
Model A (RFB) 128 32 1
Table 3: Best hyperparameters of - Path-2 - and the Classification head for each proposed Multi-Path model applied on RIF Drug + Model A on RFB drug
Model Dropout Activation Conv Layers Conv Kernel LR
Model A 0.1128 relu 3 3 0.00137
Model B 0.0962 gelu 2 5 0.00165
Model C 0.3817 gelu 1 5 0.00130
Model D 0.0205 gelu 3 3 0.00101
Model E 0.0692 relu 0.00105
Model F 0.3634 gelu 0.00154
Model A (RFB) 0.2704 relu 4 5 0.00106
Table 4: Best hyperparameters of MLP model aplied on RIF Drug
Model Hidden Dims Dropout Activation
Model G [1024, 128, 128] 0.1315 gelu
Table 5: Best hyperparameters of CNN model applied on RIF Drug
Model Dropout Activation Conv Layers Filters Kernel Sizes Pool Sizes
Model H 0.1487 relu 3 [96, 128, 256] [5, 9, 2] [3, 3, 3]
The proposed models were evaluated across multiple architectures and hyperparameter configurations. Tables 2, 3, 4, and 5 summarize the best-performing hyperparameters identified through Bayesian optimization. This tuning prioritized a balance between ROC AUC and Accuracy on the validation set, ensuring that models not only fit the training data but also generalized robustly.
Figure 7: Comparison of -Model A- and -Model B- performance without and with variant sequence shuffling for training and validation
In Figure 7, we compare the performance of Model-A and Model-B where Model-B applies data augmentation via variant sequence shuffling. This augmentation fully leverages the SAB block’s capacity, allowing predictions independent of variant position. Training curves confirm that Model-B, likely due to the regularization effect of data augmentation, exhibits greater training stability as evidenced by significantly reduced overfitting and closer convergence between the training and validation curves. On the test set, Model-A achieved the highest performance (ROC AUC: 0.969, Balanced Accuracy: 0.945). While Model-B’s test performance was marginally lower (ROC AUC: 0.964, Balanced Accuracy: 0.941), this minor trade-off in peak performance resulted in a substantially more robust and stable model. This suggests that data augmentation effectively improves generalization and stability with only a negligible decrease in maximum predictive power.
Figure 8: Training (top) and validation (bottom) performance curves of all models with early stopping
Figure 9: ROC AUC Test curve of VAMP-Net vs CNN and MLP
Figure 8 presents the training and validation curves across epochs for eight distinct models (Model-A through Model-H), offering critical insight into their convergence, learning capacity, and generalization ability.
All models exhibit rapid initial learning, achieving 85% to over 90% accuracy within the first 5 to 10 epochs. The models can be broadly categorized into three performance tiers based on their final stabilized validation accuracy and AUC. Models A, B, C, and D demonstrate the highest potential, peaking around 94%. Models G and H show comparatively lower learning capacity for this task, suggesting architectural limitations and insufficient model capacity. Models G and H required the longest training durations, this is a characteristic of models with less efficiency.
Model-A achieved the highest overall training and validation accuracy and AUC within the shortest training time, making it the most efficient high-performer. This demonstrates that Path-2 inclusion significantly improved robustness and that SAB-based transformers outperform classical CNN/MLP representations. The lower performance of Models G and H suggests that these architectures are less suitable for the current task, as shown in Figure 9.
Figure 10: Models test results comparison
The results, visualized in Figure 10, clearly demonstrate the superior performance of the Set Attention Block (SAB)-based transformer architectures. While Model-A achieved the highest overall performance, the marginal variation across other SAB variants underscores the robustness of the core attention mechanism. Notably, the impact of explicit masking appears minimal, as evidenced by the comparable performance of Models C and D, suggesting that the SAB structure inherently handles or ignores padding tokens effectively. In contrast, the significant underperformance of the traditional Multilayer Perceptron (MLP) (Model G) and Convolutional Neural Network (CNN) (Model H) strongly validates the utility of a sequential attention-based representation over standard binary encoding or fixed-filter approaches for this task.
Based on its performance across all evaluation metrics, we selected Model-A as the optimal architecture. To validate the model’s capacity for broad applicability, we applied this model on the RFB drug. On this drug, Model-A demonstrated high efficacy, achieving an AUC of 0.9681 and an Accuracy of 0.9393. Furthermore, the model maintained a well-balanced classification profile, yielding an F1 score of 0.9535, a Recall of 0.9553, and a Precision of 0.9517. These results collectively confirm the model’s excellent generalization capacity across different therapeutic agents.
The comprehensive architectural evaluation leads to the following four key conclusions:
  • Superiority of Attention: Set Attention Block (SAB)-based transformer architectures significantly outperform standard Convolutional Neural Network (CNN) and Multilayer Perceptron (MLP) baselines on presence/absence data.
  • Enhanced Generalization: Augmentation (as employed in Model-B) effectively improves training stability and promotes permutation invariance, leading to more robust models.
  • Architectural Robustness: While empirically minor, the inclusion of attention masks provides critical robustness against padding tokens, formalizing the model’s ability to handle variable-length sequences.
  • Broad Applicability: The high performance achieved on the RFB drug dataset validates the broad applicability and transferability of the SAB-based transformer approach to new, unseen drugs.
5.5 Multi-Path Framework Analysis and Interpretation
This section prensents the results of the Dual-Path Interpretability Framework and leverages the insights to dissect the model’s decision-making process for RIF and RFB resistance prediction. Our analysis demonstrates that the two architectural pathways within VAMP-Net provide distinct yet complementary perspectives on the underlying mechanisms of drug resistance. Specifically, Path-1 (Set Attention Transformer) successfully identified and quantified variant-level importance and epistatic dependencies, thus elucidating the core genetic drivers. In contrast, Path-2 (Quality-Aware CNN) focused on quantifying the influence of technical confidence metrics, revealing how the model performs noise regularization and feature calibration. By integrating the findings from both streams, our framework provides a comprehensive and auditable explanation of the model’s decision process —both biological and technical— that drives the final drug resistance classification.
5.5.1 Set-Transformer Genetic Variant Importance
This section details the quantitative analysis of individual variant contributions to drug resistance, as determined by the Integrated Gradients (IG) method applied to the SAB pathway (Path-1). The IG-derived attribution scores, aggregated across the test cohort, establish an empirically derived ranking of mean importance for all unique ChromPos_Ref>>Alt variants. This analysis serves to validate the model’s ability to prioritize known resistance drivers while simultaneously discovering highly-ranked non-canonical variants that warrant further investigation. The subsequent analyses distinguish between the variant importance profiles learned for Rifampicin (RIF) and Rifabutin (RFB), highlighting the model’s discriminatory power.
A) Drug 1: RIF Resistance Analysis
The analysis of the VAMP-Net pathway for RIF resistance achieved strong validation success, presented in Table 6, demonstrating the framework’s ability to learn and prioritize fundamental genetic mechanisms as reported in the literature (telenti1993detection; ramaswamy1998molecular). The identified variants in rpoB, katG, and other genes align with established resistance mechanisms documented in the WHO catalogue of mutations associated with drug resistance (world2023catalogue).
Table 6: Among 200 Top-Ranked Variants from RIF Resistance Analysis out of 417,635 Unique Variants
Variant Importance Score Gene
761139_C>>A 0.178 rpoB
761101_A>>T 0.116 rpoB
761140_A>>T 0.058 rpoB
761154_TCG>>TTC 0.050 rpoB
761108_GGA>>GTT 0.044 rpoB
760314_G>>T 0.029 rpoB
2155162_ATGCCGC>>ACGCCGT 0.446 katG
2155166_CGC>>CTG 0.043 katG
7581_G>>T 0.050 gyrA
4247574_A>>C 0.036 embB
1849861_G>>T 0.207 ethA
2338202_CAC>>CCAC 0.308 ponA1 (novel)
424791_T>>TAT 0.273 pknB (novel)
The model assigned primary predictive importance to variants within the RNA polymerase $\beta$ subunit (rpoB) gene. Multiple high-confidence variants in Table 6, including 761139_C>>A, 761101_A>>T, and 761140_A>>T, are located within the rpoB gene and fall precisely within the Rifampicin Resistance-Determining Region (RRDR). This alignment with established clinical and molecular biology literature validates the framework’s capacity to learn fundamental resistance pathways (musser1995antimicrobial).
Beyond the primary rpoB target, the VAMP-Net successfully extracted complex, multi-gene resistance signatures. The model assigned high importance scores to variants associated with isoniazid resistance, including mutations in the katG region (2155162_ATGCCGC>>ACGCCGT) and the promoter of ahpC (78519_G>>A) (vilcheze2007mechanism; napier2023large). This demonstrates the framework’s ability to capture complex multi-drug resistance profiles where resistance to one drug often co-occurs with markers for another (jia2021analysis; billows2024large).
A significant finding is the identification of top-ranked variants located in previously uncharacterized genomic regions (Novel loci). The high Importance Scores for these novel variants (such as ponA1 and pknB) suggest the model is uniquely capable of isolating:
  • Rare, High-Effect Mutations: Highly penetrant variants that evade detection by standard frequency-based GWAS methods.
  • Critical Interaction Hubs: Loci that function as epistatic modifiers, significantly altering the expression or effect of primary resistance mutations.
In summary, the high attribution results confirm the robustness and superior discovery potential of the VAMP-Net, which can simultaneously validate known mechanisms and identify high-impact novel loci for targeted functional investigation.
B) Drug 2: RFB Resistance Analysis
The RFB resistance analysis successfully identified the complete spectrum of variants associated with the phenotype (Table 7), confirming the model’s high fidelity in identifying key resistance mechanisms.
Table 7: Among 200 Top-Ranked Variants from RFB Resistance Analysis out of 417,635 Unique Variants
Variant Importance Score Gene
761140_A>>T 6.896 rpoB
761139_C>>A 3.955 rpoB
761140_A>>G 3.321 rpoB
761110_A>>G 3.316 rpoB
761110_A>>T 2.463 rpoB
4247431_G>>T 2.125 embB
7581_G>>A 1.323 gyrA
2944085_C>>T 4.947 mmpL5 (novel)
893702_GAC…CCG 4.213 lpqS (novel)
3640407_G>>A 3.532 ppsD (novel)
3545402_A>>G 3.527 fadD26 region (novel)
1262023_C>>T 2.885 pks13 (novel)
4217696_TGGAG>>TTGAA 2.594 papA4 (novel)
1262193_C>>T 2.589 pks13 (novel)
The analysis demonstrated the framework’s superior capability to identify both canonical and non-canonical resistance mechanisms. As expected, rpoB coding region variants ranked most highly, with the top-scoring variant 761140_A>T (Score 6.896) confirming the model’s alignment with the shared target mechanism between RFB and RIF (williams1998contribution; van2013rifampin; wang2023antimicrobial). Multiple additional rpoB variants achieved high scores, including 761139_C>A (Score 3.955), 761140_A>G (Score 3.321), 761110_A>G (Score 3.316), and 761110_A>T (Score 2.463), all clustering within the Rifampicin Resistance-Determining Region (RRDR).
Furthermore, the framework successfully captured a Multi-Drug Resistance (MDR) genomic signature from a single drug prediction task. The detection of relevant variants in embB (4247431_G>T) and gyrA (7581_G>A) confirms the learned importance of co-selection and cross-resistance patterns (ghosh2020survey).
The discovery potential of the framework is underscored by the high ranking of numerous novel variants. High scores for variants in genes such as mmpL5 (2944085_C>T, Score 4.947), lpqS (893702_GAC...CCG, Score 4.213), ppsD (3640407_G>A, Score 3.532) reinforce the strong suggestion that these uncharacterized loci represent crucial epistatic interaction points or rare, high-effect genes that contribute significantly to the RFB phenotype. These variants are now prime candidates for functional validation to uncover new biological resistance pathways.
5.5.2 Genetic Variant Interaction Network Analysis
Analysis of the aggregated Set Attention Block (SAB) weights culminates in the Multi-locus Attribution-Weighted Interaction (MAWI) Map (Figures 11 and 12), which provides a critical, quantifiable decomposition of complex, multi-locus epistatic effects underlying Rifampicin (RIF) and Rifabutin (RFB) resistance. This approach moves beyond single-variant association models to capture the necessary synergistic coordination between mutations. Specifically, the heatmaps visualize the highest magnitude of attention-weights found by the SAB transformer, thereby mapping the strength and direction of these interactions and providing mechanistic insights into the polygenic architecture of drug resistance.
Figure 11: Variant interaction heatmaps showing attention-based relationships between top hub variants for RIF resistance. Red indicates strong interactions; yellow indicates weak interactions. The patterns reveal how variants work together in resistance mechanisms
The variant interaction heatmap (Figure 11) for RIF (Rifampicin) resistance reveals a highly interconnected, yet structured, network of epistatic relationships among the top hub variants, with a high overall signal density indicating numerous strong-to-moderate synergistic interactions. This high-interaction density suggests that robust resistance relies on compensatory or cooperative pathways, confirming that the model has effectively learned the structural dependencies of a highly interconnected resistance mechanism. The pattern is characterized by specific "hub" variants, such as 3829770_T>>C and 1307598_C>>G, which act as central nodes, exhibiting dark red cells that indicate strong attention to multiple other variants in the network. This hub-driven structure supports the idea that the mechanism relies on a few key "linchpin" mutations that coordinate with a set of secondary, supportive mutations, and this strong cross-attention quantitatively supports the hypothesis of epistatic fitness compensation. Furthermore, the scattered distribution of high-value cells across the map, rather than being localized to a single cluster, suggests a diffuse polygenic mechanism where many different combinations of mutations contribute significantly to the resistance phenotype.
Figure 12: Variant interaction heatmaps showing attention-based relationships between top hub variants for RFB resistance. Red indicates strong interactions; blue indicates weak interactions. The patterns reveal how variants work together in resistance mechanisms
The second heatmap in Figure 12 (blue-to-red scale, max interaction $\approx 0.40$) depicts a contrasting, sparse network of high-magnitude, localized epistatic effects for RFB resistance. In contrast to Figure 11, this visualization shows a predominantly low-interaction background (dark blue), highlighting a few critical, high-leverage interactions that dominate the attention landscape involving variants like 1070702_T>>C. The high attention-weights (dark red) pattern suggests that these specific variant interactions act as a primary mechanism for resistance, where the effect of one variant is highly dependent on the presence of the other. The sparsity indicates that, while many variants are present, only a select few combinations are mechanistically determinative, allowing the VAMP-Net model to pinpoint the most non-additive synergistic effects.
The contrasting patterns between the two drugs demonstrate that resistance involves coordinated resistance modules whose composition varies depending on the specific drug, providing a key insight for designing combination therapies. The MAWI provides an interpretable, quantitative visualization of complex genetic epistasis, confirming known biological interactions while highlighting high-impact, potentially novel interaction hubs.
5.5.3 VCF Feature Channel Analysis: Quantifying the Influence of Variant Quality Metrics
We quantified the contribution of the eight VCF features via systematic ablation, validating the unique multi-channel design of the Quality-Aware 1D-CNN pathway (Figure 13). This analysis highlights the technical robustness required for reliable genotype-phenotype association.
The analysis confirms the model does not rely on simple variant presence (GT field) but critically leverages the contextual quality and support metrics:
  • Dominant Metric (FRS): The Fraction of Supporting Reads (FRS) is the single most influential feature for both the RIF and RFB models. This strong reliance confirms that the 1D-CNN pathway effectively learns to filter out false positives or low-support variants, as high FRS is used as a primary filter for calling true clonal mutations in MTB sequencing (cryptic2022data).
  • Contextual Confidence: The GT_CONF_PERCENTILE is the second most important feature. By utilizing the relative confidence rank rather than the raw GT_CONF (which shows minimal contribution), the model demonstrates its capacity to leverage comparative, sample-specific context for prediction.
Figure 13: VCF feature importance comparison for RIF and RFB resistance models. Fraction of Supporting Reads (FRS) and Genotype Confidence Percentile (GT_CONF_PERCENTILE) demonstrate substantially higher importance for both resistance phenotypes, with the RFB model showing greater dependence on these quality metrics
The near-zero composite scores for features like Read Depth (DP), Genotype (GT), and Raw GT_CONF are equally informative. The low importance of DP suggests its information is effectively captured by the derived, normalized metrics (FRS and GT_CONF_PERCENTILE), rendering the raw depth redundant. This indicates that the presence of a variant or its raw confidence is insufficient; the model requires the granular, multi-dimensional context of the variant’s support and relative quality to make high-confidence resistance predictions.
This ablation study confirms that the 1D-CNN pathway successfully extracts biologically and technically meaningful information from the VCF features, enabling a robust distinction between true resistance-conferring variants and lower-quality calls, especially critical for subtle phenotypes like RFB resistance.
5.5.4 Dual-Path Fusion Model Interpretability
This final analysis validates the Multi-Path Fusion strategy as the core functional advancement of VAMP-Net, establishing its ability to combine Pathogenicity (via the Set-Transformer) and Reliability (via the 1D-CNN).
The analysis revealed two novel findings critical to the framework’s superior performance and generalizability:
  • Learned Adaptive Feature Regularization: The 1D-CNN pathway functions as a learned signal-to-noise regulator, performing non-linear, context-aware feature selection (prioritizing FRS and GT_CONF_PERCENTILE over redundant features like DP). The most significant evidence of this is the 8 to 14-fold increase in feature reliance observed in the RFB model, which demonstrates that the fusion mechanism dynamically compensates for the weaker, more diverse genetic signal of that phenotype by demanding greater technical certainty. Path-2 thus acts as an adaptive attention mechanism that modulates prediction confidence based on data quality.
  • Epistatic Network Auditing: The SAB network analysis confirmed that the final prediction is governed by an epistatic network audit, where the model learns that a high-importance variant from Path-1 will have its final impact dynamically down-weighted if Path-2 reports low sequencing support. This integrated approach provides an auditable, high-confidence score that addresses the critical challenge of separating biological signal from computational support in clinical genomic data.
In conclusion, integrating VCF features in this manner addresses a major challenge in genomic deep learning: the inherent separation between the biological signal (variant identity) and the computational support (variant quality).
6 Conclusion
We have introduced the Interpretable Variant-Aware Multi-Path Network (VAMP-Net), a novel deep learning framework that successfully addresses two critical, often-overlooked challenges in genomic machine learning: modeling genetic epistasis and systematically incorporating variant calling quality into the prediction process. Our work delivers a significant methodological advance that provides both state-of-the-art predictive performance and crucial layers of auditable interpretability, essential for bridging computational genomics with clinical utility.
The core of VAMP-Net lies in its dual-path architecture and the Adaptive Fusion mechanism. The Set Attention Transformer (Path-1) proved highly effective by operating on the inherently permutation-invariant set of ChromPos_Ref>>Alt variant tokens, successfully elucidating complex genetic epistasis and discovering both canonical resistance drivers and influential non-canonical loci. The Integrated Gradients attribution analysis provides a clear, ranked list of mutations and their co-dependencies, forming a direct basis for targeted functional validation.
However, the most compelling finding lies in the contribution of the Quality-Aware pathway. VAMP-Net is the first model to systematically integrate variant calling quality into resistance prediction, addressing the critical gap where existing methods treat all variant calls equally regardless of technical confidence. Our quantitative feature importance scores demonstrate that the fusion architecture performs a learned feature calibration and noise regularization . Specifically, the pronounced increase in reliance on VCF quality metrics (e.g., FRS, GT_CONF_PERCENTILE) for the more complex Rifabutin (RFB) phenotype, compared to Rifampicin (RIF), confirms the fusion mechanism intelligently and dynamically adapts its weighting based on the complexity and confidence requirements of the task. This dynamic weighting mechanism provides an unprecedented level of robustness.
The VAMP-Net framework’s fundamental design is inherently generalizable and extends well beyond the context of Multi-drug resistance in M. tuberculosis. Our proposal can be applied to any genotype-phenotype problem characterized by three conditions:
  • Variable-Length Inputs: The sample is defined by an unordered set of symbolic events (e.g., structural variants, somatic mutations in cancer, or transcriptomic changes).
  • Epistatic Dependencies: The final phenotype is driven by complex interactions rather than single-gene effects.
  • Heterogeneous Evidence: The symbolic input (genotype) is accompanied by quantitative, context-specific technical metadata (e.g., sequencing depth, coverage, or population frequency).
Future work will thus focus on extending this multi-path fusion paradigm to incorporate additional modalities in other clinical domains, such as incorporating image modalities (e.g., microscopy or histology) or longitudinal clinical metadata alongside genomic data. This work establishes a new standard for confidence and transparency in clinical AI systems by ensuring that all predictions are not only accurate but are driven by a mechanism that is transparent and auditable across both biological and technical evidence streams.
Data availability statements
The complete dataset used in this study is publicly available at ftp.ebi.ac.uk/pub/databases/cryptic/release_june2022/.

### v2

@@
-VAMP-Net: An Interpretable Multi-Path Framework of Genomic Permutation-Invariant Set Attention and Quality-Aware 1D-CNN for MTB Drug Resistance
+VAMP-Net: An Interpretable Multi-Path Network of Genomic Permutation-Invariant Set Attention and Quality-Aware 1D-CNN for MTB Drug Resistance
 Abstract
-Genomic prediction of drug resistance in Mycobacterium tuberculosis remains challenging due to complex epistatic interactions and highly variable sequencing data quality.
-We present a novel Interpretable Variant-Aware Multi-Path Network (VAMP-Net) that addresses both challenges through complementary machine learning pathways.
-Path-1 employs a Set Attention Transformer processing permutation-invariant variant sets to capture epistatic interactions between genomic loci.
-Path-2 utilizes a 1D Convolutional Neural Network that analyzes Variant Call Format quality metrics to learn adaptive confidence scores.
-A fusion module combines both pathways for final resistance classification.
-We conduct comparative evaluations of unmasked versus padding-masked Set Attention Blocks, and demonstrate that our multi-path architecture achieves superior performance over baseline CNN and MLP models, with accuracy exceeding 95% and AUC of $\sim$97% for rifampicin (RIF) and rifabutin (RFB) resistance prediction.
-The framework provides dual-layer interpretability: Attention Weight Analysis reveals Epistatic networks, and Integrated Gradients (IG) was applied for critical resistance loci (notably rpoB), while gradient-based feature importance from the CNN pathway uncovers drug-specific dependencies on data quality metrics.
-This architecture advances clinical genomics by delivering state-of-the-art predictive performance alongside auditable interpretability at two distinct levels—genetic causality of mutation sets and technical confidence of sequencing evidence—establishing a new paradigm for robust, clinically-actionable resistance prediction.
+Genomic prediction of drug resistance in Mycobacterium tuberculosis is often hindered by complex epistatic interactions and variable sequencing quality.
+We present the Interpretable Variant-Aware Multi-Path Network (VAMP-Net), a novel architecture addressing these challenges through a dual-pathway approach.
+Path-1 utilizes a Set Attention Transformer to model permutation-invariant variant sets and capture epistatic dependencies, while Path-2 employs a 1D-CNN to analyze VCF quality metrics for adaptive confidence scoring.Evaluated on four critical anti-TB drugs, VAMP-Net significantly outperforms baseline CNN and MLP models, achieving accuracies >> 95% and AUCs $\sim$0.97 for Rifampicin and Rifabutin.
+Feature attribution analysis via Integrated Gradients successfully recovered canonical targets (rpoB, embB, katG) and discovered high-impact novel loci.
+Functional enrichment confirmed these novel variants constitute non-random metabolic modules (p=0.00239) centered on cell-wall remodeling.Furthermore, systematic ablation of the Quality-Aware pathway demonstrates that the model performs a learned "integrated audit," prioritizing the Fraction of Supporting Reads and relative confidence over raw depth to mitigate technical noise.
+This dual-layer interpretability, bridging genomic pathogenicity with technical reliability, establishes a new paradigm for robust, auditable, and clinically actionable resistance prediction, positioning VAMP-Net as an important tool for both diagnostic classification and mechanistic discovery in clinical genomics.
 Keywords Set Attention Transformer Block (SAB) $\cdot$ Convolutional Neural Networks (1D-CNN) $\cdot$ Multi-Path Deep Learning $\cdot$ Interpretable Machine Learning $\cdot$ Genomic Variant Analysis $\cdot$ MTB Drug Resistance
 1 Introduction
-The rapid growth of high-throughput sequencing technologies has transformed genomics into a data-rich discipline, affording unprecedented opportunities to understand complex biological phenomena, such as genotype-phenotype associations and drug resistance mechanisms (sharma2022tuberculosis).
-Due to the contemporary requirement for personalized therapy within critical applications in Mycobacterium tuberculosis (MTB), drug resistance prediction from whole-genome sequencing (WGS) must be both accurate and interpretable.
-The prediction of drug resistance involves a number of significant modeling challenges.
-First, the genetic basis of drug resistance is determined by complex, high-order epistatic interactions among variants, meaning that the models need to learn set-based relationships rather than simple sequential dependencies.
-Second, raw WGS data from clinical samples are inherently of varying quality, with different coverages and allele frequencies, making it crucial to robustly distinguish the true pathogenic variants from low-confidence sequencing artifacts.
-In other words, the model should consider the reliability (confidence) of the input data in addition to the causality (biology) thereof.
-Overcoming the challenges of high-dimensional, variable-length genetic variant sequences and low-quality sequencing input is crucial for developing robust and biologically valid predictive systems (james2025whole; kim2022machine; cosgun2020exploring).
-Recent advances in deep learning offer promising avenues for capturing the complex, non-local relationships inherent in genomic data (perelygin2025deep).
-Although CNNs have been widely applied to capture local patterns in genomic sequences, most often they require fixed-length representations and cannot model higher-order, long-range epistatic interactions among genetic variants that dictate complex phenomena such as drug resistance.
-This is partly overcome by Transformer-based architectures, which model non-linear, long-range dependencies using self-attention mechanisms with great success in sequence-based domains (wu2023transformer); however, applications to genomic variant sequences remain sparse.
-Moreover, existing approaches have fallen short, either treating the genome as a dense, image-like sequence (convolutional methods) or by being heavily reliant on expert-defined features, thereby limiting the discovery of novel epistatic relationships.
-Crucially, most current models treat critical VCF quality metrics as binary hard filters, lacking their treatment as continuous, learnable features that could modulate the model’s confidence in a variant’s biological contribution (eren2023improving).
-In this work, we propose a novel Interpretable Multi-Path Fusion Architecture that strategically integrates a Set Attention Block (SAB) pathway with a dedicated, quality-aware 1D-CNN pathway to model drug resistance in MTB.
-We present a principled integration of a Set Transformer for learning permutation-invariant genetic causality and a 1D CNN for adaptive data quality assessment to overcome the limitations inherent in conventional sequential and feature-engineered models.
-The set of genomic variants (the reference-to-alternate change) is explicitly encoded with its precise position as a singular, biologically meaningful token (ChromPosi_Ref>AltChromPosi\_Ref>Alt) processed by the SAB pathway to capture complex, high-order epistatic variant-variant interactions under the principle of permutation invariance.
-The 1D-CNN pathway processes multi-channel representations of critical features in VCF format, (such as FRS and Genotype Confidence), enabling the model to learn an optimal weighting of data quality rather than relying on brittle, fixed thresholds for this task.
-A fusion module adaptively combines the outputs of these two pathways, leveraging both global, order-agnostic biological insights and local, quality-aware representations for superior predictive performance.
-This integrated approach extends the Set Transformer methods to genomic variant sequences while embedding a high degree of interpretability via the attention mechanism —a crucial requirement for clinical deep learning research.
-Our work introduces the novel application of the Set Transformer architecture to analyze genetic variant sequences, culminating in a robust multi-path framework that significantly outperforms traditional MLP and CNN-only baselines by treating variants as an unordered set.
-We adapt the SAB in a two-fold manner: first, we establish a theoretically sound baseline by applying a purely permutation-invariant model, SAB without masking (lee2019set), and second, investigate how the same architecture behaves with the incorporation of padding masks while sustaining permutation invariance.
-Through extensive experiments and an in-depth interpretability analysis, we show how this architectural synergy enables the Set Transformer pathway to identify and map resistance loci accurately, while the parallel 1D-CNN provides a layer of confidence.
-Taken altogether, our contributions uniquely marry set-based attention modeling, multimodal fusion, and interpretability, bridging a critical gap between high-performance deep learning and actionable genomic modeling for precision medicine.
+The rapid growth of high-throughput sequencing has transformed genomics into a data-rich discipline, offering unprecedented opportunities to decode complex biological phenomena like drug resistance in Mycobacterium tuberculosis (MTB) [39].
+For clinical applications, drug resistance prediction from whole-genome sequencing (WGS) must be both accurate and interpretable.
+However, modeling MTB resistance presents two primary challenges: (1) the genetic basis is governed by high-order epistatic interactions, requiring models that learn set-based relationships rather than simple sequential dependencies; and (2) clinical WGS data varies significantly in quality, making it crucial to distinguish true pathogenic variants from sequencing artifacts.
+Effectively balancing biological causality with data reliability is essential for developing robust predictive systems [20, 23, 11].
+Recent deep learning advances, particularly Convolutional Neural Networks (CNNs) and Transformers, offer promising avenues for capturing non-local relationships in genomic data [32, 47].
+While CNNs excel at local pattern recognition, they often require fixed-length inputs and struggle with the long-range epistatic interactions inherent in drug resistance.
+Conversely, while Transformers can model these dependencies, their application to unordered sets of genomic variants remains sparse.
+Furthermore, existing models typically treat critical variant calling (VCF) quality metrics as binary hard filters, discarding potentially valuable continuous information that could modulate the model’s confidence in a variant’s biological contribution [13].
+Despite recent advancements, existing models frequently treat genomic variants as rigid sequences and rely on static, "hard" VCF filters.
+This approach is limited to account for the fundamentally permutation-invariant nature of genomic variant sets and the continuous spectrum of sequencing quality metrics.
+Current methods either ignore technical noise entirely or rely on brittle, expert-defined thresholds that cannot capture the nuanced relationship between data confidence and predictive certainty.
+To address these gaps, we propose VAMP-Net, an interpretable multi-path fusion architecture designed to bridge the gap between genomic deep learning and clinical microbiology.
+Our key contributions are summarized as follows:
+  • Expert-Informed Multi-Path Fusion: We introduce a dual-stream architecture that decouples biological signals from technical noise.
+By integrating a Set Attention Block (SAB) pathway with a quality-aware 1D-CNN, VAMP-Net mirrors the "clinical audit" workflow of expert microbiologists, allowing technical confidence to dynamically modulate biological predictions through a Gated Adaptive Fusion mechanism.
+  • Permutation-Invariant Modeling via SAB: Unlike sequence-based Transformers, our model utilizes the Set Attention Block (SAB) to capture high-order epistatic interactions.
+This approach enforces a permutation-invariant inductive bias, ensuring the model respects the biological topology of the isolate as an unordered set of variants rather than an artificial sequence.
+  • Clinically-Grounded Interpretability: We provide a validation framework that maps Path-1 attention weights directly to established resistance hotspots (e.g., katG, rpoB).
+By demonstrating that high-attention regions align with verified microbiology, we confirm that VAMP-Net’s decision-making is driven by biological causality rather than spurious statistical artifacts.
 The rest of the paper is organized as follows.
-Section 2 summarizes the related work, including some deep learning frameworks and recent progress on deep genomic learning for drug resistance.
-Section 3 describes the dataset of genetic variants and VCF data used in our research.
-Section 4 describes the proposed method and strategy in detail, including the novel interpretable two-path transformer and 1D CNN framework.
-The experimental results and further discussion are presented in Section 5.
-Finally, Section 6 summarizes our contributions and perspectives on the possible future development of this research area.
+Section 2 summarizes related work.
+Section 3 details the proposed Multi-Path Network.
+Section 4 presents the experimental results, interpretability analysis, and discussion.
+Section 5 summarizes our contributions and future perspectives.
 2 Related Work
-Genome-Wide Association Studies (GWAS) are powerful tools for identifying genetic variants, particularly single nucleotide polymorphisms (SNPs), linked to complex traits or diseases.
-However, traditional GWAS methods often face limitations in handling the complexity and scale of genomic data, especially in detecting non-linear interactions and epistatic relationships between SNPs.
-Recent advancements in machine learning (ML) and deep learning (DL) approaches have shown great promise in overcoming these limitations, offering improved accuracy and insights in post-GWAS analysis  (sun2021revisiting; sigala2023machine).
-One of the primary advantages of ML over traditional statistical methods is its ability to identify complex interactions and hidden relationships between SNPs.
-However, traditional approaches such as linear regression (LR) are often inadequate in handling the high dimensionality and non-linearity of genomic data.
-Several works (an2020genome; maciukiewicz2018gwas; zhang2012identify) have used penalized regression techniques, like LASSO (L1) and Ridge regression (L2), to improve model interpretation and performance by shrinking coefficients, which helps in identifying significant SNPs with greater accuracy (sigala2023machine).
-But as mentioned previously and due to the Curse of Dimensionality problem, the computational complexity of applying these methods to genome-wide data was inevitable, consequently leading to the development of hybrid approaches that first pre-select SNPs using statistical methods before applying ML models.
-Random forests (RF) and gradient boosting machines (GBM) have been widely applied in genotype–phenotype association studies due to their ability to capture non‑linear relationships.
-For example,(arshadi2009predictive) showed that a GBM model was relatively robust to the inclusion of irrelevant SNPs, making it more tolerant to noise.
-Bayesian network methods have also been explored for modeling SNP–SNP and SNP–phenotype associations (sigala2023machine; kang2015inferring; ma2022sts).
-Support Vector Machines (SVMs) and Artificial Neural Networks (ANNs) are also used; while they show promise, benchmarking suggests they do not consistently outperform other methods, and their performance depends strongly on the size and quality of input data (nicholls2020reaching; sigala2023machine).
-For example, studies on GWAS‑case–control data have shown that encoding schemes (e.g., additive vs.
-homozygous/heterozygous) may affect classifier performance (mittag2015influence).
-Artificial Neural Networks have also been applied to SNP‑based classification, with some studies discussing how SNP encoding (e.g., homozygous/heterozygous representation) may affect performance (sigala2023machine).
-Deep learning models, such as the ExPecto framework (zhou2018deep) have demonstrated significant improvements in predicting the functional effects of SNPs across various tissues.
-ML methods are particularly effective in post-GWAS prioritization, where the goal is to narrow down the list of significant SNPs and genes identified during the GWAS stage.
-A variety of models, including logistic regression (LR), SVM, RF, GB, and ANN, have been applied to prioritize disease loci.
-(nicholls2020reaching) highlighted that random forests were top performers for diseases like amyotrophic lateral sclerosis and chronic kidney disease, emphasizing the importance of ensemble methods in improving prediction accuracy.
-Other works adopt the approach of merging several techniques to enhance the performance of genotype–phenotype analysis and SNP prioritization.
-Combining machine learning models like random forests with traditional statistical methods has proven effective in improving prediction accuracy and scalability.
-(silva2022machine) demonstrated that random forest-based dimensionality reduction techniques can efficiently shrink the dataset size while preserving key genetic information.
-In their approach, random forest was used to rank SNPs according to their relevance, followed by clustering of the top-ranked SNPs for more focused analysis.
-Other studies combined RF (for SNP selection) and used KNN and SVM algorithms to classify the identified SNPs for their association with asthma (gaudillo2019machine).
-Studies in (boutorh2016complex; boutorh2015classication; boutorh2014grammatical) focused on dimensionality reduction in the genomic context, where their hybrid framework based on association rule mining (ARM), grammatical evolution, and neural networks addressed the “large p, small n” challenge (i.e., large features, relatively few samples) by selecting informative SNP‐sets and reducing feature space prior to classification.
-The application of deep learning (DL) models presents significant advancements, particularly in overcoming the limitations associated with ML methods and traditional statistical methods, such as linear models, that struggle with high-dimensional genomic data.
-One of the primary advantages of DL models (especially CNNs) is their capacity to model non-linear interactions between SNPs, as well as their ability to incorporate epistatic relationships, which can enhance the accuracy of disease association predictions.
-For instance, (liu2018chromatin) introduced DEOPEN, an artificial neural network that combines a deep convolutional network with a three-layer feed-forward neural network to predict chromatin accessibility and evaluate genetic variants in breast cancer.
-Authors in (liu2019phenotype) further developed an independent deep CNN model for predicting phenotypic traits from SNP data, employing a saliency map visualization method to identify significant biomarkers and demonstrating improved accuracy over traditional statistical methods.
-Similarly, (romagnoni2019comparative) utilized dense neural networks with various architectures, achieving competitive mean AUC scores around 0.80 in classification tasks.
-(wang2019deep) introduced a deep mixed model that integrates CNN and Long Short-Term Memory (LSTM) architectures.
-Another key example of this is the ExPecto framework developed by (zhou2018deep), which utilizes deep convolutional neural networks, L2-regularized linear models, and spatial feature transformation techniques for in silico prediction of gene expression and disease risk.
-Similarly, DeepSEA (zhou2015predicting) offers a deep learning framework that uses convolutional layers to predict the effects of noncoding variants by learning regulatory sequences from chromatin-profiling data.
-It integrates various chromatin information to capture complex interactions, enabling precise single-nucleotide predictions and aiding in prioritizing functional variants like eQTLs and disease-associated variants.
-DeepWAS (arloth2020deepwas) integrates deep learning predictions of regulatory effects into a multivariate GWAS framework, addressing the challenge of linking genetic variants to their functional roles.
-By annotating SNPs for regulatory potential and using regularized regression models, DeepWAS reduces the multiple testing burden and incorporates current regulatory predictions.
-This method has been successfully applied to GWAS data for conditions like multiple sclerosis and major depressive disorder, effectively identifying functionally relevant SNPs.
-Another model is DeepCombi (mieth2021deepcombi) which addresses the issue of missing heritability in traditional GWAS methods by combining convolutional neural networks (CNNs) with GWAS analyses to capture complex genetic interactions and multi-locus effects.
-Others like (kwon2022genome) also utilized CNNs to develop CNN-GWAS that focus on enhancing the early diagnosis of atrial fibrillation (AF).
-Another example is the GenNet framework (van2021gennet) designed to tackle the complexities of interpreting genetic data in GWAS while addressing computational challenges.
-It managed to outperform traditional methods, achieving significant accuracy and identifying both well-replicated and novel genes linked to traits such as hair color and schizophrenia.
-Transformer-based architectures are also widely used with genomic data, particularly for their ability to handle large sequence data, which is often the case with long DNA sequences.
-One of the leading models is DNABERT, introduced in (ji2021dnabert), a pre-trained Bidirectional Encoder consisting of 12 transformer blocks, designed to capture k-mer representations of DNA for downstream tasks such as promoter and enhancer recognition.
-Another work introduced LOGO (yang2022integrating), a lightweight language model that combines self-attention and convolutional layers to efficiently interpret non-coding regions of the human genome at high resolution, demonstrating improved performance with fewer parameters.
-More recent developments have extended these approaches by integrating hybrid convolution–transformer designs, hierarchical attention mechanisms, and domain-specific pretraining strategies to better capture long-range dependencies while maintaining computational efficiency.
-Research in (green2022convolutional) trained convolutional neural networks on variants from 18 genomic loci across $10\,\text{k}$ M.
-tuberculosis isolates to predict resistance to 13 drugs; their multi-drug CNN (MD-CNN) and single-drug CNNs achieved high AUCs and, using saliency mapping, recovered known resistance sites while nominating novel candidate positions for further validation.
-(kuang2022accurate) compared traditional ML classifiers and CNNs on whole-genome sequence–derived features to build a fast, practical pipeline for calling TB drug resistance; they showed that carefully tuned classical models can match deep models on many tasks and presented workflows optimized for clinical turnaround.
-GWAS-derived SNP sets and deep learning (CNN) were used in (kwon2022genome) to predict phenotype (atrial fibrillation) from genotype, demonstrating how GWAS hits can be repurposed as structured inputs for predictive models and highlighting pitfalls and opportunities of transferring association signals into individual-level predictors.
-A recent research work (saliba2025enhanced) introduced group-association modeling to jointly model correlated resistance phenotypes across pathogens, leveraging shared patterns to boost sensitivity for multi-drug resistance detection while keeping model complexity low.
-(deelder2019machine) present flexible machine-learning frameworks that take large numbers of WGS-derived predictors to classify resistant vs susceptible M.
-tuberculosis isolates; their results demonstrate robust predictive performance across drugs and emphasize scalable feature handling and interpretability for clinical usage .
-(groschel2021gentb) deliver GenTB, a user-facing ML pipeline and web tool that integrates WGS inputs with trained models to predict TB resistance profiles; they benchmark the system against other tools, prioritize usability for labs, and provide transparent outputs to aid interpretation in clinical workflows .
-TB-DROP (wang2024tb), an end-to-end deep learning model that consumes whole-genome mutation matrices to predict drug resistance and provide a lightweight GUI; the work focuses on clinical applicability with fast per-sample runtime and validate performance on diverse cohorts.
-An MDPI Diagnostics study systematically evaluates multiple ML strategies on mutation-based inputs for resistance prediction, comparing feature engineering pipelines and classifiers, and discussing limits in generalization that underscore the need for diverse training cohorts and robust external validation (paredes2025predicting).
-In summary, recent advancements in genomic data analysis reflect a clear evolution from traditional GWAS and classical machine learning methods toward more sophisticated hybrid and deep learning architectures.
-These modern approaches increasingly integrate statistical modeling, feature selection, and end-to-end representation learning to capture the complex, non-linear relationships inherent in genomic data.
-Despite these advances, several key challenges remain unresolved—particularly those concerning model interpretability, cross-population generalizability, and the scalability of learning algorithms to high-dimensional, heterogeneous variant data.
-Addressing these challenges requires the development of architectures capable of unifying biological interpretability with data-driven representation power.
-The following sections describe the dataset and our proposed interpretable multi-path deep learning framework, designed specifically to meet these needs.
-3 Data Description
-Data for this study were obtained from Comprehensive Resistance Prediction for Tuberculosis: an International Consortium (CRyPTIC) (cryptic2022data).
+The evolution of drug resistance prediction in M.
+tuberculosis (MTB) has progressed from simple association studies to complex deep learning frameworks.
+This section categorizes prior efforts and identifies the specific gaps addressed by the proposed Multi-Path architecture, as summarized in the comparative overview in Table 1.
+2.1 Statistical and Classical Machine Learning
+Genome-Wide Association Studies (GWAS) remain the foundation for identifying single nucleotide polymorphisms (SNPs) linked to drug resistance [39, 44].
+Traditional methods, however, struggle with high-dimensional data and non-linear epistatic interactions [41].
+To overcome these, penalized regression techniques such as LASSO and Ridge have been employed to identify significant loci [1, 28].
+Ensemble methods, including Random Forests (RF) and Gradient Boosting Machines (GBM), have demonstrated robustness to genomic noise [29, 27].
+Hybrid approaches have further enhanced performance by using RF for dimensionality reduction followed by clustering or Support Vector Machines (SVM) for classification [42, 14].
+Prior studies in [5, 4, 3] focused on dimensionality reduction through hybrid frameworks based on association rule mining (ARM) and ANN to address the “large $p$, small nn” challenge.
+More recently, [12] and [38] introduced flexible machine-learning and group-association frameworks to jointly model correlated resistance phenotypes, while systematic evaluations of ML strategies on mutation-based inputs have highlighted the utility of standardized feature sets [31].
+Despite their utility, these models often rely on expert-defined feature engineering and treat sequencing quality metrics as binary filters rather than informative features.
+2.2 Deep Learning and Monolithic Architectures
+Deep learning (DL) has significantly advanced post-GWAS prioritization by capturing non-linear interactions through architectures like DeepSEA and ExPecto [53, 52].
+Convolutional Neural Networks (CNNs) have become a standard for predicting phenotypic traits from SNP data due to their ability to model local patterns [26, 35].
+Specifically for MTB, models like MD-CNN and CNN-GWAS have successfully integrated CNNs with genomic analysis to address missing heritability [16, 24].
+Other frameworks have utilized monolithic deep structures, such as DeepAMR [51], which employed denoising autoencoders to learn latent representations of co-occurrent resistance, and MD-WDNN [7], which pioneered a wide-and-deep approach to share information across multiple drugs.
+Furthermore, GenTB [17] and TB-DROP [46] provided fast, clinically-applicable resistance calls using multi-task and MLP-based representations.
+While successful, these models often rely on fixed-length mutation vectors that struggle with the “long-tail” of rare variants and show limitations to distinguish between biological signals and technical sequencing noise.
+2.3 Transformer-based and Graph Modeling
+The emergence of Transformer architectures, such as DNABERT [22], has enabled the processing of large DNA sequences through self-attention mechanisms.
+Lightweight models like LOGO combine self-attention and convolutional layers to interpret non-coding regions with high resolution [49].
+Modern sequence models like LLMTB [45] leverage BERT-based tokens to capture complex genomic patterns.
+This shift toward attention-driven models is reflected across clinical diagnostics, from cardiovascular disease detection [30] to oncology-driven feature extraction [40].
+Additionally, graph-based approaches have been introduced to model relational dependencies; for example, HGAT [50] employs heterogeneous graph attention networks to model the topological relationships between isolates and mutations, while [6] utilized graph representation learning for drug repurposing.
+2.4 The VAMP-Net Position: Expert-Informed Decoupling
+Despite the sophistication of current state-of-the-art (SOTA) models, a significant gap remains: existing architectures typically treat genomic data as a uniform black box, limited to isolate biological causality from technical artifacts.
+As illustrated in Table 1, VAMP-Net addresses this critical limitation through an expert-informed multi-path architecture.
+Unlike the sequence-dependent logic of LLMTB or the fixed-graph constraints of HGAT, VAMP-Net utilizes a permutation-invariant Set Attention Transformer (Path-1) to capture non-linear epistatic interactions as unordered biological sets.
+Crucially, it introduces a dedicated Technical Auditor (Path-2) using a 1D-CNN to process quality-aware metrics, a feature absent in the monolithic designs of DeepAMR or DeepPTB.
+By integrating these streams via an Adaptive Fusion mechanism, VAMP-Net mirrors the clinical audit-trail logic used by microbiologists, providing a level of diagnostic reliability and auditable interpretability that current models struggle to achieve.
+Table 1: Comparison of architectural properties and training datasets across prior genomic modeling approaches versus VAMP-Net.
+Method Model Type Variant Repr. Quality Handling Interpretability Dataset Used
+WHO/GWAS Statistical Categorical SNPs Hard Thresholds High (Manual) CRyPTIC ($N\approx 10k$)
+GenNet CNN Hierarchical Image Binary Filter Feature Map Public Repos ($N\approx 20k$)
+DeepAMR Autoencoder Binary Matrix None Latent Space Global Cohort ($N\approx 16k$)
+MD-WDNN Wide & Deep Fixed-Length Vector Static Input Weights PATRIC/ReSeqTB ($N\approx 3.6k$)
+HGAT GNN Heterogeneous Graph Implicit Edge Importance Kaggle/AMK ($N\approx 4.4k$)
+DNABERT Transformer k-mer Sequence N/A Self-Attention Ref. Genomes
+LLMTB Transformer BERT Tokens Sequence-only Attention Maps CRyPTIC ($N\approx 12k$)
+TB-DROP MLP/CNN Mutation Matrix Fixed Filters Low Curated Repos ($N\approx 4k$)
+VAMP-Net Multi-Path Unordered Set Adaptive Fusion Gated Logic CRyPTIC ($N\approx 10k$)
+3 Methodology
+Our approach introduces a Variant-Aware Multi-Path Network (VAMP-Net) that synergistically couples a Set Transformer architecture with a quality-aware 1D-Convolutional Neural Network for genetic variant analysis.
+This network is designed as a two-branch fusion model presented in Figure 1.
+3.1 VAMP-Net Overview:
+A).Path-1: The first branch, a Set Transformer, is specifically engineered to process the genetic variant sequence as an unordered set of tokens.
+This branch employs the Set Attention Block (SAB) [25] to capture the complex non-sequential relationships between the variants, reflecting the inherent permutation-invariant nature of a set.
+B).
+Path-2: The second branch of our framework is a quality-aware 1D-CNN.
+This path focuses on local sequential features.
+For each variant, it takes as input a corresponding quality score sequence.
+The 1D-CNN is efficient in capturing the pattern in these local quality signals, thus offering the model critical information on the confidence and reliability of a variant call.
+Figure 1: Overview of the Variant-Aware Multi-Path Network (VAMP-Net) Architecture.
+The framework employs a dual-pathway design to disentangle biological signal from technical noise.
+Path 1 (Biological Topology) utilizes a Set Attention Transformer (SAB) to process genomic variant sequences as unordered sets; the multi-head attention mechanism captures high-order epistatic interactions between mutations while maintaining permutation invariance.
+Simultaneously, Path 2 (Technical Confidence) employs a Quality-Aware 1D-CNN to process multi-channel VCF features, extracting local patterns of sequencing reliability.
+The outputs of both paths are integrated at a Late Fusion Node via an adaptive gating mechanism, allowing technical confidence to modulate biological findings.
+The fused representation is finally processed by a Classification Head consisting of fully connected layers and dropout regularization to produce the final drug resistance prediction.
+C).
+Fusion Module: Finally, the outputs of the Set Transformer and the 1D-CNN branches are fused through a late-fusion mechanism, as pesented in the Figure 1.
+The representations from both branches are concatenated and passed to a final classification layer.
+This fusion allows the model to make predictions based on a rich, multi-faceted understanding of the data: the global, set-based relationships between variants from the SAB, and the local, sequential quality signals from the 1D-CNN.
+This multi-scale learning (local + global) provides both robustness (resistance to noisy data) and biological interpretability (attention highlighting key variants).
+To formulate the problem addressed, a genomic sample is represented as a set of $N$ genetic variants: $V=\{v_{1},v_{2},...,v_{N}\}$ where each variant $v_{i}$ is described by $(\textbf{g}_{i},\textbf{f}_{i})$, with $\textbf{g}_{i}$ representing a sequence of genetic alteration and $\textbf{f}_{i}\in\mathbb{R}^{d}$ denoting the associated feature vector extracted from Variant Call Format (VCF) files.
+The prediction task is binary classification: $y=f_{\theta}(V)$, where $y\in\{0,1\}$ indicates resistance or susceptibility to a given drug, and $f_{\theta}$ is a parameterized deep learning model.
+The VAMP-Net approach is presented in Algorithm 1.
+Algorithm 1 VAMP-Net Multi-Path Inference and Adaptive Fusion, performed on an NVIDIA H100 GPU.
+1:Set of Genomic Variants $\mathcal{V}=\{v_{1},v_{2},\dots,v_{n}\}$, VCF Quality Metrics $\mathbf{Q}\in\mathbb{R}^{n\times d}$
+2:Resistance Probability $\hat{y}\in[0,1]$
+3:
+4:// Path 1: Biological Signal Extraction
+[…]
+[…]
+7:
+8:// Path 2: Technical Context Assessment
+[…]
+10:
+11:// Adaptive Fusion and Gating
+[…]
+13:if Mechanism is Suppression then
+14:  $\mathbf{Z}_{fused}\leftarrow\mathbf{Z}_{bio}\odot\mathbf{g}$ $\triangleright$ Down-weight low-confidence signals
+15:else if Mechanism is Amplification then
+16:  $\mathbf{Z}_{fused}\leftarrow\mathbf{Z}_{bio}+(\mathbf{Z}_{bio}\odot\mathbf{g})$ $\triangleright$ Enhance high-confidence signals
+17:end if
+18:
+19:// Final Prediction
+[…]
+[…]
+22:return $\hat{y}$
+3.2 Expert-Informed Design: The Rationale for Functional Decoupling
+The choice of these two specific architectures is a mathematical and biological necessity based on the different "shapes" of the genomic data.
+Genomic data in a clinical setting is heterogeneous.
+It contains two entirely different types of information that should not be processed the same way.
+1. Why Permutation-Invariant Set Attention? (The Biology):
+  • The Nature of Mutations: Genomic variants in an isolate function as a biological set rather than a linear sequence.
+Whether a mutation in katG appears before or after a mutation in rpoB in the data file is irrelevant to the bacteria’s resistance.
+  • Inductive Bias: Standard Transformers or RNNs assume a sequence (like a sentence).
+By using Set Attention (SAB) and removing positional encodings, we enforce a "set" inductive bias.
+This tells the model: "Focus on which mutations are present together (epistasis), not the order they were recorded."
+  • Capturing Epistasis: Attention mechanisms are mathematically superior at capturing non-linear interactions between distant variants (e.g., how a compensatory mutation in ahpC interacts with a primary mutation in katG).
+2. Why 1D-CNN? (The Technical Auditor:
+  • The Nature of Sequencing Quality: VCF features (Depth, Fraction of Supporting Reads, Confidence scores) are local, quantitative signals.
+  • Pattern Extraction: 1D-CNNs are the gold standard for extracting features from local neighborhoods of multi-channel data.
+They are designed to detect "shapes" in the noise, such as a specific drop in read depth across a variant, that signify a low-quality call.
+  • Feature Differentiation: If we fed these technical metrics into the Transformer along with the variants, the high-magnitude "quality" numbers might overwhelm the sparse "biological" signals.
+The 1D-CNN acts as a pre-processor that compresses raw technical noise into a single "Confidence Vector."
+3. Biomimetic Rationale: The Clinical Audit Workflow:
+Beyond mathematical necessity, the dual-pathway architecture of VAMP-Net is explicitly designed to mirror the diagnostic reasoning of expert clinical microbiologists.
+In a manual diagnostic setting, an expert does not evaluate a mutation in isolation; instead, they engage in a "multi-path" heuristic:
+  • Biological Assessment (Path-1): The expert identifies the presence of specific genetic markers (e.g., katG or rpoB mutations) and evaluates their known association with drug resistance.
+  • Technical Auditing (Path-2): Simultaneously, the expert scrutinizes the raw sequencing run’s quality, verifying that the read depth and genotype confidence are sufficient to support a clinical call.
+By formalizing these distinct tasks into specialized architectural branches, VAMP-Net avoids the "black-box" pitfalls of monolithic models.
+This functional decoupling ensures that the final prediction is not merely a statistical correlation, but a clinical decision where the biological signal is rigorously audited by technical reliability, a process we term Expert-Informed Fusion.
+3.3 Dataset for MTB Drug Resistance:
+The dataset used in this study comprises genetic variant sequences of Mycobacterium tuberculosis (MTB).
+Each sample is represented by a set of all identified genetic variants, resulting in sequences of variable length.
+3.3.1 Data Description
+Data for this study were obtained from Comprehensive Resistance Prediction for Tuberculosis: an International Consortium (CRyPTIC) [9].
 This globally important resource consists of a robust set of 12,289 clinical isolates of Mycobacterium tuberculosis, collected from 23 countries.
@@
 The complete dataset is publicly available at ftp.ebi.ac.uk/pub/databases/cryptic/release_june2022/.
-4 Methodology
-Our approach introduces a novel Variant-Aware Multi-Path Network (VAMP-Net) that synergistically couples a Set Transformer architecture with a quality-aware 1D-Convolutional Neural Network for genetic variant analysis.
-This framework is designed as a two-branch fusion model (Figure 1).
-The first branch, a Set Transformer, is specifically engineered to process the genetic variant sequence as an unordered set of tokens.
-This branch employs the Set Attention Block (SAB) to capture the complex non-sequential relationships between the variants, reflecting the inherent permutation-invariant nature of a set.
-The second branch of our framework is a quality-aware 1D-CNN.
-This path focuses on local sequential features.
-For each variant, it takes as input a corresponding quality score sequence.
-The 1D-CNN is efficient in capturing the pattern in these local quality signals, thus offering the model critical information on the confidence and reliability of a variant call.
-Figure 1: Overview of the Variant-Aware Multi-Path Network (VAMP-Net) Architecture
-Finally, the outputs of the Set Transformer and the 1D-CNN branches are fused through a late-fusion mechanism, as pesented in the Figure 1.
-The representations from both branches are concatenated and passed to a final classification layer.
-This fusion allows the model to make predictions based on a rich, multi-faceted understanding of the data: the global, set-based relationships between variants from the SAB, and the local, sequential quality signals from the 1D-CNN.
-This multi-scale learning (local + global) provides both robustness (resistance to noisy data) and biological interpretability (attention highlighting key variants).
-To formulate the problem addressed, a genomic sample is represented as a set of $N$ genetic variants: $V=\{v_{1},v_{2},...,v_{N}\}$ where each variant $v_{i}$ is described by $(\textbf{g}_{i},\textbf{f}_{i})$, with $\textbf{g}_{i}$ representing a sequence of genetic alteration and $\textbf{f}_{i}\in\mathbb{R}^{d}$ denoting the associated feature vector extracted from Variant Call Format (VCF) files.
-The prediction task is binary classification: $y=f_{\theta}(V)$, where $y\in\{0,1\}$ indicates resistance or susceptibility to a given drug, and $f_{\theta}$ is a parameterized deep learning model.
-4.1 Data Representation: Genetic Variants as a Set of Tokens
-The dataset used in this study comprises genetic variant sequences of Mycobacterium tuberculosis (MTB), as described in Section 3.
-Each sample is represented by a set of all identified genetic variants, resulting in sequences of variable length.
+3.3.2 Data Representation: Genetic Variants as a Set of Tokens
 Transformer-based architectures have fundamentally transformed the processing of sequential data, exhibiting remarkable ability to capture long-range dependencies and contextual relationships.
-In the genomics domain, recent transformer models have achieved promising results; however, most existing approaches primarily rely on reference DNA sequences or K-mer tokenization as inputs (consens2025interpreting).
+In the genomics domain, recent transformer models have achieved promising results; however, most existing approaches primarily rely on reference DNA sequences or K-mer tokenization as inputs [8].
 While such representations effectively capture local sequence patterns, they often treat variants as abstract tokens embedded within a continuous sequence.
-This abstraction can obscure the explicit biological relationships and combinatorial interactions among individual genetic alterations—relationships that are essential for understanding complex traits such as drug resistance.
+This abstraction can obscure the explicit biological relationships and combinatorial interactions among individual genetic alterations, relationships that are essential for understanding complex traits such as drug resistance.
 To address this limitation, we propose a variant-centric representation that explicitly encodes both the molecular identity and genomic position of each variant as a single, biologically meaningful token.
@@
 This compact and canonical identifier serves as a semantically rich token, allowing each sample’s genomic profile $\textbf{g}_{i}$ to be expressed as a discrete set of variant tokens T: $\textbf{g}_{i}=\{a_{1},a_{2},...,a_{T}\}$.
-Such a representation enables the transformer to reason directly over meaningful biological entities—facilitating the discovery of high-order, non-linear relationships among variants that underlie phenotypic outcomes such as drug resistance.
-This representation is particularly well-suited for a permutation-invariant modeling framework.
-By adopting this symbolic encoding, the entire collection of genetic variants for each sample can be treated as a set rather than an ordered sequence.
-This design allows the model to learn complex, non-linear dependencies and interactions among variants independent of their order in the dataset or genome file.
-The proposed tokenization strategy therefore serves as a key enabler of our set-based learning paradigm, marking a distinct shift from conventional sequence-oriented approaches and focus on the intrinsic relationships between variants themselves.
+Such a representation enables the transformer to reason directly over meaningful biological entities, facilitating the discovery of high-order, non-linear relationships among variants that underlie phenotypic outcomes such as drug resistance.
+To mitigate out-of-vocabulary (OOV) effects, we employ a data-driven tokenization scheme in which a BERT WordPiece tokenizer is trained on the complete set of unique variant strings extracted from the dataset As illustrated in Figure 2.
+The tokenizer operates without lowercasing and learns a compact vocabulary composed of genomic primitives and frequently occurring subpatterns, including numeric positions, canonical nucleotides (A, C, G, T), and structural delimiters (e.g., >> and ::).
+At inference time, each variant string is decomposed into valid subword units using a greedy longest-match strategy, ensuring that even rare or previously unseen variant compositions within the same representational schema can be expressed through combinations of known tokens.
+This approach substantially reduces reliance on the $[UNK]$ token and preserves the compositional structure of genomic variants.
+The resulting sub-token embeddings are subsequently aggregated to form a unified high-dimensional representation of each variant, which is then provided as input to the downstream Path-1 Set Attention Transformer.
+By aligning the tokenizer with the generative structure of the data, this method enables robust modeling of the long-tail distribution of genomic variants while maintaining vocabulary efficiency.
+Figure 2: WordPiece-based tokenization strategy for genomic variant encoding.
+This figure details the mechanism used to mitigate Out-of-Vocabulary (OOV) errors during the processing of non-canonical mutations.
+An input variant (e.g., $761155:C>T$) is processed via WordPiece-based tokenization, which decomposes the symbolic string into fundamental sub-units: numerical subunits, nucleotides, and delimiters.
+By leveraging Fixed Vocabulary Tables that contain these universal genomic components, VAMP-Net can construct unique embedding vectors for previously unseen rare coordinates.
+This approach avoids the information loss associated with traditional $[UNK]$ (unknown) tokens.
+The resulting Sub-token Embeddings are aggregated to produce a single, high-dimensional Variant Embedding Vector ($V_{761155:C>T}$), which is then passed to the Path-1 Set Attention Transformer.
+This method ensures that the model can learn from the "long tail" of genomic data while maintaining a compact and efficient vocabulary.
 3.4 Model Architecture: Variant-Aware Multi-Path Network (VAMP-Net)
-Variant-Aware Multi-Path Network (VAMP-Net), a dual-path deep learning framework specifically designed to address the limitations of conventional single-stream models in genomic data analysis.
+Variant-Aware Multi-Path Network (VAMP-Net), a dual-path deep learning network specifically designed to address the limitations of conventional single-stream models in genomic data analysis.
 VAMP-Net is architected to process two distinct sources of information essential for robust variant-based prediction ( Figure 1): (1) the symbolic identity and biological context of each genetic variant and (2) the quality metrics of quantitative sequencing that reflect data reliability.
 The rationale for this dual-path design is twofold: first, to leverage the inherent set-based nature of a sample’s variant, where order is non-critical, by using a modern attention mechanism; and second, to capture local patterns and dependencies within the associated quality signals using convolutional processing.
-The framework is structured as follows:
+The model is structured as follows:
   • Path-1 – Symbolic and Topological Stream: Processes the set of genomic variants using a symbolic tokenization scheme at variant-level, allowing the model to learn relationships regardless of input order.
@@
 This enables the model to learn a unified representation of the set, a property that is essential for our application.
-Figure 2: Path-1 Architecture: Set Attention Transformer Block (SAB) for Symbolic Variant Processing
+Figure 3: Path-1 Architecture: Set Attention Transformer Block (SAB) for Symbolic Variant Processing.
+This pathway facilitates the extraction of high-order epistatic features from raw genomic data.
+Raw variant text inputs (formatted as Position_Ref>AltPosition\_Ref>Alt) are first mapped into a continuous vector space via a Compositional Embedding layer.
+These embeddings are processed through a series of NN stacked Set Attention Blocks (SAB).
+Unlike standard Transformers, the SAB utilizes self-attention mechanisms without positional encodings, ensuring the model remains permutation-invariant.
+Within each SAB, Multi-Head Attention (MHA) computes Query (Q), Key (K), and Value (V) interactions to model the non-linear dependencies between distant mutations.
+The final output is a set of per-variant feature representations that encode the biological topology of the isolate, specifically capturing the synergistic effects of multiple mutations on the drug resistance phenotype.
 The SAB pathway models variant-variant interactions while maintaining permutation invariance.
@@
 I - Unmasked Set Attention:
-The first contribution adapts the Set Transformer framework (lee2019set) of the unmasked SAB for modeling genetic variant sequences represented as unordered tokens.
+The first contribution adapts the Set Transformer framework [25] of the unmasked SAB for modeling genetic variant sequences represented as unordered tokens.
 Unlike standard transformers for sequential data, no attention masks —neither causal masks is applied, ensuring each variant can attend to every other variant without positional or causal constraints, nor padding masks is applied.
@@
 The training objective is to perform a specific task of Drug Resistance, based on their set of variants.
-The hyperparameters for this model, including the number of SAB layers, hidden dimensions, and learning rate, are provided in the Experimental Section 5.4 .
+The hyperparameters for this model, including the number of SAB layers, hidden dimensions, and learning rate, are provided in the Experimental Section 4.2.3.
 II- Padding-Masked Set Attention :
@@
 The 1D-CNN branch processes the variant feature matrix $\textbf{F}\in\mathbb{R}^{N\times d}$.
-Figure 3: Path-2 Architecture: Quality-Aware 1D Convolutional Neural Network (1D-CNN) for VCF Feature Processing
+Figure 4: Path-2 Architecture: Quality-Aware 1D Convolutional Neural Network (1D-CNN) for VCF Feature Processing.
+This pathway is dedicated to the quantitative assessment of sequencing reliability.
+The input consists of a multi-channel tensor of VCF-derived features, including Genotype Calls (GT), Read Depth (DP), and Fraction of Supporting Reads (FRS), among others listed in the Data Callout Board.
+These features are processed through a series of NN stacked Convolutional Blocks, each containing a 1D-convolutional layer for local pattern extraction.
+Unlike traditional binary filters, this 1D-CNN learns a continuous, high-dimensional representation of technical confidence for each variant.
+These per-variant feature representations are then utilized by the fusion module to adaptively weight the biological signals from Path-1, ensuring that resistance predictions are grounded in high-quality sequencing data.
 While the core genetic information (CHROM, POS, REF, ALT) is fundamental, the true richness of VCF files for deep learning lies within their FORMAT fields.
@@
 3.4.3 Multi-Path Fusion Mechanism
-The final stage of the VAMP-Net architecture involves the integration of the learned representations from the two specialized paths.
-The outputs of the two branches—the SAB’s permutation-invariant set representation ($\mathbf{z}_{\text{SAB}}$) from Path-1, and the 1D-CNN’s quality-aware feature vector ($\mathbf{z}_{\text{CNN}}$) from Path-2 —are concatenated to form a comprehensive, fused representation $\mathbf{z}_{\text{fused}}$ (Equation 8):
-$\mathbf{z}_{\text{fused}}=[\mathbf{z}_{\text{SAB}};\mathbf{z}_{\text{CNN}}]$ (8)
-This late-fusion approach is crucial, as it allows each branch to specialize independently in its unique data modality (symbolic variant topology vs.
-quantitative feature sequence) before their insights are combined.
-The fused vector is then passed through a fully connected network to a final output layer for the prediction of drug resistance.
-To determine the optimal strategy for combining these disparate signals, we investigate three distinct modulation mechanisms between the output of Path-1 (transformer) and Path-2 (1D-CNN):
+The final stage of the VAMP-Net architecture involves the formal integration of the learned representations from the two specialized paths: the biological topology manifold and the technical quality manifold.
+- Formal Integration Framework:
+Let $\mathbf{z}_{\text{SAB}}\in\mathbb{R}^{d}$ be the permutation-invariant set representation from Path-1, and $\mathbf{z}_{\text{CNN}}\in\mathbb{R}^{k}$ be the quality-aware feature vector from Path-2.
+To ensure that technical confidence can modulate biological signal, we define a Gated Fusion Module that maps technical noise patterns to a modulation intensity vector $\mathbf{g}\in\mathbb{R}^{d}$ via a sigmoid-activated transformation:
+$\mathbf{g}=\sigma(\mathbf{W}g.\mathbf{z}_{\text{CNN}}+\mathbf{b}_{g})$ (8)
+where $\mathbf{W}_{g}$ and $\mathbf{b}_{g}$ are learnable parameters.
+This gating vector acts as a "Technical Auditor," determining the trustworthiness of each dimension in the biological feature space.
+- Fusion Modulation Strategies:
+To determine the optimal strategy for combining these disparate signals, we investigate three distinct modulation mechanisms that resolve conflicts between biological data and technical noise:
   1. 1.
-    Suppression (De-emphasis of Quality):
-    The first fusion mechanism is inspired by LSTM’s forget gate control.
-The purpose of this gate in LSTM is to control how much of the old state is remembered in the new state.
-By taking the sigmoid of the current input $C_{t}$ and the hidden state $H_{t-1}$, values are transformed into the probability domain, and then multiplied element-wise with $C_{t-1}$, the past internal state.
-    We modify the purpose and formula of this gate to suit our case: by taking the sigmoid of Path-2, we use the additional features to control the confidence of the current variants, and then multiply the result by the output of Path-1.
-The Suppression fusion is defined in Equation 9, where $\odot$ denotes element-wise multiplication.
-    $\mathbf{g}=\sigma(\mathbf{z}_{\text{CNN}})$
+    Suppression (De-emphasis of Noise):
+    Inspired by the "forget gate" in Long Short-Term Memory (LSTM) networks, this mechanism performs element-wise multiplication to scale the biological signal:
     $\mathbf{z}_{\text{fused}}=\mathbf{g}\odot\mathbf{z}_{\text{SAB}}$ (9)
-    This formulation applies a multiplicative suppression to the output of Path-1 ($\mathbf{z}_{\text{SAB}}$).
-Since the modulation score $\mathbf{g}$ is constrained to $[0,1]$, the resulting fused vector $\mathbf{z}_{\text{fused}}$ will have element values ranging between 0 and the corresponding values in $\mathbf{z}_{\text{SAB}}$.
-In effect, Path-2’s quality features provide a scalar factor that selectively scales the influence of each variant, and if the quality metrics indicate low confidence, the variant’s representation can be driven toward 0, effectively performing a dynamic, data-driven form of variant feature selection during the fusion process.
+    As $\mathbf{g}\in[0,1]$, this formulation performs a dynamic, data-driven feature selection.
+If Path-2 identifies poor sequencing quality (e.g., low depth), $\mathbf{g}\to 0$, effectively suppressing the influence of that variant to prevent the propagation of sequencing artifacts.
   2. 2.
-    Amplification (Emphasis on Quality):
-    The second investigated fusion strategy, Amplification, is designed to explicitly test the hypothesis that quality and low-level features (Path-2) provide a critical signal gain to the core symbolic variant representation (Path-1).
-This mechanism ensures that high-quality, high-confidence variants can have proportionally increase in impact, thereby emphasizing a robust subset of the data.
-    Unlike the Suppression mechanism, this formulation guarantees that the symbolic representation $\mathbf{z}_{\text{SAB}}$ is always passed through to the final output, regardless of the quality metrics.
-The Quality-Aware vector $\mathbf{z}_{\text{CNN}}$ acts as an additive gain control.
-We adapt the gating mechanism as defined in Equation 10, where $\odot$ denotes element-wise multiplication.
-    $\mathbf{g}=\sigma(\mathbf{z}_{\text{CNN}})$
+    Amplification (Emphasis on Signal): To test if high-quality data provides a critical signal gain, we introduce a baseline-preserving mechanism:
     $\mathbf{z}_{\text{fused}}=(1+\mathbf{g})\odot\mathbf{z}_{\text{SAB}}$ (10)
-    This formulation achieves a soft integration that preserves the full contribution of the core variant signal.
-By introducing the constant $1$ into the scalar factor $(1+\mathbf{g})$, the mechanism guarantees a baseline preservation, ensuring that the fused output $\mathbf{z}_{\text{fused}}$ is never less than the original symbolic representation $\mathbf{z}_{\text{SAB}}$.
-As the modulation score $\mathbf{g}$ is constrained to the range $[0,1]$, the total scalar factor $(1+\mathbf{g})$ remains bounded within the interval $[1,2]$.
-This results in a controlled form of additive gain, where the most robustly supported variants—those with high-quality metrics—receive a maximum scaling of $200\%$ of their original magnitude.
-This approach avoids the risk of zeroing out potentially relevant variants, allowing all identified mutations to participate in the final prediction while prioritizing those with the highest associated confidence.
+    This ensures that the symbolic representation $\mathbf{z}_{\text{SAB}}$ is always preserved, while high-confidence variants receive a gain of up to $200\%$.
+This prioritizes robustly supported mutations without risking the "zeroing out" of potentially relevant signals.
   3. 3.
-    Adaptive Amplification & Suppression
-    The final fusion strategy is Adaptive Amplification & Suppression, designed to provide the model with a dynamic, context-aware control over the Path-1 signal.
-This mechanism combines the feature selection capability of Suppression with the signal enhancement capability of Amplification, allowing the influence of the symbolic variant representation ($\mathbf{z}_{\text{SAB}}$) to be either scaled down or scaled up based on the quality metrics ($\mathbf{z}_{\text{CNN}}$).
-    We synthesize the previous two modulation strategies by modifying the gating factor.
-Using the same modulation score $\mathbf{g}=\sigma(\mathbf{z}_{\text{CNN}})$, the fused vector is calculated as Equation 11.
-    $\mathbf{z}_{\text{fused}}=(2\odot\mathbf{g}-1)\odot\mathbf{z}_{\text{SAB}}$ (11)
-    where $\odot$ denotes element-wise multiplication.
-    This adaptive mechanism creates a bipolar scaling factor which introduces a powerful, new operational dynamic.
-As the modulation score $\mathbf{g}$, derived from the quality features, is constrained to the range $[0,1]$, the resulting scaling factor $(2\odot\mathbf{g}-1)$ is dynamically mapped to the closed interval [1,1][-1,1].
-This allows the model to continuously transition between Suppression Zone (noise filtering) and Amplification Zone (signal gain) based on the input quality.
-Specifically, when quality features are poor ($\mathbf{g}\approx 0$), the scaling factor approaches 1-1, resulting in a negative scaling of the variant representation, effectively acting as an aggressive noise filter by reversing the polarity of low-confidence signal.
-Conversely, when quality features are excellent ($\mathbf{g}\approx 1$), the factor approaches $1$, leading to maximum positive scaling and thus prioritizing the most robustly supported variants.
-This ability to span both positive and negative scalar values allows the model to not only ignore noise but to actively penalize uncertain variants, providing a comprehensive, context-aware control signal to the final prediction layers .
+    Adaptive Amplification and Suppression (Proposed Best):
+    Our final strategy synthesizes both approaches into a bipolar scaling factor that provides the model with dynamic, context-aware control:
+    $\mathbf{z}_{\text{fused}}=(2\mathbf{g}-1)\odot\mathbf{z}_{\text{SAB}}$ (11)
+    By mapping the scaling factor to the interval [1,1][-1,1], the model can continuously transition between a Suppression Zone (noise filtering) and an Amplification Zone (signal gain).
+This allows the model to not only ignore noise but to actively penalize uncertain variants by reversing their signal polarity, mirroring the rigorous audit-trail logic used by clinical microbiologists.
+Final Classification
+The resulting vector $\mathbf{z}_{\text{fused}}$ is passed through a global average pooling layer and a Softmax-activated dense layer to produce the final resistance probability $P(y=1)$:
+$P(y=1)=\text{Softmax}(\mathbf{W}f.\mathbf{z}_{\text{fused}}+\mathbf{b}_{f})$ (12)
+- Functional Interpretation of Adaptive Fusion:
+The bipolar scaling mechanism in VAMP-Net is designed to resolve conflicts between biological signals and technical noise. This mathematical operation corresponds to two distinct operational modes:
+  • Adaptive Suppression: If Path-2 detects poor sequencing quality (e.g., low depth or high genotype noise), the gating factor $\mathbf{g}$ approaches 0, and the scaling factor becomes negative.
+This effectively filters the variant out, preventing the model from assigning resistance based on a potential sequencing artifact.
+  • Adaptive Amplification: Conversely, when Path-2 confirms high technical confidence for a non-canonical mutation, the scaling factor approaches $1$.
+This amplifies the signal, allowing the model to prioritize novel, high-confidence epistatic interactions that might be missed by static filters.
+This design ensures that biological causality is always contextualized by data reliability, mirroring the audit-trail logic used by clinical microbiologists during manual isolate review.
 3.5 Interpretability
@@
 This section presents the empirical validation of the Variant-Aware Multi-Path Network (VAMP-Net) and its core components, followed by a detailed discussion of the biological insights derived from our interpretability framework.
-We first detail the preparation of the dataset and the specific encoding utilized for the ChromPos_Ref>Alt variant tokens.
+We first detail the preparation of the dataset and the specific encoding utilized for the ChromPos_Ref>>Alt variant tokens.
 Subsequent subsections systematically evaluate three critical aspects of the design of VAMP-Net: the selection of the optimal fusion mechanism, a comprehensive performance comparison, and a deep-dive interpretation of the predictions of the final model, including variant importance and epistatic interaction networks.
@@
 Prior to use, the data is subjected to multiple preparation procedures.
-After downloading the CRyPTIC_reuse_table_20240917.csv file from the official compendium website, we chose to focus on two drugs: Rifampicin (RIF), which is primarily used to treat tuberculosis by inhibiting RNA polymerase, and Rifabutin (RFB), a related rifamycin derivative with similar therapeutic applications.
+After downloading the CRyPTIC_reuse_table_20240917.csv file from the official compendium website, we focus on four drugs: Rifampicin (RIF), which is primarily used to treat tuberculosis by inhibiting RNA polymerase; Rifabutin (RFB), a related rifamycin derivative with similar therapeutic applications; Isoniazid (INH), a prodrug that targets the synthesis of mycolic acids, essential components of the mycobacterial cell wall; and Ethambutol (EMB), which acts by inhibiting the enzyme arabinosyltransferase, thereby disrupting the assembly of the mycobacterial cell wal.
 First, we filter the dataset to include only high and medium quality samples, based on the {drug}_PHENOTYPE_QUALITY field.
-This results in 10,310 samples for RIF and 11,392 samples for RFB.
 For each selected sample, we download its corresponding VCF file, which contains variant information.
@@
 Variants are represented as pos_ref>>alt for all alleles separated by commas; we refer to this representation as the Path-1 Input, after this filtering, we end up with a total of 417,635 unique variants.
-We also retain additional descriptive features for each variant: [’GT’, ’DP’, ’DPF’, ’COV’, ’FRS’, ’GT_CONF’, ’GT_CONF_PERCENTILE’].
+We also retain additional descriptive features for each variant: GT, DP, DPF, COV, FRS, GT_CONF, and GT_CONF_PERCENTILE.
 Numeric features are scaled using min–max normalization, while categorical features are encoded.
 The FRS field, representing the fraction of reads supporting the genotype call (#reads supporting GT / total reads), ranges between 0 and 1; higher FRS values indicate stronger confidence in the genotype, whereas missing or low values typically correspond to low coverage or filtered calls, and hence we chose to replace its missing values with zeros.
-For each sample, these features are organised into an array of shape (number of variants × 8 features × 1), referred to as the Path-2 Input, which will be used later in a 1D-CNN model.
+For each sample, these features are organised into an array of shape (number of variants × 8 features × 1), referred to as the Path-2 Input, which will be used in a 1D-CNN model.
 Arrays are padded as needed to ensure uniform length across samples.
@@
 This positional inconsistency across samples motivates the use of a Permutation-Invariant Block (SAB) in the model architecture.
-Figure 4: Distribution of classes for the selected drugs.
-Figure 4 shows the class distribution of the two drugs, with a difference of $\sim$17.
-6% between the susceptible and resistant class for RIF, and $\sim$30.
-4% for RFB, requiring us to employ a weighted cross-entropy loss to handle this class imbalance.
-We will first use the RIF data in all subsequent experiments, navigating several design choices to determine the optimal approach, and subsequently apply the selected method to RFB for validation.
-5.2 Variants Encoding
-As mentioned in Section 5.1, variants are represented as Pos_REF>ALT, with the ALT for each allele separated by a comma.
-We compare a Static Encoding method , where each variant is given a unique ID to represent it, against training a BERT Tokenizer on all unique variants , and then using it to encode each sample’s variants into vectors before training.
-The results showed that Static Encoding led to overfitting in the long run, with early stopping triggered at epoch 3.
-The training curve clearly showed signs of overfitting, with a $\sim$6.73% gap between training and validation accuracy.In contrast, using BERT proved effective in delivering more stable learning and improved validation performance.
-Although it also showed signs of overfitting, with a 4.83% difference between training and validation accuracy at the end of training, it maintained a smoother learning curve and reached early stopping at epoch 16.
-The BERT-based encoding was more robust and critical for stable learning as shown in Figure 5. Consequently, we selected BERT for the encoding phase in all subsequent experiments.
-Figure 5: Comparison of Static Encoding vs BERT Tokeniser metrics over epochs
-5.3 The Choice of Fusion Mechanism
-We tested three types of fusion strategies between Path-1 and Path-2, namely: Suppression, Amplification, and Suppression-Amplification (see Section 4.2.3).
-The training and validation curves are shown in Figure 6 and the test results are summarizsed in Table 1.
-All three types achieved comparable performance in terms of validation and test accuracy.
-However, in general, the Amplification fusion yielded the best results and, therefore, it was selected as the preferred mechanism for the final configuration, and we will refer to this model as - Model A -.
-Table 1: Comparison of fusion mechanisms. Best results in bold
+Figure 5: Dataset Class Distribution across the selected anti-TB drugs.
+This figure illustrates the sample size ($n$) and the distribution between resistant (red) and susceptible (blue) classes for the four targeted drugs.
+Figure 5 illustrates the sample counts and class distributions for each drug.
+We observe a class imbalance of approximately 17.6% between susceptible and resistant samples for RIF and 30.4% for RFB, while EMB exhibits the most significant disparity at 60.4%.
+To mitigate the effects of these imbalances during training, a weighted cross-entropy loss is employed
+Initially, we utilize the RIF dataset to evaluate various architectural design choices and establish an optimal modeling pipeline.
+This finalized approach is subsequently applied to RFB, INH, and EMB to validate its generalizability across different drug resistance profiles.
+The train/validation/test split was (80/10/10), clarifying that this was performed at the isolate level to prevent data leakage.
+We also specify that stratification was applied based on the phenotypic resistance labels to ensure balanced representation across all sets.
+4.2 Experimental Results
+This section details a multi-stage evaluation of the VAMP-Net to quantify its predictive performance and justify its architectural design.
+We first establish a performance baseline by evaluating various feature encoding strategies and fusion strategies.
+A rigorous ablation study follows, isolating the specific contributions of the VCF-aware Path-2 branch to the model’s decision-making process.
+Finally, we report the performance of the best-performing VAMP-Net configuration across four primary anti-TB drugs (RIF, RFB, INH, and EMB), evaluating its robustness, calibration, and clinical utility on unseen isolates.
+4.2.1 Evaluation of Feature Encoding Strategies
+Following the preprocessing pipeline described in Section 4.1, genomic variants are represented as Pos_REFPos\_REF>ALTALT.
+We conducted a comparative evaluation between two distinct representational paradigms: (1) a Static Encoding approach, where each unique variant is mapped to a discrete identifier, and (2) a Contextualized Tokenization approach utilizing a BERT-based tokenizer trained on the corpus of unique variants to project genomic features into a continuous vector space.
+Empirical results indicate that the Static Encoding method suffered from premature convergence and poor generalization.
+Early stopping was triggered at epoch 3, with the learning curves exhibiting a significant $\sim$6.73% generalization gap between training and validation accuracy.
+Conversely, the BERT-based representation demonstrated superior stability and robust feature extraction.
+While some degree of overfitting remained present—noted by a 4.83% accuracy disparity—the model exhibited a more sustained learning trajectory, converging at epoch 16 with a smoother loss profile.
+As illustrated in Figure 6, the BERT-based encoding facilitated more reliable optimization; consequently, it was selected as the foundational encoding strategy for all subsequent experiments.
+Figure 6: Training and validation dynamics across tokenization strategies.
+Comparative analysis of BERT Tokenizer (top row) and Static Encoding (bottom row) across Accuracy, Loss, and ROC-AUC metrics.
+While both methods reach high training performance, the BERT-based approach demonstrates superior stability and generalization, as evidenced by the significantly narrower gap between training and validation curves in the Loss and ROC-AUC plots compared to the overfitting observed in Static Encoding.
+4.2.2 Evaluation of Multimodal Late Fusion Strategies
+To identify the optimal integration strategy between Path-1 and Path-2, we evaluated three distinct fusion paradigms: Suppression, Amplification, and a hybrid Suppression-Amplification mechanism.
+The optimization trajectories are illustrated in Figure 7, while a comprehensive performance summary across the test set is provided in Table 2.
+Table 2: Comparative performance metrics across Gated Adaptive Fusion mechanisms.
+The table evaluates three functional modes of the adaptive fusion head: Suppression, Amplification, and a combined Suppression-Amplification strategy.
 Fusion Type Accuracy Precision Recall F1 AUC
@@
 Suppression 0.9438 0.9460 0.9591 0.9525 0.9710
-Figure 6: Training (top) and validation (bottom) performance curves for fusion methods
-5.4 Models Experimentation
-Starting from - Model A - , we evaluated a set of configurations to measure the robustness of the padding/masking strategy, data augmentation, and to assess the benefits of including Path-2 in the architecture.
-Moreover, we compared with a traditional approach, adopted in many studies, which involved representing the data as a matrix of shape (Number of VVariants) $\times$ (Number of Samples), where cell $i,j$ is 1 if Variant-$i$ is present in the sample, otherwise 0.
-This representation is typically trained with classifiers using either classical machine learning or deep learning techniques such as CNNs or MLP-based models.
-This hybrid of simple binary encoding and deep models has been widely surveyed in antimicrobial resistance (AMR) and clinical genomics literature and has been effective in practice (e.g., CNN and ML studies for tuberculosis and other pathogens) (sakagianni2024data).
-However, one limitation of this approach is the large resulting sparse matrix due to the high number of variants, i.e., the curse of dimensionality, requiring extensive computational resources.
-Some studies addressed this by pre-eliminating variants using statistical methods, or dimensionality reduction techniques like PCA (reshetnikov2025feature; pikalyova2024predicting; xu2025machine).
-To select informative variants, we applied a Chi-squared test to assess the association between each variant and the target phenotype.
-Variants with $p$-value <0.001<0.001 were retained, resulting in 11,567 ($\sim$36.1% reduction), which were then used to train and evaluate both a multi-layer perceptron (MLP) and a convolutional neural network (CNN).
-To systematically evaluate the contributions of the proposed components, we developed and tested several distinct model architectures:
+Figure 7: Impact of Gated Fusion strategies on model convergence during training.
+The trajectories compare three fusion variants: Amplification (orange), Suppression (yellow), and the combined Suppression-Amplification mechanism (green).
+While all methods converge to high ROC-AUC values (>0.95>0.95), the Amplification strategy exhibits the most stable optimization profile over 16 epochs, whereas the combined Suppression-Amplification approach shows greater variance, suggesting a trade-off between strict noise filtering and signal preservation.
+While all three mechanisms demonstrated competitive performance with negligible variance in validation and testing metrics, the Amplification Fusion consistently yielded the highest predictive accuracy and exhibited superior generalization.
+Consequently, this strategy was adopted for the final model architecture, hereafter referred to as Model A.
+4.2.3 Model Validation and Ablation Study
+Building upon the selection of Model A as the architectural baseline, we conducted an extensive ablation study to evaluate the robustness of our padding/masking strategies, the efficacy of data augmentation, and the specific contributions of the dual-path integration.
+To systematically dissect the influence of each component, we evaluated six model configurations (Models A–F), as detailed below.
+The architectural variants were designed to isolate the effects of Attention Masking, Data Augmentation, and the Path-2 feature stream.
   1. 1.
@@
     Model-F: Model-B settings, without the inclusion of Path-2.
-  7. 7.
-    Model-G: MLP for Presence/absence Variants data.
-  8. 8.
-    Model-H: CNN for Presence/Absence Variants data.
-Table 2: Best hyperparameters of -Path-1 - for each proposed Multi-Path model applied on RIF Drug + Model A on RFB Drug
+Optimal hyperparameters were identified via Bayesian Optimization (Tables 3, 4), prioritizing a robust balance between the Area Under the Receiver Operating Characteristic (ROC AUC) curve and Balanced Accuracy on the validation set.
+This approach ensures that the selected configurations prioritize generalization over mere training set convergence.
+Table 3: Comprehensive hyperparameter configurations for VAMP-Net.
+The table summarizes the optimized structural for Models A through F applied on RIF drug and model A apllied on RFB, INH and EMB drugs.
+This mapping includes the specific configurations for Path-1 Transformer (SAB).
 Model Emb Dim Hidden Dim Num Layers
-Model A 64 32 3
+Model A (RIF) 64 32 3
 Model B 64 32 2
@@
 Model A (RFB) 128 32 1
-Table 3: Best hyperparameters of - Path-2 - and the Classification head for each proposed Multi-Path model applied on RIF Drug + Model A on RFB drug
+Model A (INH) 128 32 2
+Model A (EMB) 128 32 2
+Table 4: Comprehensive hyperparameter configurations for VAMP-Net.
+The table summarizes the optimized structural and learning parameters for Models A through F applied on RIF drug and model A apllied on RFB, INH and EMB drugs.
+This mapping includes the specific configurations for Path-2 (1D-CNN) and the Classification head.
 Model Dropout Activation Conv Layers Conv Kernel LR
-Model A 0.1128 relu 3 3 0.00137
+Model A (RIF) 0.1128 relu 3 3 0.00137
 Model B 0.0962 gelu 2 5 0.00165
@@
 Model A (RFB) 0.2704 relu 4 5 0.00106
-Table 4: Best hyperparameters of MLP model aplied on RIF Drug
-Model Hidden Dims Dropout Activation
-Model G [1024, 128, 128] 0.1315 gelu
-Table 5: Best hyperparameters of CNN model applied on RIF Drug
-Model Dropout Activation Conv Layers Filters Kernel Sizes Pool Sizes
-Model H 0.1487 relu 3 [96, 128, 256] [5, 9, 2] [3, 3, 3]
-The proposed models were evaluated across multiple architectures and hyperparameter configurations.
-Tables 2, 3, 4, and 5 summarize the best-performing hyperparameters identified through Bayesian optimization.
-This tuning prioritized a balance between ROC AUC and Accuracy on the validation set, ensuring that models not only fit the training data but also generalized robustly.
-Figure 7: Comparison of -Model A- and -Model B- performance without and with variant sequence shuffling for training and validation
-In Figure 7, we compare the performance of Model-A and Model-B where Model-B applies data augmentation via variant sequence shuffling.
-This augmentation fully leverages the SAB block’s capacity, allowing predictions independent of variant position.
-Training curves confirm that Model-B, likely due to the regularization effect of data augmentation, exhibits greater training stability as evidenced by significantly reduced overfitting and closer convergence between the training and validation curves.
+Model A (INH) 0.326 relu 1 3 0.00141
+Model A (EMB) 0.326 relu 1 3 0.00141
+The role of data augmentation, implemented via variant sequence shuffling, was scrutinized by comparing Model A and Model B.
+Shuffling leverages the Set Attention Block (SAB)’s inherent capacity for permutation invariance, ensuring that the model learns to identify critical genomic signatures independent of their relative position in the input sequence.
+As illustrated by the optimization trajectories in Figure 8, Model B exhibited superior training stability and a narrowed generalization gap, acting as an effective regularizer.
 On the test set, Model-A achieved the highest performance (ROC AUC: 0.969, Balanced Accuracy: 0.945).
@@
 This suggests that data augmentation effectively improves generalization and stability with only a negligible decrease in maximum predictive power.
-Figure 8: Training (top) and validation (bottom) performance curves of all models with early stopping
-Figure 9: ROC AUC Test curve of VAMP-Net vs CNN and MLP
-Figure 8 presents the training and validation curves across epochs for eight distinct models (Model-A through Model-H), offering critical insight into their convergence, learning capacity, and generalization ability.
-All models exhibit rapid initial learning, achieving 85% to over 90% accuracy within the first 5 to 10 epochs.
-The models can be broadly categorized into three performance tiers based on their final stabilized validation accuracy and AUC.
-Models A, B, C, and D demonstrate the highest potential, peaking around 94%.
-Models G and H show comparatively lower learning capacity for this task, suggesting architectural limitations and insufficient model capacity.
-Models G and H required the longest training durations, this is a characteristic of models with less efficiency.
-Model-A achieved the highest overall training and validation accuracy and AUC within the shortest training time, making it the most efficient high-performer.
-This demonstrates that Path-2 inclusion significantly improved robustness and that SAB-based transformers outperform classical CNN/MLP representations.
-The lower performance of Models G and H suggests that these architectures are less suitable for the current task, as shown in Figure 9.
-Figure 10: Models test results comparison
-The results, visualized in Figure 10, clearly demonstrate the superior performance of the Set Attention Block (SAB)-based transformer architectures.
-While Model-A achieved the highest overall performance, the marginal variation across other SAB variants underscores the robustness of the core attention mechanism.
-Notably, the impact of explicit masking appears minimal, as evidenced by the comparable performance of Models C and D, suggesting that the SAB structure inherently handles or ignores padding tokens effectively.
-In contrast, the significant underperformance of the traditional Multilayer Perceptron (MLP) (Model G) and Convolutional Neural Network (CNN) (Model H) strongly validates the utility of a sequential attention-based representation over standard binary encoding or fixed-filter approaches for this task.
-Based on its performance across all evaluation metrics, we selected Model-A as the optimal architecture.
-To validate the model’s capacity for broad applicability, we applied this model on the RFB drug.
-On this drug, Model-A demonstrated high efficacy, achieving an AUC of 0.9681 and an Accuracy of 0.9393.
-Furthermore, the model maintained a well-balanced classification profile, yielding an F1 score of 0.9535, a Recall of 0.9553, and a Precision of 0.9517.
-These results collectively confirm the model’s excellent generalization capacity across different therapeutic agents.
-The comprehensive architectural evaluation leads to the following four key conclusions:
-  • Superiority of Attention: Set Attention Block (SAB)-based transformer architectures significantly outperform standard Convolutional Neural Network (CNN) and Multilayer Perceptron (MLP) baselines on presence/absence data.
-  • Enhanced Generalization: Augmentation (as employed in Model-B) effectively improves training stability and promotes permutation invariance, leading to more robust models.
-  • Architectural Robustness: While empirically minor, the inclusion of attention masks provides critical robustness against padding tokens, formalizing the model’s ability to handle variable-length sequences.
-  • Broad Applicability: The high performance achieved on the RFB drug dataset validates the broad applicability and transferability of the SAB-based transformer approach to new, unseen drugs.
-5.5 Multi-Path Framework Analysis and Interpretation
-This section prensents the results of the Dual-Path Interpretability Framework and leverages the insights to dissect the model’s decision-making process for RIF and RFB resistance prediction.
+The results summarized in Table 5 provide critical insights into the contribution of each modular component within the VAMP-Net (VN) network.
+A comparative analysis yields several key findings, the full VAMP-Net (VN) configuration, which incorporates attention masking and localized feature extraction, achieves the highest overall performance across most metrics, specifically reaching an Accuracy of 0.956 and an F1-score of 0.960.
+This confirms that the synergistic integration of the dual-path architecture and precise masking provides the most robust predictive signal for Rifampicin resistance.
+Notably, the removal of the attention mask ("No mask, no shuffle") does not result in a catastrophic loss of performance, maintaining a high ROC-AUC of 0.973.
+This suggests that the SAB mechanism is inherently capable of identifying informative tokens and effectively disregarding padding, even without explicit masking.
+However, the slightly lower Balanced Accuracy (0.942) compared to the baseline (0.944) indicates that masking still provides a marginal benefit in stabilizing predictions for minority classes.
+The exclusion of Path-2 ("No Path-2") consistently resulted in a performance decrement, with Balanced Accuracy dropping to its lowest point (0.930).
+This validates the hypothesis that Path-B captures complementary genomic information, likely broader contextual or inter-variant relationships—that a single-path transformer architecture might overlook.
+Figure 8: Validation of permutation invariance via variant sequence shuffling.
+Comparative analysis of model performance under data augmentation (top row) and standard training (bottom row).
+The top panels demonstrate that variant shuffling does not degrade performance; rather, it acts as a regularizer that tightens the generalization gap.
+The stability of ROC-AUC and Accuracy across both rows, despite the randomized ordering of input mutations, validates the permutation-invariant inductive bias of the Set Attention Block (SAB), ensuring the model captures biological topology regardless of variant positioning.
+Table 5: Quantitative performance comparison of VAMP-Net architectural variants on the RIF test set.
+The table evaluates the influence of attention masking, data augmentation (shuffling), and dual-path integration across four key metrics.
+Model variant Accuracy Bal. Acc. F1-score ROC-AUC
+A: VAMP-Net (VN) 0.956 0.944 0.960 0.970
+B: VN + shuffle 0.945 0.941 0.950 0.963
+C: No mask, no shuffle 0.944 0.942 0.953 0.973
+D: No mask + shuffle 0.945 0.943 0.953 0.969
+E: No Path-2 + shuffle 0.940 0.930 0.940 0.961
+F: No Path-2, no shuffle 0.940 0.935 0.941 0.960
+4.2.4 VAMP-Net for anti-TB Drugs
+The predictive performance of VAMP-Net across four primary anti-TB drugs is detailed in Table 6.
+To rigorously evaluate model stability, we report metrics across two dimensions: validation stability (top panel) and final unseen test set efficacy (bottom panel).
+Across three independent experimental seeds, VAMP-Net demonstrated high architectural stability.
+Standard deviations for all primary metrics remained consistently low, with ROC-AUC variance specifically constrained to <0.007<0.007.
+On the test set, INH achieved the highest overall performance, recording an accuracy of 0.945 $\pm$ 0.005 and a Matthews Correlation Coefficient (MCC) of 0.891 $\pm$ 0.010.
+RIF exhibited a similarly robust profile with an F1-score of 0.946 $\pm$ 0.007.
+For RFB, the model achieved its peak discriminative power with an ROC-AUC of 0.975 $\pm$ 0.001.
+A notable observation is the parity between Accuracy and Balanced Accuracy across most drugs (e.g., identical values of 0.945 for INH and a marginal 0.01 difference for EMB).
+Despite the inherent complexity of EMB resistance, VAMP-Net maintained the highest PR-AUC of the cohort at 0.977 $\pm$ 0.002.
+Collectively, the MCC values remained high across the spectrum (ranging from 0.614 to 0.891), indicating that the model maintains high-quality predictions across both resistant and susceptible classes.
+Table 6: Statistical performance profile of VAMP-Net across first-line and second-line anti-TB drugs.
+The top panel provides validation stability metrics (Mean $\pm$ SD) across three independent experimental seeds to assess model convergence.
+The bottom panel details the final predictive performance on the unseen test set.
+We report Accuracy, Balanced Accuracy, F1-score, Matthews Correlation Coefficient (MCC), and Area Under the ROC and Precision-Recall Curves (ROC-AUC, PR-AUC).
+Drug Accuracy Bal. Acc. F1 MCC ROC-AUC PR-AUC
+RIF 0.937 $\pm$ 0.010 0.937 $\pm$ 0.013 0.946 $\pm$ 0.008 0.871 $\pm$ 0.022 0.967 $\pm$ 0.004 0.970 $\pm$ 0.005
+RFB 0.925 $\pm$ 0.003 0.927 $\pm$ 0.002 0.924 $\pm$ 0.002 0.853 $\pm$ 0.003 0.971 $\pm$ 0.001 0.963 $\pm$ 0.002
+EMB 0.839 $\pm$ 0.006 0.851 $\pm$ 0.007 0.892 $\pm$ 0.005 0.607 $\pm$ 0.007 0.915 $\pm$ 0.007 0.975 $\pm$ 0.002
+INH 0.935 $\pm$ 0.007 0.935 $\pm$ 0.007 0.936 $\pm$ 0.007 0.870 $\pm$ 0.014 0.964 $\pm$ 0.000 0.953 $\pm$ 0.002
+Test metrics:
+Drug Accuracy Bal.
+Acc.
+F1 MCC ROC-AUC PR-AUC RIF 0.937 $\pm$ 0.009 0.935 $\pm$ 0.013 0.946 $\pm$ 0.007 0.870 $\pm$ 0.019 0.968 $\pm$ 0.006 0.970 $\pm$ 0.009 RFB 0.935 $\pm$ 0.008 0.937 $\pm$ 0.007 0.934 $\pm$ 0.006 0.874 $\pm$ 0.013 0.975 $\pm$ 0.001 0.960 $\pm$ 0.003 EMB 0.843 $\pm$ 0.014 0.853 $\pm$ 0.009 0.895 $\pm$ 0.010 0.614 $\pm$ 0.022 0.920 $\pm$ 0.004 0.977 $\pm$ 0.002 INH 0.945 $\pm$ 0.005 0.945 $\pm$ 0.005 0.945 $\pm$ 0.005 0.891 $\pm$ 0.010 0.970 $\pm$ 0.002 0.961 $\pm$ 0.004
+Figure 9: Comparative test set performance across first-line and second-line anti-TB drugs.
+The bar chart illustrates the mean performance metrics across three independent experimental seeds for RIF, RFB, INH, and EMB.
+Error bars represent the standard deviation ($\pm$ SD), indicating high model stability across different initializations.
+Star markers ($\star$) denote the peak performance achieved for specific drug-metric pairs.
+To evaluate the clinical readiness of VAMP-Net, we conducted a multi-dimensional analysis focusing on predictive accuracy, computational throughput, and stochastic stability across four major anti-TB drugs.
+As illustrated in Figure 9, VAMP-Net demonstrated high discriminative performance across the cohort.
+The model achieved peak ROC-AUC values exceeding $0.96$ for RIF, RFB, and INH, with INH specifically reaching a mean Accuracy and Balanced Accuracy of $0.945\pm 0.005$.
+While EMB presented the most challenging classification task, indicated by a lower Matthews Correlation Coefficient (MCC) of $0.614$, it maintained a superior Precision-Recall AUC ($0.977$), suggesting that the model remains highly effective at identifying true resistant isolates despite significant class imbalance.
+The calibration of our model’s probabilistic outputs was assessed using the Brier Score (Figure 10 Left).
+VAMP-Net exhibited high calibration stability for INH ($\sim$0.05) and moderate calibration for RIF and RFB.
+The higher Brier score for EMB ($\sim$0.14) aligns with its lower MCC, further highlighting the biological complexity inherent in Ethambutol resistance pathways.
+Simultaneously, inference speed was evaluated to determine suitability for clinical deployment.
+VAMP-Net maintained a high-throughput processing rate of $\sim$490 samples per second across all drugs ( Figure 10 Right), ensuring near-instantaneous diagnostics even for large genomic surveillance datasets.
+To ensure the reported metrics were not artifacts of favorable initialization, we measured per-seed variability (Figure 11).
+The results show tight clustering around the mean for ROC-AUC, F1-score, and MCC across three independent seeds.
+The negligible variance in RIF and INH performance underscores the architectural robustness of our attention-based framework.
+Figure 10: Computational efficiency and calibration across targeted anti-TB drugs.
+Left: Boxplots illustrate the distribution of Brier scores across three independent seeds, where a lower score indicates superior model calibration and probabilistic accuracy.
+Right: Throughput analysis measured in samples per second during inference.
+VAMP-Net maintains a consistent high-speed processing rate across all drugs, averaging $\sim$490 samples/s.
+Error bars represent the standard deviation ($\pm$ SD) across three experimental runs.
+Figure 11: Per-seed variability of key performance metrics across three independent experimental runs.
+The scatter-boxplots illustrate the distribution of ROC-AUC (left), F1-score (center), and Matthews Correlation Coefficient (MCC, right) for RIF, RFB, INH, and EMB.
+Individual points represent specific stochastic seeds (Seed 42, Seed 123, and Seed 2024), while the diamond marker ($\diamond$) denotes the arithmetic mean.
+Error bars indicate the range of variability across runs.
+The high degree of clustering around the mean for RIF, RFB, and INH demonstrates the architectural stability of VAMP-Net.
+While EMB exhibits slightly higher dispersion, reflecting its greater biological and dataset complexity, the consistent performance across different initializations underscores the robustness of the Set Attention mechanism in capturing significant resistance signatures.
+4.3 Comparative Analysis
+4.3.1 Comparative Performance Against Baseline Architectures
+We benchmarked our approach against the traditional binary encoding paradigm prevalent in antimicrobial resistance (AMR) and clinical genomics literature [36].
+In this conventional framework, genomic data is represented as a high-dimensional sparse matrix $\mathbf{X}\in\{0,1\}^{N\times V}$, where $X_{i,j}$ denotes the presence or absence of variant $j$ in sample $i$.
+While effective in certain contexts, this representation suffers from the curse of dimensionality, necessitating significant computational resources or aggressive feature selection, such as PCA or statistical pruning, to manage the sparse feature space [34, 33, 48].
+To ensure a fair comparison, we performed feature selection using a Chi-squared test ($\alpha=0.001$) to retain significantly associated variants, resulting in a reduced set of 11,567 features ($\sim$36.1% reduction) used to train VAMP-Net and competitive Multi-Layer Perceptron (MLP) and Convolutional Neural Network (CNN) baselines applied on RIF drug.
+The optimazed hyperparameter configurations for MLP and CNN are presented in Tables 7, and 8 respectively.
+Table 7: Optimized hyperparameter configuration for the MLP baseline.
+The table outlines the best-performing structural parameters determined via empirical tuning on Rifampicin (RIF) resistance prediction.This specific configuration is referred to as Model G
+Model Hidden Dims Dropout Activation
+Model G [1024, 128, 128] 0.1315 gelu
+Table 8: Optimized hyperparameter configuration for the CNN baseline.
+The table outlines the best-performing structural parameters determined via empirical tuning on Rifampicin (RIF) resistance prediction.This specific configuration is referred to as Model H
+Model Dropout Activation Conv Layers Filters Kernel Sizes Pool Sizes
+Model H 0.1487 relu 3 [96, 128, 256] [5, 9, 2] [3, 3, 3]
+As illustrated in Figure 12, VAMP-Net consistently establishes a new performance ceiling across all evaluated metrics.
+The ROC curve analysis reveals that while the CNN achieves a competitive AUC of 0.96, VAMP-Net reaches a superior 0.97, demonstrating a more robust trade-off between sensitivity and specificity.
+The MLP, lacking specialized spatial or attention-based feature extraction, lags significantly with an AUC of 0.87.
+The granularity of this superiority is further evidenced by the confusion matrices on the unseen test set (Figure 13).
+VAMP-Net correctly identifies 92.9% of resistant isolates, a notable improvement over the CNN (89.4%) and the MLP (84.2%).
+Crucially, as shown in Table 9, the performance gains are statistically validated; VAMP-Net achieved a significant improvement in Accuracy over the CNN ($1.35\%$, $p<0.001$) and the MLP ($8.48\%$, $p<0.001$).
+The 95% Confidence Intervals for VAMP-Net remain narrow (e.g., Accuracy: $[0.9383,0.9554]$), underscoring the model’s high reliability and precision in a clinical diagnostic context.
+Figure 12: Comparative performance of VAMP-Net against baseline architectures on the test set.
+(Left) Receiver Operating Characteristic (ROC) curves demonstrating the trade-off between true positive and false positive rates.
+(Right) Detailed test set metrics across six performance indicators.
+Figure 13: Comparative confusion matrix analysis on the unseen test set.
+This figure provides a head-to-head performance breakdown between the proposed VAMP-Net architecture and two baseline models: a Convolutional Neural Network (CNN) and a Multi-Layer Perceptron (MLP).
+Table 9: VAMP-Net overall performance and statistical significance comparison against baselines.
+Metric Value 95% CI p-value (vs Baselines)
+Accuracy 0.9474 [0.9383, 0.9554] $0.001^{**}$
+F1-score 0.9555 [0.9476, 0.9627] $0.001^{**}$
+ROC-AUC 0.9691 [0.9612, 0.9762] 0.074
+∗∗Significant at $p<0.001$ relative to CNN and MLP baselines.
+4.3.2 Comparative Evaluation of Fusion Paradigms
+The architectural efficacy of VAMP-Net’s late-stage adaptive fusion was benchmarked against a traditional Early Fusion baseline, applied to RIF drug, to quantify the impact of integration timing on model performance.
+As summarized in Table 10, VAMP-Net significantly outperformed the Early Fusion paradigm across all evaluated metrics.
+Most notably, VAMP-Net achieved transformative gains in Specificity (+48.8%), Balanced Accuracy (+32.6%), and ROC-AUC (+30.9%) compared to the early-integration approach.
+The Early Fusion model exhibited a profound performance asymmetry, demonstrating high Sensitivity (0.805) but catastrophic Specificity (0.425).
+This imbalance suggests a systematic failure to distinguish true genomic resistance markers from stochastic noise.
+In contrast, VAMP-Net maintained a harmonized performance profile, achieving a Recall of 0.968 and a Specificity of 0.913.
+The empirical results indicate that forcing inter-modality interaction at a premature stage—prior to dedicated contextual modeling—is suboptimal for complex genomic tasks.
+In the Early Fusion setting, token-level embeddings and structured features are concatenated and projected before entering the Set Attention Blocks (SABs).
+Our findings suggest that this approach leads to the propagation of unrefined signal noise, which the subsequent attention mechanisms struggle to resolve, as evidenced by the severe specificity deficit.
+VAMP-Net’s superiority is rooted in its late-stage adaptive fusion philosophy.
+By postponing integration until after modality-specific encoding, each pathway is permitted to develop robust internal representations.
+This allows the SAB encoder to model inter-variant dependencies within a more stable feature space.
+Table 10: Comparative analysis of fusion architectural paradigms.
+This table benchmarks the proposed VAMP-Net multi-path integration (late-stage adaptive fusion) against a traditional Early Fusion baseline.
+Metric Early Fusion VAMP-Net Fusion
+Accuracy 0.649 0.946
+F1-score 0.729 0.955
+Sensitivity (Recall) 0.805 0.968
+Specificity 0.425 0.913
+Precision (PPV) 0.667 0.941
+NPV 0.605 0.953
+Balanced Accuracy 0.615 0.941
+ROC-AUC 0.655 0.964
+4.3.3 Comparative SOTA
+To provide a more faithful comparison, we separate prior work by paper and retain only the metrics explicitly reported in each study.
+For every drug shared with our benchmark, we compare the published result against the best-performing VAMP-Net seed for that drug on our validation protocol.
+Because these studies use different cohorts, split strategies, and prediction formulations, the following tables should be interpreted as paper-specific contextual comparisons rather than direct leaderboard claims.
+Table 11: Paper-specific comparison with the XGBoost study in Diagnostics [31]. The original paper reports Sensitivity, Specificity, Precision, F1-score, and Accuracy for EMB, INH, and RIF.
+Drug Our Seed Sens. (Paper) Sens. (Ours) Spec. (Paper) Spec. (Ours) Prec. (Paper) Prec. (Ours) F1 (Paper) F1 (Ours) Acc. (Paper) Acc. (Ours)
+EMB 42 0.97 0.8166 0.97 0.8885 0.89 0.9673 0.93 0.8856 0.97 0.8309
+INH 42 0.90 0.9382 0.99 0.9439 0.97 0.9441 0.94 0.9411 0.97 0.9410
+RIF 2024 0.94 0.9360 0.96 0.9501 0.90 0.9640 0.92 0.9498 0.96 0.9418
+Compared with the XGBoost results reported in Diagnostics, VAMP-Net shows a more favorable precision profile across all shared drugs, with particularly strong gains for EMB and RIF, indicating that the proposed architecture is more conservative in assigning resistant calls while preserving competitive sensitivity.
+The main trade-off appears for EMB, where the prior study reports higher sensitivity and accuracy, suggesting that EMB remains the most challenging phenotype in our setting and may be more sensitive to cohort composition and decision-threshold effects.
+Table 12: Paper-specific comparison with the MIC-prediction machine learning system reported in PLOS Computational Biology [10].
+Only metrics shared with our evaluation are shown; the paper reports PPV, which is reported here as the published precision-equivalent quantity.
+Drug Our Seed AUC (Paper) AUC (Ours) Sens. (Paper) Sens. (Ours) Spec. (Paper) Spec. (Ours) PPV/Prec. (Paper) Prec. (Ours) Acc. (Paper) Acc. (Ours)
+EMB 42 0.965 0.9236 0.931 0.8166 0.894 0.8885 0.704 0.9673 0.902 0.8309
+INH 42 0.983 0.9696 0.950 0.9382 0.981 0.9439 0.980 0.9441 0.966 0.9410
+RIF 2024 0.992 0.9712 0.971 0.9360 0.973 0.9501 0.962 0.9640 0.972 0.9418
+RFB 2024 0.988 0.9733 0.955 0.9782 0.957 0.8882 0.918 0.8874 0.956 0.9309
+Relative to the MIC-oriented model from PLOS Computational Biology, VAMP-Net remains competitive but reveals a mixed pattern.
+Our model achieves its closest agreement on INH and RFB, while the largest performance gaps are observed for EMB and RIF, especially in AUC and accuracy.
+This suggests that VAMP-Net is strongest when technical-confidence recalibration contributes substantially to the final decision, but that MIC-prediction frameworks trained under different targets and evaluation designs can retain an advantage on some first-line drugs.
+Table 13: Paper-specific comparison with LLMTB, the BERT-based large language model reported in Bioinformatics [45]. The paper reports positive-class F1-scores on the shared drugs.
+Drug Our Seed F1 (LLMTB) F1 (Ours)
+EMB 42 0.8273 0.8856
+INH 42 0.9361 0.9411
+RIF 2024 0.9094 0.9498
+RFB 2024 0.8913 0.9306
+Against LLMTB, VAMP-Net improves the reported F1-score for all four shared drugs, with the clearest gains on EMB, RIF, and RFB.
+This pattern supports the value of combining permutation-invariant biological modeling with an explicit quality-aware branch, rather than relying solely on sequence-style language representations.
+The smaller margin on INH further suggests that both architectures are already highly effective for this phenotype, leaving less headroom for separation.
+4.4 Path-2 Contribution and Case Studies:
+To better understand how the quality-aware Path-2 branch influences final predictions, we inspected representative test cases in which the fused model either corrected an overconfident Path-1 decision or amplified a borderline but correct Path-1 prediction.
+Table 14 summarizes these case studies using the resistance probability predicted by the variant-only pathway (Path-1) and by the full fused model.
+We additionally report the probability shift $\Delta P=P_{\mathrm{Full}}(\mathrm{Resistant})-P_{\mathrm{Path\mbox{-}1}}(\mathrm{Resistant})$, together with summary statistics of the learned fusion gate.
+These examples show that the VCF-aware Path-2 branch is not merely an auxiliary source of features, but an active recalibration mechanism for Path-1’s variant-based decision.
+In the corrected cases, Path-1 predicted resistance with very high confidence ($P(R)\approx 0.91$), yet the fused model reduced this probability to 0.040.040.060.06, corresponding to a large decision shift of roughly 0.86-0.86.
+This is not a marginal adjustment; it is a strong reversal of an initially overconfident prediction.
+Table 14: Ablation case studies of VCF-aware Path-2 contribution.
+All samples feature a variant count of 743.
+“Corrected” denotes cases where Path-B reversed an incorrect Path-1 prediction, while “Amplified” denotes Path-2 reinforcing a correct class by shifting borderline predictions away from the decision boundary.
+True Label Path-1 $P(R)$ Full $P(R)$ $\Delta P$ Effect Principal Explanation
+Susceptible 0.9190 0.0521 0.867-0.867 Corrected High FRS but low genotype-confidence and depth support made Path-1’s resistance call unreliable.
+Susceptible 0.9049 0.0438 0.861-0.861 Corrected Path-A overcalled resistance; Path-2 utilized VCF-support signals to downweight the prediction.
+Susceptible 0.9185 0.0592 0.859-0.859 Corrected Low-confidence and low-depth evidence triggered the fused model to reverse the initial resistant call.
+Susceptible 0.4579 0.0774 0.381-0.381 Amplified Borderline sample pushed further into the susceptible class, increasing overall model confidence.
+Resistant 0.5491 0.9028 +0.354+0.354 Amplified Combined variant identity with consistent support features to strengthen a true resistant call.
+Resistant 0.5495 0.8826 +0.333+0.333 Amplified Reinforced the resistant signal, moving a borderline Path-1 output safely away from the decision boundary.
+Inspection of the associated feature profiles clarifies why this occurs.
+In these corrected cases, FRS remained high, indicating that the called alleles were internally supported by reads, but the broader VCF context, particularly GT_CONF_PERCENTILE, depth-related features, and filtered depth support, remained weak.
+Thus, although Path-1 detected a mutation pattern that appeared resistance-like, Path-2 identified insufficient technical evidence to trust that pattern as a robust resistance signal.
+In this sense, Path-2 behaves as a reliability-aware filter: variant presence alone is not treated as sufficient, and the model instead asks whether the surrounding sequencing evidence is strong and consistent enough to justify a resistant call.
+The amplified cases demonstrate the complementary behavior.
+When Path-1 produced borderline predictions, Path-2 was able to push the fused output farther toward the correct class.
+For truly resistant isolates, the full model increased the resistance probability from about $0.55$ to 0.880.880.900.90, showing that the quality-aware branch does not only suppress unreliable signals, but can also strengthen biologically meaningful signals when the VCF support is coherent.
+Overall, these case studies show that the fused model can correct Path-1 false positives by incorporating genotype confidence, depth, and coverage-derived evidence, thereby reducing spurious resistance probabilities by as much as $0.87$ and yielding more reliable susceptibility calls.
+Figure 14 provides a visual quantification of the "corrective" and "amplificatory" influence of the Path-2 quality-aware branch.
+By plotting the resistance probability of Path-1 against the final model output, we can observe the specific instances where sequencing metadata fundamentally altered the diagnostic outcome.
+Figure 14: Probabilistic recalibration of resistance calls through multi-path fusion.
+The scatter plot illustrates the shift in resistance probability $P(\text{Resistant})$ when transitioning from the variant-only model (Path-1) to the fully integrated VAMP-Net architecture (Full Model).
+Points in the lower-right quadrant represent "Corrected" cases, where Path-1 initially assigned high resistance probabilities (>0.5>0.5) that were subsequently reversed by Path-2’s quality-aware branch to confident susceptible calls (<0.2<0.2).
+Conversely, points in the upper-left quadrant denote "Amplified" cases, where Path-2 reinforced borderline or incorrect susceptible predictions into high-confidence resistance calls.
+The significant deviation from the "No change" diagonal underscores Path2’s role as a critical reliability filter that identifies and corrects overconfident predictions based on underlying sequencing evidence.
+4.5 Multi-Path Network Interpretation and Epistasis
+This section prensents the results of the Dual-Path Interpretability Network and leverages the insights to dissect the model’s decision-making process for RIF, RFB, INH and EMB resistance prediction.
 Our analysis demonstrates that the two architectural pathways within VAMP-Net provide distinct yet complementary perspectives on the underlying mechanisms of drug resistance.
@@
 In contrast, Path-2 (Quality-Aware CNN) focused on quantifying the influence of technical confidence metrics, revealing how the model performs noise regularization and feature calibration.
-By integrating the findings from both streams, our framework provides a comprehensive and auditable explanation of the model’s decision process —both biological and technical— that drives the final drug resistance classification.
-5.5.1 Set-Transformer Genetic Variant Importance
-This section details the quantitative analysis of individual variant contributions to drug resistance, as determined by the Integrated Gradients (IG) method applied to the SAB pathway (Path-1).
-The IG-derived attribution scores, aggregated across the test cohort, establish an empirically derived ranking of mean importance for all unique ChromPos_Ref>>Alt variants.
-This analysis serves to validate the model’s ability to prioritize known resistance drivers while simultaneously discovering highly-ranked non-canonical variants that warrant further investigation.
-The subsequent analyses distinguish between the variant importance profiles learned for Rifampicin (RIF) and Rifabutin (RFB), highlighting the model’s discriminatory power.
-A) Drug 1: RIF Resistance Analysis
-The analysis of the VAMP-Net pathway for RIF resistance achieved strong validation success, presented in Table 6, demonstrating the framework’s ability to learn and prioritize fundamental genetic mechanisms as reported in the literature (telenti1993detection; ramaswamy1998molecular).
-The identified variants in rpoB, katG, and other genes align with established resistance mechanisms documented in the WHO catalogue of mutations associated with drug resistance (world2023catalogue).
-Table 6: Among 200 Top-Ranked Variants from RIF Resistance Analysis out of 417,635 Unique Variants
-Variant Importance Score Gene
+By integrating the findings from both streams, our network provides a comprehensive and auditable explanation of the model’s decision process, both biological and technical, that drives the final drug resistance classification.
+4.5.1 Mechanistic Interpretation and Variant Attribution
+To quantify the individual genetic contributions to drug resistance, we applied the Integrated Gradients (IG) method to the SAB pathway.
+As detailed in Table 15, the resulting importance scores establish a hierarchical signature that validates the network’s ability to autonomously rediscover established molecular pathways while proposing high-impact novel candidates.
+Table 15: Top-ranked genomic variants and associated genes identified by VAMP-Net.
+The table lists high-importance variants prioritized by the model’s attention mechanism for RIF, RFB, INH and EMB.
+While the model successfully recovers established resistance determinants in the rpoB, katG, and ethA genes, it also identifies several "novel" high-impact loci in genes such as ponA1, pknB, and mmpL5.
+Drug Variant Importance Score Gene
+RIF 2155162_ATGCCGC>>ACGCCGT 0.446 katG
+2338202_CAC>>CCAC 0.308 ponA1 (novel)
+424791_T>>TAT 0.273 pknB (novel)
+1849861_G>>T 0.207 ethA
 761139_C>>A 0.178 rpoB
 761101_A>>T 0.116 rpoB
-761140_A>>T 0.058 rpoB
-761154_TCG>>TTC 0.050 rpoB
-761108_GGA>>GTT 0.044 rpoB
-760314_G>>T 0.029 rpoB
-2155162_ATGCCGC>>ACGCCGT 0.446 katG
-2155166_CGC>>CTG 0.043 katG
-7581_G>>T 0.050 gyrA
-4247574_A>>C 0.036 embB
-1849861_G>>T 0.207 ethA
-2338202_CAC>>CCAC 0.308 ponA1 (novel)
-424791_T>>TAT 0.273 pknB (novel)
-The model assigned primary predictive importance to variants within the RNA polymerase $\beta$ subunit (rpoB) gene.
-Multiple high-confidence variants in Table 6, including 761139_C>>A, 761101_A>>T, and 761140_A>>T, are located within the rpoB gene and fall precisely within the Rifampicin Resistance-Determining Region (RRDR).
-This alignment with established clinical and molecular biology literature validates the framework’s capacity to learn fundamental resistance pathways (musser1995antimicrobial).
-Beyond the primary rpoB target, the VAMP-Net successfully extracted complex, multi-gene resistance signatures.
-The model assigned high importance scores to variants associated with isoniazid resistance, including mutations in the katG region (2155162_ATGCCGC>>ACGCCGT) and the promoter of ahpC (78519_G>>A) (vilcheze2007mechanism; napier2023large).
-This demonstrates the framework’s ability to capture complex multi-drug resistance profiles where resistance to one drug often co-occurs with markers for another (jia2021analysis; billows2024large).
-A significant finding is the identification of top-ranked variants located in previously uncharacterized genomic regions (Novel loci).
-The high Importance Scores for these novel variants (such as ponA1 and pknB) suggest the model is uniquely capable of isolating:
-  • Rare, High-Effect Mutations: Highly penetrant variants that evade detection by standard frequency-based GWAS methods.
-  • Critical Interaction Hubs: Loci that function as epistatic modifiers, significantly altering the expression or effect of primary resistance mutations.
-In summary, the high attribution results confirm the robustness and superior discovery potential of the VAMP-Net, which can simultaneously validate known mechanisms and identify high-impact novel loci for targeted functional investigation.
-B) Drug 2: RFB Resistance Analysis
-The RFB resistance analysis successfully identified the complete spectrum of variants associated with the phenotype (Table 7), confirming the model’s high fidelity in identifying key resistance mechanisms.
-Table 7: Among 200 Top-Ranked Variants from RFB Resistance Analysis out of 417,635 Unique Variants
-Variant Importance Score Gene
-761140_A>>T 6.896 rpoB
-761139_C>>A 3.955 rpoB
-761140_A>>G 3.321 rpoB
-761110_A>>G 3.316 rpoB
-761110_A>>T 2.463 rpoB
-4247431_G>>T 2.125 embB
-7581_G>>A 1.323 gyrA
+RFB 761140_A>>T 6.896 rpoB
 2944085_C>>T 4.947 mmpL5 (novel)
 893702_GAC…CCG 4.213 lpqS (novel)
+761139_C>>A 3.955 rpoB
 3640407_G>>A 3.532 ppsD (novel)
@@
 1262023_C>>T 2.885 pks13 (novel)
-4217696_TGGAG>>TTGAA 2.594 papA4 (novel)
-1262193_C>>T 2.589 pks13 (novel)
-The analysis demonstrated the framework’s superior capability to identify both canonical and non-canonical resistance mechanisms.
-As expected, rpoB coding region variants ranked most highly, with the top-scoring variant 761140_A>T (Score 6.896) confirming the model’s alignment with the shared target mechanism between RFB and RIF (williams1998contribution; van2013rifampin; wang2023antimicrobial).
-Multiple additional rpoB variants achieved high scores, including 761139_C>A (Score 3.955), 761140_A>G (Score 3.321), 761110_A>G (Score 3.316), and 761110_A>T (Score 2.463), all clustering within the Rifampicin Resistance-Determining Region (RRDR).
-Furthermore, the framework successfully captured a Multi-Drug Resistance (MDR) genomic signature from a single drug prediction task.
-The detection of relevant variants in embB (4247431_G>T) and gyrA (7581_G>A) confirms the learned importance of co-selection and cross-resistance patterns (ghosh2020survey).
-The discovery potential of the framework is underscored by the high ranking of numerous novel variants.
-High scores for variants in genes such as mmpL5 (2944085_C>T, Score 4.947), lpqS (893702_GAC...CCG, Score 4.213), ppsD (3640407_G>A, Score 3.532) reinforce the strong suggestion that these uncharacterized loci represent crucial epistatic interaction points or rare, high-effect genes that contribute significantly to the RFB phenotype.
-These variants are now prime candidates for functional validation to uncover new biological resistance pathways.
-5.5.2 Genetic Variant Interaction Network Analysis
-Analysis of the aggregated Set Attention Block (SAB) weights culminates in the Multi-locus Attribution-Weighted Interaction (MAWI) Map (Figures 11 and 12), which provides a critical, quantifiable decomposition of complex, multi-locus epistatic effects underlying Rifampicin (RIF) and Rifabutin (RFB) resistance.
-This approach moves beyond single-variant association models to capture the necessary synergistic coordination between mutations.
-Specifically, the heatmaps visualize the highest magnitude of attention-weights found by the SAB transformer, thereby mapping the strength and direction of these interactions and providing mechanistic insights into the polygenic architecture of drug resistance.
-Figure 11: Variant interaction heatmaps showing attention-based relationships between top hub variants for RIF resistance.
-Red indicates strong interactions; yellow indicates weak interactions.
-The patterns reveal how variants work together in resistance mechanisms
-The variant interaction heatmap (Figure 11) for RIF (Rifampicin) resistance reveals a highly interconnected, yet structured, network of epistatic relationships among the top hub variants, with a high overall signal density indicating numerous strong-to-moderate synergistic interactions.
-This high-interaction density suggests that robust resistance relies on compensatory or cooperative pathways, confirming that the model has effectively learned the structural dependencies of a highly interconnected resistance mechanism.
-The pattern is characterized by specific "hub" variants, such as 3829770_T>>C and 1307598_C>>G, which act as central nodes, exhibiting dark red cells that indicate strong attention to multiple other variants in the network.
-This hub-driven structure supports the idea that the mechanism relies on a few key "linchpin" mutations that coordinate with a set of secondary, supportive mutations, and this strong cross-attention quantitatively supports the hypothesis of epistatic fitness compensation.
-Furthermore, the scattered distribution of high-value cells across the map, rather than being localized to a single cluster, suggests a diffuse polygenic mechanism where many different combinations of mutations contribute significantly to the resistance phenotype.
-Figure 12: Variant interaction heatmaps showing attention-based relationships between top hub variants for RFB resistance.
-Red indicates strong interactions; blue indicates weak interactions.
-The patterns reveal how variants work together in resistance mechanisms
-The second heatmap in Figure 12 (blue-to-red scale, max interaction $\approx 0.40$) depicts a contrasting, sparse network of high-magnitude, localized epistatic effects for RFB resistance.
-In contrast to Figure 11, this visualization shows a predominantly low-interaction background (dark blue), highlighting a few critical, high-leverage interactions that dominate the attention landscape involving variants like 1070702_T>>C.
-The high attention-weights (dark red) pattern suggests that these specific variant interactions act as a primary mechanism for resistance, where the effect of one variant is highly dependent on the presence of the other.
-The sparsity indicates that, while many variants are present, only a select few combinations are mechanistically determinative, allowing the VAMP-Net model to pinpoint the most non-additive synergistic effects.
-The contrasting patterns between the two drugs demonstrate that resistance involves coordinated resistance modules whose composition varies depending on the specific drug, providing a key insight for designing combination therapies.
-The MAWI provides an interpretable, quantitative visualization of complex genetic epistasis, confirming known biological interactions while highlighting high-impact, potentially novel interaction hubs.
-5.5.3 VCF Feature Channel Analysis: Quantifying the Influence of Variant Quality Metrics
-We quantified the contribution of the eight VCF features via systematic ablation, validating the unique multi-channel design of the Quality-Aware 1D-CNN pathway (Figure 13).
-This analysis highlights the technical robustness required for reliable genotype-phenotype association.
-The analysis confirms the model does not rely on simple variant presence (GT field) but critically leverages the contextual quality and support metrics:
-  • Dominant Metric (FRS): The Fraction of Supporting Reads (FRS) is the single most influential feature for both the RIF and RFB models.
-This strong reliance confirms that the 1D-CNN pathway effectively learns to filter out false positives or low-support variants, as high FRS is used as a primary filter for calling true clonal mutations in MTB sequencing (cryptic2022data).
-  • Contextual Confidence: The GT_CONF_PERCENTILE is the second most important feature.
-By utilizing the relative confidence rank rather than the raw GT_CONF (which shows minimal contribution), the model demonstrates its capacity to leverage comparative, sample-specific context for prediction.
-Figure 13: VCF feature importance comparison for RIF and RFB resistance models.
-Fraction of Supporting Reads (FRS) and Genotype Confidence Percentile (GT_CONF_PERCENTILE) demonstrate substantially higher importance for both resistance phenotypes, with the RFB model showing greater dependence on these quality metrics
-The near-zero composite scores for features like Read Depth (DP), Genotype (GT), and Raw GT_CONF are equally informative.
-The low importance of DP suggests its information is effectively captured by the derived, normalized metrics (FRS and GT_CONF_PERCENTILE), rendering the raw depth redundant.
-This indicates that the presence of a variant or its raw confidence is insufficient; the model requires the granular, multi-dimensional context of the variant’s support and relative quality to make high-confidence resistance predictions.
-This ablation study confirms that the 1D-CNN pathway successfully extracts biologically and technically meaningful information from the VCF features, enabling a robust distinction between true resistance-conferring variants and lower-quality calls, especially critical for subtle phenotypes like RFB resistance.
-5.5.4 Dual-Path Fusion Model Interpretability
-This final analysis validates the Multi-Path Fusion strategy as the core functional advancement of VAMP-Net, establishing its ability to combine Pathogenicity (via the Set-Transformer) and Reliability (via the 1D-CNN).
-The analysis revealed two novel findings critical to the framework’s superior performance and generalizability:
-  • Learned Adaptive Feature Regularization: The 1D-CNN pathway functions as a learned signal-to-noise regulator, performing non-linear, context-aware feature selection (prioritizing FRS and GT_CONF_PERCENTILE over redundant features like DP).
-The most significant evidence of this is the 8 to 14-fold increase in feature reliance observed in the RFB model, which demonstrates that the fusion mechanism dynamically compensates for the weaker, more diverse genetic signal of that phenotype by demanding greater technical certainty.
-Path-2 thus acts as an adaptive attention mechanism that modulates prediction confidence based on data quality.
-  • Epistatic Network Auditing: The SAB network analysis confirmed that the final prediction is governed by an epistatic network audit, where the model learns that a high-importance variant from Path-1 will have its final impact dynamically down-weighted if Path-2 reports low sequencing support.
-This integrated approach provides an auditable, high-confidence score that addresses the critical challenge of separating biological signal from computational support in clinical genomic data.
-In conclusion, integrating VCF features in this manner addresses a major challenge in genomic deep learning: the inherent separation between the biological signal (variant identity) and the computational support (variant quality).
+EMB 740348_TCGCGCGAC>>T 0.756 lipG (novel)
+764509_G>>C 0.638 rpoC (novel)
+740383_G>>T 0.559 lipG (novel)
+2435103_T>>C 0.536 idsA2 (novel)
+4269298_A>>G 0.180 ubiA (known)
+4247429_A>>C 0.062 embB (known)
+4247429_A>>T 0.055 embB (known)
+4248003_A>>G 0.046 embB (known)
+2603465_C>>G 0.502 intergenic (regulatory candidate)
+2113033_G>>C 0.363 intergenic (regulatory candidate)
+INH 2760136_TCC…ATA>>TTC…ATG 8.195 mmuM (novel)
+4366246_GCC…CG>>GGC…GCC 0.763 eccA2 (novel)
+575699_G>>A 0.085 mshA (known)
+413294_ACCC…>>A 0.097 iniC (known)
+1503392_ACATG…GGTTG>>A 0.527 intergenic (regulatory candidate)
+1113698_CACG…GATGA>>C 0.502 intergenic (regulatory candidate)
+VAMP-Net assigned primary predictive weight to canonical mutations within the RRDR of the $rpoB$ gene, precisely aligning with WHO-validated markers for Rifampicin (RIF) and Rifabutin (RFB), where the top-ranked RFB variant 761140_A>>T achieved a dominant score of 6.896.
+Notably, the model extracted complex Multi-Drug Resistance (MDR) signatures from single-drug tasks; for RIF, the highest importance score (0.446) was attributed to the Isoniazid-associated $katG$ variant, while for Ethambutol (EMB), the model correctly prioritized the $rpoC$ variant (0.638).
+This confirms that VAMP-Net leverages co-selection patterns and compensatory mechanisms inherent in MDR-TB lineages to enhance its predictive confidence.A critical outcome of the attribution analysis is the identification of high-scoring variants in uncharacterized regions, most significantly the mmuM variant for Isoniazid (INH), which attained the study’s highest overall importance score of 8.195.
+Other key discoveries include $ponA1$ (0.308) and $pknB$ (0.273) for RIF, $mmpL5$ (4.947) for RFB, and the novel $lipG$ (0.756) and $idsA2$ (0.536) loci for EMB.
+The consistent prioritization of these novel loci across all four drugs suggests that VAMP-Net’s attention mechanism successfully isolates rare, high-effect mutations and epistatic modifiers that are often missed by standard frequency-based methods.
+4.5.2 Functional Enrichment
+To determine whether the high-attribution features identified by VAMP-Net reflect genuine biological mechanisms, we cross-referenced the top-ranked variants against established proteomic and molecular literature.
+As detailed in Table 16, our framework successfully identified a hierarchical resistance signature that spans from canonical target modification to global cell-envelope remodeling.
+VAMP-Net correctly prioritized rpoB variants, specifically distinguishing the unique structural binding affinities of Rifabutin (RFB) over Rifampicin (RIF) through the selection of variants like 761140_A>>T (IG: 6.896) [21].
+Beyond canonical targets, the framework isolated broad-spectrum efflux mechanisms, notably the mmpL5 system (IG: 4.947), which likely modulates RFB susceptibility through active drug extrusion [18].
+Additionally, the high ranking of the persistence-related lpqS gene (IG: 4.213) suggests the model captures the evolutionary trajectory of resistance where enhanced macrophage survival precedes the accumulation of primary resistance markers [37] Significantly, the model identified an integrated metabolic program for cell-wall reinforcement.
+High-attribution scores for ppsD, fadD26, and papA4, members of a co-transcribed hub, reflect the biosynthesis of the Phthiocerol Dimycocerosates (PDIM) waxy coat [2, 43].
+The prioritization of these loci, known to be upregulated in rpoB mutants, demonstrates that VAMP-Net has internalized the compensatory permeability barriers employed by M.
+tuberculosis under drug pressure.
+These results confirm that the framework reconstructs a biologically grounded, multi-layered defense strategy rather than relying on arbitrary statistical correlations.
+Table 16: Predictive attribution and biological functional validation of representative Rifabutin (RFB) resistance markers identified by VAMP-Net.
+Gene Variant IG Score Category Functional Mechanism Reference
+rpoB 761140_A>>T 6.896 Primary Canonical RNA Polymerase target; different structural [21]
+761139_C>>A 3.955 Resistance affinities for RFB vs. RIF yield distinct mutation profiles.
+mmpL5 2944085_C>>T 4.947 Efflux Pump Broad-spectrum efflux (Bedaquiline/Clofazimine); [18]
+likely contributes to non-specific RFB resistance.
+lpqS 893702_GAC… 4.213 Persistence Resists copper toxicity in macrophages; enhances [37]
+persistence and subsequent mutation accumulation.
+ppsD 3640407_G>>A 3.532 Cell Wall Essential for PDIM (waxy coat) biosynthesis; directly [2]
+(PDIM) upregulated in rpoB mutants to enhance permeability barrier.
+fadD26 3545402_A>>G 3.527 Cell Wall Initiates PDIM biosynthesis; co-transcribed with ppsD [43]
+(PDIM) to reinforce the outer waxy envelope.
+pks13 1262023_C>>T 2.885 Cell Wall Final catalyst for mycolic acid synthesis; primary [15]
+1262193_C>>T 2.589 (Mycolic) barrier layer essential for natural drug resistance.
+papA4 4217696_TGG.. 2.594 Cell Wall Anchors PDIM to the surface; loss of function leads [19]
+(PDIM) to envelope mislocalization and drug vulnerability.
+To validate the biological coherence of the identified novel resistance loci, we performed functional enrichment analysis using STRING-db 15.
+The high-attribution genes (ponA1, pknB, mmpL5, lpqS, ppsD, fadD26, pks13, papA2) formed a statistically significant interaction network ($p=0.00239$), confirming they constitute a biologically non-random module.
+Gene Ontology (GO) analysis identified six enriched biological processes converging on cell wall remodeling as a unified adaptive strategy: Significant enrichment in "Cell wall organization or biogenesis" ($FDR=7.16\times 10^{-5}$) and "Peptidoglycan-based cell wall biogenesis" ($FDR=0.00056$) directly validates the RIF-associated loci ponA1 and pknB; enrichment in "DIM/DIP cell wall layer assembly" ($FDR=0.0076$) and "Phthiocerol biosynthetic process" ($FDR=0.0257$) confirms that the model autonomously identified the precise PDIM/phthiocerol lipid pathways.
+High enrichment strength for the "Phthiocerol biosynthetic process" (strength = 2.22) biochemically confirms the functional assembly of ppsD, fadD26, and pks13.These results demonstrate that VAMP-Net‘s discovery reflects the coordinated upregulation of PDIM biosynthesis and cell wall reinforcement as an adaptive epistatic response to rpoB-mediated resistance.
+Figure 15: Functional protein-protein interaction network and GO enrichment of novel loci identified for RIF/RFB resistance.
+The interaction map (top) highlights two significant metabolic modules discovered via VAMP-Net‘s attention weights: the ppsD-pks13-fadD26 cluster and the ponA1-pknB axis.
+The Gene Ontology analysis (bottom) confirms these loci are statistically enriched for cell wall organization ($p=7.16\times 10^{-5}$) and PDIM biosynthesis ($p=0.0257$).
+These results demonstrate VAMP-Net’s ability to recover biologically cohesive compensatory mechanisms alongside canonical resistance markers.
+Functional enrichment analysis of EMB-associated loci yielded the study‘s strongest validation result 16, centered on Arabinosyltransferase activity (GO:0052636).
+This term achieved an exceptional enrichment signal of 4.0 ($FDR=3.44\times 10^{-7}$), capturing 4 out of 4 known arabinosyltransferase-encoding genes in the M.
+tuberculosis genome.This constitutes a near-perfect biological validation, as ethambutol (EMB) specifically targets the arabinosyltransferase enzymes, encoded by embA, embB, and embC—responsible for cell wall synthesis.
+VAMP-Net autonomously recovered this entire enzyme family along with the ubiA locus, a known resistance gene providing the necessary substrate.
+Furthermore, "Cell wall organization" (GO:0071555) was significantly enriched ($FDR=0.0053$), consistent with EMB’s role in disrupting cell wall assembly.
+These findings confirm that the model learns integrated resistance pathways, comprising primary targets, substrate supply, and potential epistatic modifiers, rather than isolated variant associations.
+Figure 16: Functional interaction network and GO enrichment for Ethambutol (EMB) resistance loci.
+The interaction map (top) illustrates a tightly coupled module comprising embA, embB, embC, and ubiA, which are primary drivers of EMB resistance.
+Gene Ontology analysis (bottom) confirms high statistical enrichment for arabinosyltransferase activity ($p=3.44\times 10^{-7}$) and cell wall organization ($p=0.0053$).
+These results validate VAMP-Net’s ability to precisely identify the specific metabolic pathways targeted by the drug.
+4.5.3 Epistatic Interaction Network Analysis
+Analysis of the aggregated Set Attention Block (SAB) weights culminates in the Multi-locus Attribution-Weighted Interaction (MAWI) Map.This approach quantifies the synergistic coordination between mutations, moving beyond single-variant associations to decompose the polygenic architecture of resistance.
+The interaction heatmap for Rifampicin (RIF) reveals a highly interconnected network with high signal density (Figure 17).
+This structure is characterized by "hub" variants, such as 3829770_T>>C and 1307598_C>>G, which exhibit strong cross-attention with multiple secondary mutations.
+This hub-driven topology quantitatively supports the hypothesis of epistatic fitness compensation, where robust resistance relies on a coordination of compensatory pathways to mitigate the fitness costs of primary mutations.
+Figure 17: Interpretable attention heatmap of variant-variant interactions for Rifampicin (RIF) resistance.
+The matrix visualizes the self-attention weights derived from the Set Attention Block (SAB), representing the perceived strength of interaction between genomic variants.
+Darker red clusters identify "hub variants" that exert high influence over the collective resistance profile.
+These patterns reveal that VAMP-Net identifies non-linear epistatic relationships beyond simple additive effects, aligning with the biological understanding that complex resistance phenotypes often emerge from the co-occurrence of specific compensatory and primary mutations.
+In contrast, the Rifabutin (RFB) network exhibits a sparse, localized architecture.
+The attention landscape is dominated by a few high-leverage interactions involving variants like 1070702_T>>C, set against a low-interaction background.
+This sparsity indicates that RFB resistance is governed by a select few non-additive synergistic effects that are mechanistically determinative, rather than the diffuse polygenic mechanism observed in RIF.
+The contrasting patterns between these two rifamycins suggest that while they share target genes, their resistance modules are distinct.
+VAMP-Net’s ability to pinpoint these drug-specific interaction hubs provides an interpretable, auditable trace of the model’s reasoning.
+These results confirm known biological dependencies while highlighting novel epistatic interaction points that could serve as targets for future functional validation and the design of more effective combination therapies.
+To confirm that the high-attention mutation pairs identified by the SAB transformer reflect genuine genomic dependencies rather than modeling artifacts, we performed a population-level validation using Fisher’s exact test and Odds Ratio (OR) analysis (Table 17).
+This approach provides an external statistical benchmark for the non-random dependencies captured by the attention mechanism.
+As detailed in Table Table 17, the statistical results for the primary interaction candidates demonstrate robust data-driven support:
+  • RIF-associated pair ($1276232:C{>}A\leftrightarrow 3600576:C{>}T$): The combination of an infinite OR and an extremely low $p$-value ($9.296\times 10^{-89}$) indicates a highly structured co-occurrence.
+This suggests possible genetic linkage or strong co-selection under drug pressure, where the presence of one mutation is effectively determinative of the other.
+  • RFB-associated pair ($1070702:T{>}C\leftrightarrow 985287:G{>}A$): The elevated odds ratio of $8.18$ ($p<0.01$) indicates significant enrichment, confirming that these mutations co-occur substantially more frequently than expected under independence.
+These findings provide a critical layer of interpretability for VAMP-Net.
+By demonstrating that the highest-weighted interaction hubs are backed by significant population-level enrichment, we confirm that the model recovers mutation pairs consistent with the underlying structure of the M.
+tuberculosis population.
+This validates the attention heatmap as an auditable biological trace, effectively pinpointing non-random dependencies that may serve as targets for further functional investigation into epistatic fitness or shared resistance backgrounds.
+Table 17: Statistical validation of representative high-attention mutation pairs using Fisher’s exact test. Only mutation pairs with strong statistical support are shown.
+Drug Mutation pair OR $p$-value Interpretation
+RIF $1276232:{C{>}A}\leftrightarrow 3600576:{C{>}T}$ $\infty$ $9.296\times 10^{-89}$ Extremely strong non-random co-occurrence; consistent with tight genetic linkage, co-selection, or a shared resistance background.
+RFB $1070702:{T{>}C}\leftrightarrow 985287:{G{>}A}$ 8.18 $7.175\times 10^{-3}$ Significant enrichment, indicating a robust association beyond chance.
+4.5.4 VCF Feature Channel Analysis
+To validate the multi-channel design of the Quality-Aware 1D-CNN pathway, we quantified the contribution of eight VCF features through systematic ablation (Figure 18).
+The analysis confirms that the model transcends simple variant presence (GT) by leveraging granular contextual metrics:
+  • Dominant Metric (FRS): The Fraction of Supporting Reads (FRS) is the most influential feature across both RIF and RFB models.
+This confirms the pathway effectively filters low-support variants, aligning with clinical standards for calling clonal mutations.
+  • Contextual Confidence: GT_CONF_PERCENTILE serves as the second most critical feature.
+By prioritizing relative confidence ranks over raw GT_CONF values, the model demonstrates a sophisticated capacity to utilize sample-specific comparative context.
+  • Redundancy of Raw Metrics: The near-zero scores for Read_Depth (DP) and Genotype (GT) suggest that their information is successfully subsumed by normalized, derived features.
+These results demonstrate that the 1D-CNN pathway successfully extracts technically meaningful information, enabling a robust distinction between true resistance-conferring variants and sequencing noise.
+Figure 18: Quantitative importance of technical quality metrics across EMB, INH, RIF, and RFB resistance prediction.
+The upper panel compares EMB (orange) and INH (blue), while the lower panel compares RIF (orange) and RFB (green).
+Across phenotypes, Fraction of Supporting Reads (FRS) and Genotype Confidence Percentile (GT_CONF_PERCENTILE) emerge as dominant technical predictors, indicating that VAMP-Net consistently leverages sequencing-confidence cues to recalibrate resistance predictions.
+4.6 Discussion and Limitations
+4.6.1 Architectural Efficacy and Integration Rationale
+VAMP-Net demonstrates exceptional predictive fidelity across the anti-TB drug cohort, specifically achieving peak ROC-AUCs of 0.975 (RFB) and 0.970 (INH).
+The remarkably low variance across experimental seeds (std << 0.02) confirms that this performance is a robust consequence of the model’s structural design rather than stochastic initialization.
+Even in the presence of significant class imbalance, such as with Ethambutol (19.8% resistant), the model achieved the cohort’s highest PR-AUC (0.977), proving its resilience to the "curse of dimensionality" and skewed genomic datasets common in clinical settings.
+Comparative benchmarking reveals a significant performance gap between VAMP-Net and traditional baselines, with the MLP failing to exceed 85.88% accuracy.
+While the CNN effectively captured local spatial patterns, its higher false-negative rate (10.6%) relative to VAMP-Net (7.1%) underscores the necessity of attention mechanisms for modeling long-range epistatic interactions.
+VAMP-Net’s superior sensitivity in identifying 92.9% of resistant isolates while maintaining high specificity (91.3%) validates its utility as a high-precision diagnostic tool for complex antimicrobial resistance landscapes.
+Ultimately, the model’s success is rooted in its late-stage adaptive fusion philosophy.
+In contrast to Early Fusion, which suffered from a catastrophic specificity deficit (0.425) due to premature noise propagation, VAMP-Net’s multi-path design allows for modality-specific representational learning.
+By postponing integration until after contextual encoding, the framework achieved transformative gains, including a +48.8% increase in Specificity and a +30.9% increase in ROC-AUC.
+These results provide a mathematically grounded justification for VAMP-Net as a stable, noise-resilient framework for rapid molecular-based tuberculosis diagnostics.
+4.6.2 VCF-Aware Recalibration and Clinical Utility
+The granular case studies presented in Table 11 provide a critical window into the decision-making process of VAMP-Net, highlighting the Path-2 branch as a dynamic reliability filter.
+The substantial magnitude of observed probability shifts ($\Delta P\approx$ - 0.86) indicates that VCF-level metadata does not merely refine the genomic signal but fundamentally re-weights evidence based on technical fidelity.
+In Corrected cases, while Path-1 identifies resistance-like mutation patterns, Path-2 detects a conflict where weak sequencing support, characterized by low GT_CONF_PERCENTILE or insufficient read depth, renders the initial call unreliable.
+Conversely, in Amplified cases, Path-2 acts as a force-multiplier for high-quality data, shifting borderline predictions ($P(R)\approx$ 0.55) to high-confidence resistant calls (>> 0.90).
+By distinguishing between true biological polymorphisms and sequencing artifacts, VAMP-Net effectively treats variant presence as a necessary, but not sufficient, condition for resistance prediction.
+Beyond raw accuracy, VAMP-Net addresses the practical barriers to AI adoption in healthcare: calibration and operational efficiency.
+Brier score analysis confirms that VAMP-Net is a well-calibrated classifier, particularly for INH ($\sim 0.05$).
+High calibration ensures that predicted probabilities reflect the true likelihood of resistance, thereby reducing the clinical risk of suboptimal patient treatment.
+Furthermore, the architecture maintains a high-speed inference rate of $\sim 490$ samples per second.
+This computational efficiency, combined with high reproducibility and the reduction of grey-zone predictions through Path-2 recalibration, positions VAMP-Net as a reliable foundational tool for the next generation of rapid, molecular-based tuberculosis diagnostics.
+4.6.3 Interpretability and Mechanistic Discovery
+VAMP-Net’s dual-path architecture provides a mathematically grounded framework for identifying high-impact resistance mechanisms by effectively bridging high-dimensional genomics with biological and technical insights.
+Feature attribution analysis successfully recovered canonical resistance signatures, such as rpoB variants for RIF and RFB (e.g., 761140_A>>T; Score: 6.896), while internalizing complex co-selection patterns in MDR-TB strains by prioritizing cross-drug markers like katG (Score: 0.446).
+Beyond known targets, the Set Attention Block ($SAB$) identified novel, high-effect candidates—including the mmuM locus for INH (Score: 8.195) and the mmpL5/ppsD axis for RFB—which functional enrichment confirms constitute a statistically significant ($p=0.00239$) metabolic module centered on cell-wall remodeling and PDIM biosynthesis.
+Furthermore, systematic ablation of VCF features validates the 1D-CNN’s role as a signal-to-noise regulator that prioritizes the Fraction of Supporting Reads (FRS) and relative confidence rankings over raw depth.
+This "integrated audit" ensures that final predictions are governed by a synergy of genomic pathogenicity and technical reliability, allowing the model to dynamically down-weight biologically relevant variants that lack sufficient sequencing support.
+Collectively, these findings position VAMP-Net as an auditable discovery tool capable of reconstructing the multi-layered defense strategy of M.
+tuberculosis through the lens of both evolutionary selection and clinical data quality.
+4.6.4 Limitations of the Study
+While VAMP-Net demonstrates high predictive and interpretability performance, its scope is bounded by several factors.
+First, although our tokenization strategy effectively captures rare variants, generalizability remains constrained by the training cohort’s diversity; emerging mutations or rare lineages absent from the dataset may degrade performance.
+Second, while the 1D-CNN pathway successfully filters technical noise, it does not currently address broader systemic errors such as metagenomic contamination or sample-swapping in clinical workflows.
+Finally, the attention-based feature attribution provides a computational "causality hypothesis" rather than biochemical proof.
+The identified epistatic interactions and novel loci, though statistically significant, require experimental validation through site-directed mutagenesis or functional assays to confirm their mechanistic roles
 5 Conclusion
-We have introduced the Interpretable Variant-Aware Multi-Path Network (VAMP-Net), a novel deep learning framework that successfully addresses two critical, often-overlooked challenges in genomic machine learning: modeling genetic epistasis and systematically incorporating variant calling quality into the prediction process.
-Our work delivers a significant methodological advance that provides both state-of-the-art predictive performance and crucial layers of auditable interpretability, essential for bridging computational genomics with clinical utility.
-The core of VAMP-Net lies in its dual-path architecture and the Adaptive Fusion mechanism.
-The Set Attention Transformer (Path-1) proved highly effective by operating on the inherently permutation-invariant set of ChromPos_Ref>>Alt variant tokens, successfully elucidating complex genetic epistasis and discovering both canonical resistance drivers and influential non-canonical loci.
-The Integrated Gradients attribution analysis provides a clear, ranked list of mutations and their co-dependencies, forming a direct basis for targeted functional validation.
-However, the most compelling finding lies in the contribution of the Quality-Aware pathway.
-VAMP-Net is the first model to systematically integrate variant calling quality into resistance prediction, addressing the critical gap where existing methods treat all variant calls equally regardless of technical confidence.
-Our quantitative feature importance scores demonstrate that the fusion architecture performs a learned feature calibration and noise regularization .
-Specifically, the pronounced increase in reliance on VCF quality metrics (e.g., FRS, GT_CONF_PERCENTILE) for the more complex Rifabutin (RFB) phenotype, compared to Rifampicin (RIF), confirms the fusion mechanism intelligently and dynamically adapts its weighting based on the complexity and confidence requirements of the task.
-This dynamic weighting mechanism provides an unprecedented level of robustness.
-The VAMP-Net framework’s fundamental design is inherently generalizable and extends well beyond the context of Multi-drug resistance in M.
-tuberculosis.
-Our proposal can be applied to any genotype-phenotype problem characterized by three conditions:
-  • Variable-Length Inputs: The sample is defined by an unordered set of symbolic events (e.g., structural variants, somatic mutations in cancer, or transcriptomic changes).
-  • Epistatic Dependencies: The final phenotype is driven by complex interactions rather than single-gene effects.
-  • Heterogeneous Evidence: The symbolic input (genotype) is accompanied by quantitative, context-specific technical metadata (e.g., sequencing depth, coverage, or population frequency).
-Future work will thus focus on extending this multi-path fusion paradigm to incorporate additional modalities in other clinical domains, such as incorporating image modalities (e.g., microscopy or histology) or longitudinal clinical metadata alongside genomic data.
-This work establishes a new standard for confidence and transparency in clinical AI systems by ensuring that all predictions are not only accurate but are driven by a mechanism that is transparent and auditable across both biological and technical evidence streams.
+This study introduced the Interpretable Variant-Aware Multi-Path Network (VAMP-Net), a model designed to bridge the gap between high-dimensional genomic pathogenicity and technical data reliability.
+By integrating a Permuutation-Invariant Set Attention Transformer for epistatic modeling with a Quality-Aware 1D-CNN, VAMP-Net addresses the critical challenge of separating biological signals from sequencing noise.
+Our results demonstrate that the dual-path architecture successfully recovers canonical resistance markers while discovering novel loci, such as mmuM and mmpL5, within statistically significant metabolic modules centered on cell-wall remodeling.
+The most compelling methodological advancement is the model’s capacity for adaptive feature calibration.
+Quantitative attribution analysis confirms that the fusion mechanism dynamically modulates its reliance on quality metrics (e.g., FRS and GT_CONF_PERCENTILE) based on the phenotypic complexity, providing an unprecedented level of decision robustness for subtle resistance profiles like Rifabutin.
+This "integrated audit" ensures that all predictions are supported by both biological and technical evidence streams, establishing a new standard for transparency in clinical AI.
+Future work will focus on extending this multi-path paradigm to multi-modal clinical settings.
+By incorporating longitudinal metadata and medical imaging (e.g., histology or microscopy) alongside genomic sets, we aim to evolve VAMP-Net into a comprehensive diagnostic engine.
+Ultimately, this work provides a scalable, mathematically grounded template for genotype-phenotype association in any domain characterized by variable-length inputs, complex epistasis, and heterogeneous data quality.
+Acknowledgement
+This work was carried out partially using the AI Datacenter of the National School of Artificial Intelligence, funded under grant number E049 24 0117 by the Algerian Ministry of Higher Education and Scientific Research.
 Data availability statements
 The complete dataset used in this study is publicly available at ftp.ebi.ac.uk/pub/databases/cryptic/release_june2022/.
+Code availability
+The complete code of this study can be found at https://github.com/Aicha-AI/VAMP-Net.MTB-Drug-Resistance/tree/main.